### v1

# Differentially Private Adaptation of Diffusion Models via Noisy Aggregated Embeddings

Abstract
We introduce novel methods for adapting diffusion models under differential privacy (DP) constraints, enabling privacy-preserving style and content transfer without fine-tuning. Traditional approaches to private adaptation, such as DP-SGD, incur significant computational overhead and degrade model performance when applied to large, complex models. Our approach instead leverages embedding-based techniques: Universal Guidance and Textual Inversion (TI), adapted with differentially private mechanisms. We apply these methods to Stable Diffusion for style adaptation using two private datasets: a collection of artworks by a single artist and pictograms from the Paris 2024 Olympics. Experimental results show that the TI-based adaptation achieves superior fidelity in style transfer, even under strong privacy guarantees, while both methods maintain high privacy resilience by employing calibrated noise and subsampling strategies. Our findings demonstrate a feasible and efficient pathway for privacy-preserving diffusion model adaptation, balancing data protection with the fidelity of generated images, and offer insights into embedding-driven methods for DP in generative AI applications.
1 Introduction
In recent years, diffusion models ho2020denoising ; song2020score , particularly latent diffusion models rombach2022high , have spearheaded high quality text-to-image generation, and have been widely adopted by researchers and the general public alike. Trained on massive datasets like LAION-5B schuhmann2022laion , these models have developed a broad understanding of visual concepts, enabling new creative and practical applications. Notably, tools like Stable Diffusion rombach2022high ; podell2023sdxl have been made readily accessible for general use. Building on this foundation, efficient adaptation methods such as parameter efficient fine-tuning (PEFT) hu2022lora ; von-platen-etal-2022-diffusers ; ruiz2023dreambooth , guidance based approaches ho2021classifierfree ; kim2022diffusionclip ; bansal2023universal , and pseudo-word generation gal2023an enable users to leverage this extensive pretraining for customizing models that can specialize on downstream tasks with smaller datasets.
However, the rapid adoption of diffusion models has also raised significant privacy, ethical and legal concerns. One critical issue is the vulnerability of these models to privacy attacks, from membership inference duan2023diffusion , where an attacker determines whether a specific data point was used to train a particular model, to data extraction carlini2023extracting , which enables an attacker to reconstruct particular images from the training dataset. This issue is even more severe during the fine-tuning phase where the model is fine-tuned on smaller specialized datasets from a possibly different domain and each data record has more impact on the final model. This risk underscores the importance of privacy-preserving technologies, particularly as diffusion models often rely on vast datasets scraped from the internet without explicit consent from content owners. Copyright concerns further complicate the landscape vyas2023provable . In particular, diffusion models can closely reproduce original artworks without credit or compensation to the artist, as evident in data extraction attacks carlini2023extracting . Such concerns intensify the debate on balancing technological advancement with the rights and privacy of content owners, underscoring the need for clearer ethical and legal guidelines in the training and deployment of generative models.
Differential privacy (DP) dwork2006calibrating ; dwork2006differential is a widely adopted approach to addressing these challenges, where controlled noise is added during training or inference to prevent information leakage from individual data points while still enabling the model to learn effectively from the overall dataset. The standard method for ensuring DP in deep learning is Differentially Private Stochastic Gradient Descent (DP-SGD) abadi2016deep , which modifies traditional SGD by adding noise to clipped gradients.
While, in theory, DP-SGD can be applied to train diffusion models, it poses several significant practical challenges. First, it is well known that DP-SGD introduces considerable computational and memory overhead due to the need for per-sample gradient clipping, which is crucial for bounding sensitivity dwork2006calibrating ; abadi2016deep . Second, DP-SGD is fundamentally incompatible with batch-wise operations, such as batch normalization, since these operations inherently link samples together, rendering sensitivity analysis infeasible. Moreover, training a large network with DP-SGD typically leads to significant performance degradation (i.e. the curse of dimensionality), even under relatively lenient privacy budgets. To the best of our knowledge, existing diffusion models trained with DP-SGD are constrained to relatively small-scale datasets dockhorn2023differentially ; ghalebikesabi2023differentially .
As a result, recent research has increasingly turned to alternative strategies for adapting diffusion models in a privacy-preserving manner without the need for training from scratch under DP-SGD. A promising approach leverages large, publicly pre-trained models, adapting them to new domains under differential privacy (DP) constraints, which allows models to capitalize on existing representational strengths while avoiding the computational and memory burdens of full DP training ghalebikesabi2023differentially . Similarly, parameter-efficient fine-tuning (PEFT) methods, such as DP-LoRA yu2021differentially , fine-tune only a subset of parameters, allowing for efficient, targeted adaptation with limited privacy costs. Methods like DP-RDM (Differentially Private Retrieval-Augmented Diffusion Model) lebensold2024dp illustrate this trend by enabling model adaptation through a retrieval mechanism that conditions image generation on private data retrieved at inference time, avoiding direct model updates. These approaches collectively demonstrate how large models can be adapted privately, circumventing the complexities and high costs associated with full DP training.
Meanwhile, alternative methods such as Universal Guidance bansal2023universal and Textual Inversion (TI) gal2023an enable model adaptation to specific content or style without modifying the diffusion model itself, instead relying on external embedding vectors. Universal Guidance’s style guidance (SG) mechanism generates a target CLIP embedding by passing the reference image through a CLIP image encoder. During the reverse diffusion process, the cosine distance between this target embedding and the CLIP embedding of the generated image is minimized, guiding the generation process to match the desired style. TI, in contrast, learns a new token embedding representing a set of target images, which is then incorporated into the text prompt to adapt the model to the style or content associated with those images. Both SG and TI avoid direct model optimization, reducing memory and computation demands while also involving fewer trainable parameters, making them highly compatible with DP constraints on smaller datasets.
In this work, we propose two novel privacy-preserving adaptation methods for smaller datasets, leveraging SG and TI to bypass the need for DP-SGD or DP-PEFT approaches that require extensive model updating. Our first approach adapts Universal Guidance for privacy by using a differentially private CLIP embedding. Although SG is typically applied to a single target image, we extend it to handle multiple images under DP by computing an embedding for each target image and aggregating these embeddings via a centroid calculation. To ensure differential privacy, we add calibrated noise to the centroid, which protects individual data points in the dataset while preserving the model’s guidance capabilities.
Our second approach enhances Textual Inversion (TI) with differential privacy. While standard TI compresses training data into a single, low-dimensional vector, offering some intrinsic privacy benefits, it lacks the rigorous guarantees required by DP. We address this by introducing a private variant of TI, in which we decouple interactions among samples by learning a separate embedding for each target image. These embeddings are then aggregated into a noisy centroid, similar to our SG approach, ensuring differential privacy while preserving TI’s data compression advantages.
Our experimental results underscore the practicality and effectiveness of the proposed differentially private embedding-based adaptation methods, with Textual Inversion (TI) proving particularly robust for preserving stylistic fidelity even under privacy constraints. In applying our approach to two distinct datasets–a private collection of artworks by @eveismyname and pictograms from the Paris 2024 Olympics Paris2024Pictograms –we demonstrate that differentially private TI can successfully capture nuanced stylistic elements while maintaining privacy guarantees. Our findings reveal a trade-off between privacy levels (determined by the DP parameter $\varepsilon$) and image quality, where lower $\varepsilon$ values yield diminished stylistic fidelity, yet still preserve the target style under moderate noise. Furthermore, subsampling offers an effective means of improving privacy resilience by controlling sensitivity to individual data points, significantly mitigating the impact of added noise on image quality. Ultimately, our approach provides a versatile framework for privacy-preserving model adaptation, enabling diffusion models to generalize effectively to novel styles and domains without risking exposure of sensitive data points.
2 Background and Related Works
2.1 Diffusion Models
Diffusion models ho2020denoising ; song2020score ; song2020denoising ; rombach2022high leverage an iterative denoising process to generate high-quality images that aligns with a given conditional input from random noise. In text-to-image generation, this conditional input is based on a textual description (a prompt) that guides the model in shaping the image to reflect the content and style specified by the text. To convert the text prompt into a suitable conditional format, it is first tokenized into discrete tokens, each representing a word or sub-word unit. These tokens are then converted into a sequence of embedding vectors $v_{i}$ that encapsulate the meaning of each token within the model’s semantic space. Next, these embeddings pass through a transformer text encoder, such as CLIP radford2021learning , outputting a single text-conditional vector $y$ that serves as the conditioning input. This text-conditioning vector $y$ is then incorporated at each denoising step, guiding the model in aligning the output image with the specific details outlined in the prompt.
The image generation process, also known as the reverse diffusion process, comprises of $T$ discrete timesteps and starts with pure Gaussian noise $x_{T}$. At each decreasing timestep $t$, the denoising model, which often utilizes a U-Net structure with cross-attention layers, takes a noisy image $x_{t}$ and text conditioning $y$ as inputs and predicts the noise component $\epsilon_{\theta}(x_{t},y,t)$, where $\theta$ denotes the denoising model’s parameters. The predicted noise is then used to make a reverse diffusion step from $x_{t}$ to $x_{t-1}$, iteratively refining the noisy image closer to a coherent output $x_{0}$ that aligns with the text conditional $y$.
In particular, denoising diffusion implicit models (DDIM) sampling song2020denoising uses the predicted noise $\epsilon_{\theta}(x_{t},y,t)$ and a noise schedule represented by an array of scalars $\{\alpha_{t}\}_{t=1}^{T}$ to first predict a clean image $\hat{x}_{0}$, then makes a small step in the direction of $\hat{x}_{0}$ to obtain $x_{t-1}$. The reverse diffusion process for DDIM sampling can be formalized as follows:
$\hat{x}_{0}=\frac{x_{t}-\sqrt{1-\alpha_{t}}\epsilon_{\theta}(x_{t},y,t)}{\sqrt{\alpha_{t}}}$ (1)
$x_{t-1}=\sqrt{\alpha_{t-1}}\hat{x}_{0}+\sqrt{1-\alpha_{t-1}}\epsilon_{\theta}(x_{t},y,t)$ (2)
The training of a text-conditioned diffusion model involves minimizing a loss function that guides the model to predict the noise added to an image, given both the noisy image $x_{t}$ and the text conditioning $y$. The objective function is typically a mean squared error (MSE) between the true noise $\epsilon$ and the predicted noise $\epsilon_{\theta}(x_{t},y,t)$. The denoising model is therefore trained over the following optimization problem:
$\theta^{*}=\arg\min_{\theta}\mathbb{E}_{x,\epsilon\sim\mathcal{N}(0,I),t\sim[T]}\left[\|\epsilon-\epsilon_{\theta}(x_{t},y,t)\|^{2}\right]$ (3)
2.1.1 Style Guidance
Bansal et al. bansal2023universal propose a universal guidance algorithm that allows the reverse diffusion process to be guided by arbitrary guidance functions, including object detection, image segmentation and facial recognition. In this section, we formally describe CLIP style guidance, one of the many processes explored in the universal guidance study. We also only describe the forward guidance process. For a generalized and complete description of universal guidance process, including the backward guidance process and per-step self-recurrence, we defer the reader to the original paper.
Let $x_{c}$ denote a target image whose style we wish to replicate and $\textsf{CLIP}(\cdot)$ denote a CLIP image encoder. To incorporate style guidance, given a noisy image $x_{t}$, we first compute $\hat{x}_{0}$ using the predicted noise $\epsilon_{\theta}(x_{t},y,t)$ and Equation (1). Using $\hat{x}_{0}$, we then compute the following:
$\hat{\epsilon}_{\theta}(x_{t},y,t)=\epsilon_{\theta}(x_{t},y,t)+w\sqrt{1-\alpha_{t}}\nabla_{x_{t}}\ell_{\cos}(\textsf{CLIP}(x_{c}),\textsf{CLIP}(\hat{x}_{0}))$ (4)
where $w$ is a guidance weight parameter and $\ell_{\cos}(\cdot,\cdot)$ is the negative cosine loss.
We then perform the reverse diffusion step using the same procedure given in Equations (1) and (2), but replacing $\epsilon_{\theta}(x_{t},y,t)$ with $\hat{\epsilon}_{\theta}(x_{t},y,t)$. Using this reverse diffusion procedure, we generate a high-quality image $x_{0}$ that aligns with the text conditioning $y$ while also optimizing the cosine similarity between the CLIP embeddings of $x_{c}$ and $x_{0}$, resulting in $x_{0}$ having the stylistic characteristics of $x_{c}$.
2.1.2 Textual Inversion
Textual Inversion (TI) gal2023an is an adaptation technique that enables personalization using a small dataset of typically only 3-5 images. This approach essentially learns a new token that encapsulates the semantic meaning of the training images, allowing the model to associate specific visual features with a custom token.
To achieve this, TI trains a new token embedding, denoted as $u$, which represents a placeholder token, often denoted as $S$. During training, images are conditioned of phrases such as “A photo of $S$” or “A painting in the style of $S$”. However, unlike the fixed embeddings of typical tokens $v_{i}$, $u$ is a learnable parameter. Let $y_{u}$ denote the text conditioning vector resulting from a prompt containing the token $S$. Through gradient descent, TI minimizes the diffusion model loss with respect to $u$, while keeping the diffusion model parameters $\theta$ fixed, iteratively refining this embedding to capture the unique characteristics of the training images. The resulting optimal embedding $u_{*}$ is formalized as follows:
$u_{*}=\arg\min_{u}\mathbb{E}_{x,\epsilon\sim\mathcal{N}(0,I),t\sim[T]}[\|\epsilon-\epsilon_{\theta}(x_{t},y_{u},t)\|^{2}].$ (5)
Hence, $u_{*}$ represents an optimized placeholder token $S_{*}$, which can employed in prompts such as “A photo of $S_{*}$ floating in space” or “A drawing of a capybara in the style of $S_{*}$”, enabling the generation of personalized images that reflect the learned visual characteristics.
2.2 Differential Privacy
In this work, we adopt differential privacy (DP) dwork2006calibrating ; dwork2006differential as our privacy framework. Over the past decade, DP has become the gold standard for privacy protection in both research and industry. It measures the stability of a randomized algorithm concerning changes in an input instance, thereby quantifying the extent to which an adversary can infer the existence of a specific input based on the algorithm’s output.
Definition 2.1 ((Approximate) Differential Privacy)
For $\varepsilon,\delta\geq 0$, a randomized mechanism $\mathcal{M}:\mathcal{X}^{n}\rightarrow\mathcal{Y}$ satisfies $(\varepsilon,\delta)$-DP if for all neighboring datasets $\mathcal{D},\mathcal{D}^{\prime}\in\mathcal{X}^{n}$ which differ in a single record (i.e., $\left\lVert\mathcal{D}-\mathcal{D}^{\prime}\right\rVert_{\mathsf{H}}\leq 1$ where $\left\lVert\cdot\right\rVert_{\mathsf{H}}$ is the Hamming distance) and all measurable $\mathcal{S}$ in the range of $\mathcal{M}$, we have that
$\mathbb{P}\left(\mathcal{M}(\mathcal{D})\in\mathcal{S}\right)\leq e^{\varepsilon}\mathbb{P}\left(\mathcal{M}(\mathcal{D}^{\prime})\in\mathcal{S}\right)+\delta.$
When $\delta=0$, we say $\mathcal{M}$ satisfies $\varepsilon$-pure DP or ($\varepsilon$-DP).
To achieve DP, the Gaussian mechanism is often applied dwork2014algorithmic ; balle2018improving , adding Gaussian noise scaled by the sensitivity of the function $f$ and privacy parameters $\varepsilon$ and $\delta$. Specifically, noise with standard deviation $\sigma=\frac{\Delta_{f}\sqrt{2\ln(1.25/\delta)}}{\varepsilon}$ is added to the output [^1] In practice, we use numerical privacy accountant such as balle2018improving ; mironov2017renyi to calibrate the noise. balle2018improving , where $\Delta_{f}$ represents $\ell_{2}$-sensitivity of the target function $f(\cdot)$. When the context is clear, we may omit the subscript $f$. This mechanism enables a smooth privacy-utility tradeoff and is widely used in privacy-preserving machine learning, including in DP-SGD abadi2016deep , which applies Gaussian noise during model updates to achieve DP.
2.2.1 Privacy Amplification by Subsampling
Subsampling is a standard technique in DP, where a full dataset of size $n$ is first subsampled to $m$ records without replacement (typically with $m\ll n$), and the privatization mechanism (such as the Gaussian mechanism) is then applied. Specifically, if a mechanism provides $(\varepsilon,\delta)$-DP on a dataset of size $m$, it achieves $(\varepsilon^{\prime},\delta^{\prime})$-DP on the subsampled dataset, where $\delta^{\prime}=\frac{m}{n}\delta$ and
$\varepsilon^{\prime}=\log\left(1+\frac{m}{n}\left(e^{\varepsilon}-1\right)\right)=O\left(\frac{m}{n}\varepsilon\right).$ (6)
This result is well-known (see, for example, (steinke2022composition, , Theorem 29)). Notably, if the DP mechanism is Gaussian, even tighter amplification bounds can be applied mironov2017renyi .
2.3 Differentially Private Adaptation of Diffusion Models
Recent advancements in applying differential privacy (DP) to diffusion models have aimed to balance privacy preservation with the high utility of generative outputs. Dockhorn et al. dockhorn2023differentially proposed a differentially private diffusion model (DP diffusion) that enables privacy-preserving generation of realistic samples, setting a foundational approach for adapting diffusion processes under DP constraints. Another common strategy involves training a model on a large public dataset, followed by differentially private finetuning on a private dataset, as explored by Ghalebikesabi et al. ghalebikesabi2023differentially . While effective in certain contexts, this approach raises privacy concerns, particularly around risks of information leakage during the finetuning phase pmlr-v235-tramer24a .
In response to these limitations, various adaptation techniques have emerged. Although not specific to diffusion models, some methods focus on training models on synthetic data followed by DP-constrained finetuning, as in the VIP approach yu2024vip , which demonstrates the feasibility of applying DP in later adaptation stages. Other approaches explore differentially private learning of feature representations sander2024differentially , aiming to distill private information into a generalized embedding space while maintaining DP guarantees. Although these adaptations are not yet implemented for diffusion models, they lay essential groundwork for developing secure and efficient privacy-preserving generative models.
3 Differentially Private Adaptation via Noisy Aggregated Embeddings
3.1 General Approach
Let $x^{(1)},\dots,x^{(n)}$ represent a target dataset of images whose characteristics we wish to privately adapt our image generation towards. Assuming that we can encode each image $x^{(i)}$ as a directional embedding vector $u^{(i)}$, we can aggregate the embeddings $u^{(1)},\dots,u^{(n)}$ by calculating the centroid. The purpose of this aggregation is to limit the sensitivity of the final output to each $x^{(i)}$. In order to provide DP guanrantees, we also add isotropic Gaussian noise to the centroid. We can therefore define the resulting embedding vector $u^{*}$ as follows:
$u^{*}=\frac{1}{n}\sum_{i=1}^{n}u^{(i)}+\mathcal{N}(0,\sigma^{2}I)$ (7)
where the minimum $\sigma$ required to provide $(\varepsilon,\delta)$-DP is given by the following expression based on (balle2018improving, , Theorem 1):
$\sigma=\frac{\Delta}{n}\cdot\frac{\sqrt{2\log(1.25)/\delta}}{\varepsilon}.$ (8)
In the context of our problem, $\Delta=\sup_{i,j}\|u^{(i)}-u^{(j)}\|$. Since our embedding vectors are directional, we can normalize each $u^{(i)}$, allowing us to set $\Delta=2$.
The noisy centroid embedding $u^{*}$ can then be used to adapt the downstream image generation process. In Sections 3.3 and 3.4, we describe how to apply our method to TI and SG respectively, outlining the encoding of $x^{(i)}$ as $u^{(i)}$ and the role of $u^{*}$ in the image generation process. To reduce the amount of noise needed to provide the same level of DP, we employ subsampling: instead of computing the centroid over all $n$ embedding vectors, we randomly sample $m\leq n$ embedding vectors without replacement and compute the centroid over only the sampled vectors. Then the standard privacy amplification by subsampling bounds (such as (6)) can be applied. Formally, we sample $D_{\text{sub}}\subseteq\{u^{(1)},\dots,u^{(n)}\}$ where $|D_{\text{sub}}|=m$, and compute the output embedding as follows:
$u^{*}=\frac{1}{m}\sum_{u^{(i)}\in D_{\text{sub}}}u^{(i)}+\mathcal{N}(0,\sigma^{2}I),$ (9)
where $\sigma$ can be computed numerically for any target $\varepsilon,\delta$ and subsampling rate $\frac{m}{n}$.
3.2 Using Style Guidance
Applying our method to style guidance is straightforward. To obtain the embeddings, we simply use the CLIP image encoder to obtain the embedding for each image. More formally, for each $x^{(i)}$, let $u^{(i)}=\textsf{CLIP}(x^{(i)})$. We can then aggregate the embeddings $u^{(i)}$ to obtain $u^{*}$ via Equation (7) or (9), depending on whether we wish to incorporate subsampling. Finally, we use $u^{*}$ to guide the reverse diffusion process, as outlined in Section 2.1.1, by computing $\hat{\epsilon}_{\theta}(x_{t},y,t)$ via a modification of Equation (2):
$\hat{\epsilon}_{\theta}(x_{t},y,t)=\epsilon_{\theta}(x_{t},y,t)+w\sqrt{1-\alpha_{t}}\nabla_{x_{t}}\ell_{\cos}(u^{*},\textsf{CLIP}(\hat{x}_{0}))$ (10)
3.3 Using Textual Inversion
Textual Inversion (TI) is inherently parameter-efficient and offers certain privacy benefits, as information from an entire dataset of images is compressed into a single token embedding vector. This compression limits the model’s capacity to memorize specific images, making data extraction attacks difficult. However, this privacy is merely heuristic and yet to be proven, so TI may still be vulnerable to privacy attacks such as membership inference. A similar adapation technique with privacy guarantees may therefore be desirable.
To incorporate our approach into TI, instead of training a single token embedding on the entire dataset, we instead train a separate embedding $u^{(i)}$ on each $x^{(i)}$ to obtain a set of embeddings $u^{(1)},\dots,u^{(n)}$. We can formalize the encoding process as follows:
$u^{(i)}=\arg\min_{u}\mathbb{E}_{\epsilon\sim\mathcal{N}(0,I),t}[\|\epsilon-\epsilon_{\theta}(x_{t}^{(i)},y_{u},t)\|^{2}]$ (11)
We can then aggregate the embeddings $u^{(i)}$ to obtain $u^{*}$ via Equation (7) or (9), depending on whether we wish to utilize subsampling. Similar to TI’s $u_{*}$, we can use $u^{*}$ to represent a new placeholder token $S^{*}$ that can be incorporated into prompts for personalized image generation. While $u^{*}$ may not fully solve the TI optimization problem presented in Equation (5), it provides provable privacy guarantees, with only a minimal trade-off in accurately representing the content or style of the training data.
4 Experimental Results
4.1 Datasets
We compiled two datasets for our style adaptation experiments, designed to proxy for private datasets unrecognized by the base Stable Diffusion model. The first dataset consists of 158 artworks created by the artist @eveismyname, who has provided consent for non-commercial use. This dataset serves to examine how models can capture and replicate artistic styles without directly copying individual works. Although we acknowledge the possibility that a few of these artworks may have been inadvertently included in the pre-training of Stable Diffusion due their public presence on social media platforms, the artist’s limited recognition and modest portfolio size make it unlikely that their unique style could be effectively extracted from the pre-trained model. The second dataset consists of 47 pictograms from the Paris 2024 Olympics Paris2024Pictograms , approved strictly for non-commercial editorial use IOCProperties . These pictograms were unveiled in February 2023, several months after the release of Stable Diffusion v1.5, which serves as the base model for our experiments. Through this dataset, we hope to evaluate how effectively our approach adapts to novel and unfamiliar visual patterns. Representative samples from both datasets are presented in Figure 1.
Figure 1: Samples of images used in our style adaptation experiments. Left: artwork by @eveismyname ($n=158$). Right: Paris 2024 Olympic pictograms ($n=47$), © International Olympic Committee, 2023.
4.2 Style Guidance Results
We applied our SG-based bansal2023universal approach, detailed in Section 3.2 to both datasets. While this method successfully provided privacy protection by obfuscating embedding details, the resulting image quality and style capture was notably less compared to the TI-based approach. The CLIP embeddings, even without noise or subsampling, led to images that retained only generalized stylistic elements, lacking the detailed fidelity and coherence achieved with TI. Results of our attempt can be seen in Figure 2. This result underscores the superiority of our TI method for maintaining both privacy and high-quality image generation.
A possible explanation for the difference in style transfer effectiveness is that SG, as with other guidance methods, suffers from instability and sensitive dependency to hyperparameters such as guidance weight $w$. While Bansal et al. proposed using backward guidance and per-step self-recurrence to remedy these issues, we found that they were insufficient for our application. Another possible explanation is that the stylistic signal of the image embedding vectors $\textsf{CLIP}(x^{(i)})$ are simply not resilient to our aggregation process.
Figure 2: Attempts of using universal guidance to generate drawings of Taylor Swift and icons of the Eiffel Tower in the styles of @eveismyname and Paris 2024 Pictograms respectively. Here, we apply no subsampling or DP-noise.
4.3 Textual Inversion Results
Figure 3: Images generated by Stable Diffusion v1.5 using the prompt “A painting of Taylor Swift in the style of <@eveismyname>”, with the embedding <@eveismyname> trained using different values of $m$ and $\varepsilon$.
Figure 4: Images generated by Stable Diffusion v1.5 using the prompt “Icon of the Eiffel Tower in the style of <Paris 2024 Pictograms>”, with the embedding <Paris 2024 Pictograms> trained using different values of $m$ and $\varepsilon$.
Using both the @eveismyname and Paris 2024 pictograms dataset, we trained TI gal2023an embeddings on Stable Diffusion v1.5 rombach2022high using our proposed method detailed in Section 3.3. Our primary goal is to investigate how DP configurations, specifically the $\varepsilon$ and subsampling sizes ($m$), affect the generated images’ quality and privacy resilience. For regular TI, we utilize Stable Diffusion’s default process to embed the private dataset without any additional noise. For the DP-enhanced experiments using our method, we test multiple configurations of $m$ and $\varepsilon$ to analyze the trade-off between image fidelity and privacy resilience.
Figures 3 and  4 present generated images across three key configurations: (1) regular TI without DP, (2) centroid TI with DP at various epsilon values, and (3) centroid TI with DP at different subsampling sizes within the private dataset. We used the same random seed to generate embeddings and sample the images for ease of visual comparision between different configurations. As with common practice, we set $\delta=1/n$. Since $\sigma$ is undefined for $\varepsilon=0$, we demonstrate the results of $\varepsilon\approx 0$, in other words, infinite noise, by setting $\varepsilon=10^{-5}$. Images generated without DP closely resemble the unique stylistic elements of the target dataset. In particular, images adapted using @eveismyname images displayed crisp details and nuanced color gradients characteristic of the artist’s work, while those of Paris 2024 pictograms captured the logo’s original structure. In contrast, DP configurations introduce a discernible degradation in image quality, with lower epsilon values and smaller subsampling sizes resulting in more noticeable noise and diminished stylistic fidelity.
As $\varepsilon\rightarrow 0$, the resulting token embedding $u^{*}$ gradually loses its semantic meaning, leading to a loss of stylistic fidelity. In particular, $y_{u^{*}}$ tends towards $y$ (a conditioning vector independent of the learnable embedding). In our results, this manifests as a painting of Taylor Swift devoid of the artist-specific stylistic elements, or a generic icon of the Eiffel Tower (with color, as opposed to the black and white design of the pictograms). With this in mind, $\varepsilon$ can be interpreted as a drift parameter, representing the progression from the optimal $u^{*}$ towards infinity, gradually steering the generated image away from the target style in exchange for stronger privacy guarantees.
Meanwhile, reducing $m$ also reduces the sensitivity of the generated image to $\varepsilon$, as evident by the observation that, on both datasets at $m=4$, (subsampling rate below $0.1$) image generation can tolerate $\varepsilon$ as low as $0.5$ without significant changes in visual characteristics, and retaining stylistic elements of the target dataset at $\varepsilon$ as low as $0.1$. This strong boost in robustness comes at a small price of base style capture fidelity. As observed in figures 3 and 4, we can also treat subsampling as an introduction of noise. Mathematically, the subsample centroid is an unbiased estimate of the true centroid, and so the subsampling process itself defines a distribution centered at the true centroid. However, the amount of noise introduced by the subsampling process is limited by the individual image embeddings, as a subsample centroid can only stray from the true centroid as much as the most anomalous point in the dataset.
5 Conclusion
We presented a differentially private adaptation approach for diffusion models using embedding-based methods–Universal Guidance and Textual Inversion (TI)–to enable privacy-preserving style transfer from small, sensitive datasets. Through experiments on a private artwork dataset and the Paris 2024 Olympics pictograms, we demonstrated that our approach maintains high stylistic fidelity while providing strong privacy guarantees. Notably, TI proved more effective in capturing detailed stylistic elements compared to SG, especially with added noise and subsampling to protect privacy.
Our findings suggest that embedding-driven methods can serve as efficient and scalable alternatives to DP-SGD for privacy-preserving model adaptation, minimizing computational demands and offering resilience against privacy attacks. This approach provides a balance between style quality and privacy, enabling generative models to adapt responsibly to new domains.
6 Acknowledgments
We sincerely thank Tatchamon Wongworakul (@eveismyname) for graciously providing her artwork for use in this study.

### v2

@@
 Abstract
-We introduce novel methods for adapting diffusion models under differential privacy (DP) constraints, enabling privacy-preserving style and content transfer without fine-tuning.
-Traditional approaches to private adaptation, such as DP-SGD, incur significant computational overhead and degrade model performance when applied to large, complex models.
-Our approach instead leverages embedding-based techniques: Universal Guidance and Textual Inversion (TI), adapted with differentially private mechanisms.
-We apply these methods to Stable Diffusion for style adaptation using two private datasets: a collection of artworks by a single artist and pictograms from the Paris 2024 Olympics.
-Experimental results show that the TI-based adaptation achieves superior fidelity in style transfer, even under strong privacy guarantees, while both methods maintain high privacy resilience by employing calibrated noise and subsampling strategies.
-Our findings demonstrate a feasible and efficient pathway for privacy-preserving diffusion model adaptation, balancing data protection with the fidelity of generated images, and offer insights into embedding-driven methods for DP in generative AI applications.
+We introduce a novel method for adapting diffusion models under differential privacy (DP) constraints, enabling privacy-preserving style and content transfer without fine-tuning model weights.
+Traditional approaches to private adaptation, such as DP-SGD, incur significant computational and memory overhead when applied to large, complex models.
+In addition, when adapting to small-scale specialized datasets, DP-SGD incurs large amount of noise that significantly degrades the performance.
+Our approach instead leverages an embedding-based technique derived from Textual Inversion (TI) and adapted with differentially private mechanisms.
+We apply TI to Stable Diffusion for style adaptation using two private datasets: a collection of artworks by a single artist and pictograms from the Paris 2024 Olympics.
+Experimental results show that the TI-based adaptation achieves superior fidelity in style transfer, even under strong privacy guarantees.
+Machine Learning, ICML
+Figure 1: Overview of DPAgg-TI.
+We first apply Textual Inversion to extract embeddings for each image in the private dataset.
+These embeddings are then aggregated with differentially private mechansim, incorporating subsampling to produce a private embedding $u^{\star}$.
+Finally, images are generated using the corresponding token <$S^{\star}$>.
 1 Introduction
-In recent years, diffusion models ho2020denoising ; song2020score , particularly latent diffusion models rombach2022high , have spearheaded high quality text-to-image generation, and have been widely adopted by researchers and the general public alike.
-Trained on massive datasets like LAION-5B schuhmann2022laion , these models have developed a broad understanding of visual concepts, enabling new creative and practical applications.
-Notably, tools like Stable Diffusion rombach2022high ; podell2023sdxl have been made readily accessible for general use.
-Building on this foundation, efficient adaptation methods such as parameter efficient fine-tuning (PEFT) hu2022lora ; von-platen-etal-2022-diffusers ; ruiz2023dreambooth , guidance based approaches ho2021classifierfree ; kim2022diffusionclip ; bansal2023universal , and pseudo-word generation gal2023an enable users to leverage this extensive pretraining for customizing models that can specialize on downstream tasks with smaller datasets.
-However, the rapid adoption of diffusion models has also raised significant privacy, ethical and legal concerns.
-One critical issue is the vulnerability of these models to privacy attacks, from membership inference duan2023diffusion , where an attacker determines whether a specific data point was used to train a particular model, to data extraction carlini2023extracting , which enables an attacker to reconstruct particular images from the training dataset.
-This issue is even more severe during the fine-tuning phase where the model is fine-tuned on smaller specialized datasets from a possibly different domain and each data record has more impact on the final model.
-This risk underscores the importance of privacy-preserving technologies, particularly as diffusion models often rely on vast datasets scraped from the internet without explicit consent from content owners.
-Copyright concerns further complicate the landscape vyas2023provable .
-In particular, diffusion models can closely reproduce original artworks without credit or compensation to the artist, as evident in data extraction attacks carlini2023extracting .
-Such concerns intensify the debate on balancing technological advancement with the rights and privacy of content owners, underscoring the need for clearer ethical and legal guidelines in the training and deployment of generative models.
-Differential privacy (DP) dwork2006calibrating ; dwork2006differential is a widely adopted approach to addressing these challenges, where controlled noise is added during training or inference to prevent information leakage from individual data points while still enabling the model to learn effectively from the overall dataset.
-The standard method for ensuring DP in deep learning is Differentially Private Stochastic Gradient Descent (DP-SGD) abadi2016deep , which modifies traditional SGD by adding noise to clipped gradients.
-While, in theory, DP-SGD can be applied to train diffusion models, it poses several significant practical challenges.
-First, it is well known that DP-SGD introduces considerable computational and memory overhead due to the need for per-sample gradient clipping, which is crucial for bounding sensitivity dwork2006calibrating ; abadi2016deep .
-Second, DP-SGD is fundamentally incompatible with batch-wise operations, such as batch normalization, since these operations inherently link samples together, rendering sensitivity analysis infeasible.
-Moreover, training a large network with DP-SGD typically leads to significant performance degradation (i.e.
-the curse of dimensionality), even under relatively lenient privacy budgets.
-To the best of our knowledge, existing diffusion models trained with DP-SGD are constrained to relatively small-scale datasets dockhorn2023differentially ; ghalebikesabi2023differentially .
-As a result, recent research has increasingly turned to alternative strategies for adapting diffusion models in a privacy-preserving manner without the need for training from scratch under DP-SGD.
-A promising approach leverages large, publicly pre-trained models, adapting them to new domains under differential privacy (DP) constraints, which allows models to capitalize on existing representational strengths while avoiding the computational and memory burdens of full DP training ghalebikesabi2023differentially .
-Similarly, parameter-efficient fine-tuning (PEFT) methods, such as DP-LoRA yu2021differentially , fine-tune only a subset of parameters, allowing for efficient, targeted adaptation with limited privacy costs.
-Methods like DP-RDM (Differentially Private Retrieval-Augmented Diffusion Model) lebensold2024dp illustrate this trend by enabling model adaptation through a retrieval mechanism that conditions image generation on private data retrieved at inference time, avoiding direct model updates.
-These approaches collectively demonstrate how large models can be adapted privately, circumventing the complexities and high costs associated with full DP training.
-Meanwhile, alternative methods such as Universal Guidance bansal2023universal and Textual Inversion (TI) gal2023an enable model adaptation to specific content or style without modifying the diffusion model itself, instead relying on external embedding vectors.
-Universal Guidance’s style guidance (SG) mechanism generates a target CLIP embedding by passing the reference image through a CLIP image encoder.
-During the reverse diffusion process, the cosine distance between this target embedding and the CLIP embedding of the generated image is minimized, guiding the generation process to match the desired style.
-TI, in contrast, learns a new token embedding representing a set of target images, which is then incorporated into the text prompt to adapt the model to the style or content associated with those images.
-Both SG and TI avoid direct model optimization, reducing memory and computation demands while also involving fewer trainable parameters, making them highly compatible with DP constraints on smaller datasets.
-In this work, we propose two novel privacy-preserving adaptation methods for smaller datasets, leveraging SG and TI to bypass the need for DP-SGD or DP-PEFT approaches that require extensive model updating.
-Our first approach adapts Universal Guidance for privacy by using a differentially private CLIP embedding.
-Although SG is typically applied to a single target image, we extend it to handle multiple images under DP by computing an embedding for each target image and aggregating these embeddings via a centroid calculation.
-To ensure differential privacy, we add calibrated noise to the centroid, which protects individual data points in the dataset while preserving the model’s guidance capabilities.
-Our second approach enhances Textual Inversion (TI) with differential privacy.
-While standard TI compresses training data into a single, low-dimensional vector, offering some intrinsic privacy benefits, it lacks the rigorous guarantees required by DP.
-We address this by introducing a private variant of TI, in which we decouple interactions among samples by learning a separate embedding for each target image.
-These embeddings are then aggregated into a noisy centroid, similar to our SG approach, ensuring differential privacy while preserving TI’s data compression advantages.
-Our experimental results underscore the practicality and effectiveness of the proposed differentially private embedding-based adaptation methods, with Textual Inversion (TI) proving particularly robust for preserving stylistic fidelity even under privacy constraints.
-In applying our approach to two distinct datasets–a private collection of artworks by @eveismyname and pictograms from the Paris 2024 Olympics Paris2024Pictograms –we demonstrate that differentially private TI can successfully capture nuanced stylistic elements while maintaining privacy guarantees.
-Our findings reveal a trade-off between privacy levels (determined by the DP parameter $\varepsilon$) and image quality, where lower $\varepsilon$ values yield diminished stylistic fidelity, yet still preserve the target style under moderate noise.
-Furthermore, subsampling offers an effective means of improving privacy resilience by controlling sensitivity to individual data points, significantly mitigating the impact of added noise on image quality.
-Ultimately, our approach provides a versatile framework for privacy-preserving model adaptation, enabling diffusion models to generalize effectively to novel styles and domains without risking exposure of sensitive data points.
-2 Background and Related Works
+In recent years, diffusion models (Ho et al., 2020; Song et al., 2021b), particularly latent diffusion models (Rombach et al., 2022), have spearheaded high quality text-to-image generation, and have been widely adopted by researchers and the general public alike.
+Trained on massive datasets like LAION-5B (Schuhmann et al., 2022), these models have developed a broad understanding of visual concepts, enabling new creative and practical applications.
+Notably, tools such as Stable Diffusion (Rombach et al., 2022; Podell et al., 2023) have been made readily accessible for general use.
+Building on this foundation, efficient adaptation methods such as parameter efficient fine-tuning (PEFT) methods (Hu et al., 2022; von Platen et al., 2022; Ruiz et al., 2023), guidance based approaches (Ho & Salimans, 2021; Kim et al., 2022; Bansal et al., 2024), and pseudo-word generation (Gal et al., 2023) enable users to leverage this extensive pretraining for customizing models that can specialize on downstream tasks with smaller datasets.
+The rapid adoption of diffusion models has raised significant privacy and legal concerns.
+These models are vulnerable to privacy attacks, such as membership inference (Duan et al., 2023), where attackers determine if a specific data point was used for training, and data extraction (Carlini et al., 2023), which enables reconstruction of training data.
+This risk is amplified during fine-tuning on smaller, domain-specific datasets, where each record has a greater impact.
+Additionally, reliance on large datasets scraped without consent raises copyright concerns (Vyas et al., 2023), as diffusion models can reproduce original artworks without credit or compensation.
+These issues highlight the urgent need for privacy-preserving technologies and clearer ethical and legal guidelines for generative models.
+Differential privacy (DP) (Dwork et al., 2006; Dwork, 2006) is a widely adopted approach to addressing these challenges, where controlled noise is added during training or inference to prevent information leakage from individual data points while still enabling the model to learn effectively from the overall dataset.
+One standard approach for ensuring DP in deep learning is Differentially Private Stochastic Gradient Descent (DP-SGD) (Abadi et al., 2016), which modifies traditional SGD by adding noise to clipped gradients.
+However, applying DP-SGD to train diffusion models poses several challenges.
+It introduces significant computational and memory overhead due to per-sample gradient clipping (Hoory et al., 2021), which is essential for bounding gradient sensitivity (Dwork et al., 2006; Abadi et al., 2016).
+DP-SGD is also incompatible with batch-wise operations like batch normalization, as these link samples and hinder sensitivity analysis.
+Furthermore, training large models with DP-SGD often leads to substantial performance degradation, particularly under realistic privacy budgets since the required noise scales with the gradient norm.
+Consequently, existing diffusion models trained with DP-SGD are limited to relatively small-scale images.
+(Dockhorn et al., 2023; Ghalebikesabi et al., 2023).
+As a result, recent research has focused on privacy-preserving strategies for adapting diffusion models without the need for full DP-SGD training.
+One approach adapts large, publicly pre-trained models to new domains under DP constraints, leveraging their representational strengths while reducing computational and memory costs (Ghalebikesabi et al., 2023).
+Similarly, PEFT methods like DP-LoRA (Yu et al., 2022) fine-tune a small subset of parameters, enabling efficient adaptation with lower privacy costs.
+Methods like DP-RDM (Lebensold et al., 2024) avoid direct model updates by using retrieval mechanisms that condition image generation on private data retrieved during inference.
+However, these alternate approaches often fail to capture the detail of style, underscoring the challenges of balancing privacy, efficiency, and generative performance.
+Independent of privacy concerns, Textual Inversion (TI) (Gal et al., 2023) provides an effective method for adapting diffusion models to specific styles or content without modifying the model.
+Instead, TI learns an external embedding vector that captures the style or content of a target image set, which is then incorporated into text prompts to guide the model’s outputs.
+A key advantage of TI is its ability to compress a style into a compact vector, reducing computational and memory demands while simplifying the application of privacy-preserving mechanisms, as privacy constraints can be applied directly to embeddings rather than the full model.
+Additionally, since TI avoids direct model optimization, it remains efficient and compatible with DP constraints on smaller datasets.
+In this work, we propose a novel privacy-preserving adaptation method for smaller datasets, leveraging TI to avoid the extensive model updates required by DP-SGD or DP-PEFT approaches.
+Standard TI inherently compresses the target dataset into a single, low-dimensional vector, providing some obfuscation benefits, but it does not offer formal privacy guarantees.
+To address this limitation, we introduce a private variant of TI, called Differentially Private Aggregation via Textual Inversion (DPAgg-TI) and summarize it in Figure 1.
+Our method decouples interactions among samples by learning a separate embedding for each target image, which are then aggregated into a noisy centroid.
+This approach ensures efficient and secure adaptation to private datasets.
+Our experiments demonstrate the effectiveness of DPAgg-TI, showing that TI remains robust in preserving stylistic fidelity even under privacy constraints.
+Applying our method to a private artwork collection by @eveismyname and Paris 2024 Olympics pictograms (Paris 2024, ), we show that DPAgg-TI captures nuanced stylistic elements while ensuring privacy.
+We observe a trade-off between privacy (controlled by DP parameter $\varepsilon$) and image quality: lower $\varepsilon$ reduces fidelity but maintains the target style under moderate noise.
+Subsampling further amplifies privacy by reducing sensitivity to individual data points, mitigating noise impact on image quality.
+This framework enables privacy-preserving adaptation of diffusion models to new styles and domains while protecting sensitive data.
+Our contributions can be summarized as follows:
+  • We propose DPAgg-TI that ensures privacy by learning separate embeddings for individual images and aggregating them into a noisy centroid.
+  • Our approach enables style adaptation without extensive model updates, reducing computational overhead while preserving privacy.
+  • We analyze the trade-off between privacy and image quality, showing that moderate noise maintains stylistic fidelity while protecting sensitive data.
+  • We validate our method on diverse datasets, demonstrating its effectiveness in capturing stylistic elements under privacy constraints.
+2 Preliminaries and Related Work
 2.1 Diffusion Models
-Diffusion models ho2020denoising ; song2020score ; song2020denoising ; rombach2022high leverage an iterative denoising process to generate high-quality images that aligns with a given conditional input from random noise.
+Diffusion models (Ho et al., 2020; Song et al., 2021b, a; Rombach et al., 2022) leverage an iterative denoising process to generate high-quality images that align with a given conditional input from random noise.
 In text-to-image generation, this conditional input is based on a textual description (a prompt) that guides the model in shaping the image to reflect the content and style specified by the text.
-To convert the text prompt into a suitable conditional format, it is first tokenized into discrete tokens, each representing a word or sub-word unit.
+To convert the text prompt into a suitable conditional format, it is first broken down into discrete tokens, each representing a word or sub-word unit.
 These tokens are then converted into a sequence of embedding vectors $v_{i}$ that encapsulate the meaning of each token within the model’s semantic space.
-Next, these embeddings pass through a transformer text encoder, such as CLIP radford2021learning , outputting a single text-conditional vector $y$ that serves as the conditioning input.
-This text-conditioning vector $y$ is then incorporated at each denoising step, guiding the model in aligning the output image with the specific details outlined in the prompt.
+Next, these embeddings pass through a transformer text encoder, such as CLIP (Radford et al., 2021), outputting a single text-conditional vector $y$ that serves as the conditioning input.
+This vector $y$ is then incorporated at each denoising step, guiding the model to align the output image with the specific details outlined in the prompt.
 The image generation process, also known as the reverse diffusion process, comprises of $T$ discrete timesteps and starts with pure Gaussian noise $x_{T}$.
@@
 The predicted noise is then used to make a reverse diffusion step from $x_{t}$ to $x_{t-1}$, iteratively refining the noisy image closer to a coherent output $x_{0}$ that aligns with the text conditional $y$.
-In particular, denoising diffusion implicit models (DDIM) sampling song2020denoising uses the predicted noise $\epsilon_{\theta}(x_{t},y,t)$ and a noise schedule represented by an array of scalars $\{\alpha_{t}\}_{t=1}^{T}$ to first predict a clean image $\hat{x}_{0}$, then makes a small step in the direction of $\hat{x}_{0}$ to obtain $x_{t-1}$.
-The reverse diffusion process for DDIM sampling can be formalized as follows:
-$\hat{x}_{0}=\frac{x_{t}-\sqrt{1-\alpha_{t}}\epsilon_{\theta}(x_{t},y,t)}{\sqrt{\alpha_{t}}}$ (1)
-$x_{t-1}=\sqrt{\alpha_{t-1}}\hat{x}_{0}+\sqrt{1-\alpha_{t-1}}\epsilon_{\theta}(x_{t},y,t)$ (2)
-The training of a text-conditioned diffusion model involves minimizing a loss function that guides the model to predict the noise added to an image, given both the noisy image $x_{t}$ and the text conditioning $y$.
-The objective function is typically a mean squared error (MSE) between the true noise $\epsilon$ and the predicted noise $\epsilon_{\theta}(x_{t},y,t)$.
+The objective function for a text-conditioned diffusion model, given both the noisy image $x_{t}$ and the text conditioning $y$, is typically a mean squared error (MSE) between the true noise $\epsilon$ and the predicted noise $\epsilon_{\theta}(x_{t},y,t)$.
 The denoising model is therefore trained over the following optimization problem:
-$\theta^{*}=\arg\min_{\theta}\mathbb{E}_{x,\epsilon\sim\mathcal{N}(0,I),t\sim[T]}\left[\|\epsilon-\epsilon_{\theta}(x_{t},y,t)\|^{2}\right]$ (3)
-2.1.1 Style Guidance
-Bansal et al.
-bansal2023universal propose a universal guidance algorithm that allows the reverse diffusion process to be guided by arbitrary guidance functions, including object detection, image segmentation and facial recognition.
-In this section, we formally describe CLIP style guidance, one of the many processes explored in the universal guidance study.
-We also only describe the forward guidance process.
-For a generalized and complete description of universal guidance process, including the backward guidance process and per-step self-recurrence, we defer the reader to the original paper.
-Let $x_{c}$ denote a target image whose style we wish to replicate and $\textsf{CLIP}(\cdot)$ denote a CLIP image encoder.
-To incorporate style guidance, given a noisy image $x_{t}$, we first compute $\hat{x}_{0}$ using the predicted noise $\epsilon_{\theta}(x_{t},y,t)$ and Equation (1).
-Using $\hat{x}_{0}$, we then compute the following:
-$\hat{\epsilon}_{\theta}(x_{t},y,t)=\epsilon_{\theta}(x_{t},y,t)+w\sqrt{1-\alpha_{t}}\nabla_{x_{t}}\ell_{\cos}(\textsf{CLIP}(x_{c}),\textsf{CLIP}(\hat{x}_{0}))$ (4)
-where $w$ is a guidance weight parameter and $\ell_{\cos}(\cdot,\cdot)$ is the negative cosine loss.
-We then perform the reverse diffusion step using the same procedure given in Equations (1) and (2), but replacing $\epsilon_{\theta}(x_{t},y,t)$ with $\hat{\epsilon}_{\theta}(x_{t},y,t)$.
-Using this reverse diffusion procedure, we generate a high-quality image $x_{0}$ that aligns with the text conditioning $y$ while also optimizing the cosine similarity between the CLIP embeddings of $x_{c}$ and $x_{0}$, resulting in $x_{0}$ having the stylistic characteristics of $x_{c}$.
-2.1.2 Textual Inversion
-Textual Inversion (TI) gal2023an is an adaptation technique that enables personalization using a small dataset of typically only 3-5 images.
+$\theta^{*}=\arg\min_{\theta}\mathbb{E}_{x,\epsilon\sim\mathcal{N}(0,I),t\sim[T]}[\|\epsilon-\epsilon_{\theta}(x_{t},y,t)\|^{2}].$ (1)
+Textual Inversion.
+Textual Inversion (TI) (Gal et al., 2023) is an adaptation technique that enables personalization using a small dataset of typically 3-5 images.
 This approach essentially learns a new token that encapsulates the semantic meaning of the training images, allowing the model to associate specific visual features with a custom token.
-To achieve this, TI trains a new token embedding, denoted as $u$, which represents a placeholder token, often denoted as $S$.
-During training, images are conditioned of phrases such as “A photo of $S$” or “A painting in the style of $S$”.
+To achieve this, TI trains a new token embedding, denoted as $u$, representing a placeholder token, denoted as $S$.
+During training, images are conditioned on phrases such as “A photo of $S$” or “A painting in the style of $S$”.
 However, unlike the fixed embeddings of typical tokens $v_{i}$, $u$ is a learnable parameter.
 Let $y_{u}$ denote the text conditioning vector resulting from a prompt containing the token $S$.
-Through gradient descent, TI minimizes the diffusion model loss with respect to $u$, while keeping the diffusion model parameters $\theta$ fixed, iteratively refining this embedding to capture the unique characteristics of the training images.
+Through gradient descent, TI minimizes the diffusion model loss given in (1) with respect to $u$, while keeping the diffusion model parameters $\theta$ fixed, iteratively refining this embedding to capture the unique characteristics of the training images.
 The resulting optimal embedding $u_{*}$ is formalized as follows:
@@
 2.2 Differential Privacy
-In this work, we adopt differential privacy (DP) dwork2006calibrating ; dwork2006differential as our privacy framework.
+In this work, we adopt differential privacy (DP) (Dwork et al., 2006; Dwork, 2006) as our privacy framework.
 Over the past decade, DP has become the gold standard for privacy protection in both research and industry.
-It measures the stability of a randomized algorithm concerning changes in an input instance, thereby quantifying the extent to which an adversary can infer the existence of a specific input based on the algorithm’s output.
-Definition 2.1 ((Approximate) Differential Privacy)
+It measures the stability of a randomized algorithm with respect to changes in an input instance, thereby quantifying the extent to which an adversary can infer the existence of a specific input based on the algorithm’s output.
+Definition 2.1 ((Approximate) Differential Privacy).
 For $\varepsilon,\delta\geq 0$, a randomized mechanism $\mathcal{M}:\mathcal{X}^{n}\rightarrow\mathcal{Y}$ satisfies $(\varepsilon,\delta)$-DP if for all neighboring datasets $\mathcal{D},\mathcal{D}^{\prime}\in\mathcal{X}^{n}$ which differ in a single record (i.e., $\left\lVert\mathcal{D}-\mathcal{D}^{\prime}\right\rVert_{\mathsf{H}}\leq 1$ where $\left\lVert\cdot\right\rVert_{\mathsf{H}}$ is the Hamming distance) and all measurable $\mathcal{S}$ in the range of $\mathcal{M}$, we have that
@@
 When $\delta=0$, we say $\mathcal{M}$ satisfies $\varepsilon$-pure DP or ($\varepsilon$-DP).
-To achieve DP, the Gaussian mechanism is often applied dwork2014algorithmic ; balle2018improving , adding Gaussian noise scaled by the sensitivity of the function $f$ and privacy parameters $\varepsilon$ and $\delta$.
-Specifically, noise with standard deviation $\sigma=\frac{\Delta_{f}\sqrt{2\ln(1.25/\delta)}}{\varepsilon}$ is added to the output [^1] In practice, we use numerical privacy accountant such as balle2018improving ; mironov2017renyi to calibrate the noise.
-balle2018improving , where $\Delta_{f}$ represents $\ell_{2}$-sensitivity of the target function $f(\cdot)$.
+To achieve DP, the Gaussian mechanism is often applied (Dwork et al., 2014; Balle & Wang, 2018), adding Gaussian noise scaled by the sensitivity of the function $f$ and privacy parameters $\varepsilon$ and $\delta$.
+Specifically, noise with standard deviation $\sigma=\frac{\Delta_{f}\sqrt{2\ln(1.25/\delta)}}{\varepsilon}$ is added to the output [^1] In practice, we use numerical privacy accountant such as (Balle & Wang, 2018; Mironov, 2017) to calibrate the noise.
+(Balle & Wang, 2018), where $\Delta_{f}$ represents $\ell_{2}$-sensitivity of the target function $f(\cdot)$.
 When the context is clear, we may omit the subscript $f$.
-This mechanism enables a smooth privacy-utility tradeoff and is widely used in privacy-preserving machine learning, including in DP-SGD abadi2016deep , which applies Gaussian noise during model updates to achieve DP.
-2.2.1 Privacy Amplification by Subsampling
-Subsampling is a standard technique in DP, where a full dataset of size $n$ is first subsampled to $m$ records without replacement (typically with $m\ll n$), and the privatization mechanism (such as the Gaussian mechanism) is then applied.
+This mechanism enables a smooth privacy-utility tradeoff and is widely used in privacy-preserving machine learning, including in DP-SGD (Abadi et al., 2016), which applies Gaussian noise during model updates to achieve DP.
+Privacy Amplification by Subsampling.
+Subsampling is a standard technique in DP, where a full dataset of size $n$ is first subsampled to $m$ records without replacement (typically with $m\ll n$) before the privatization mechanism (such as the Gaussian mechanism) is applied.
 Specifically, if a mechanism provides $(\varepsilon,\delta)$-DP on a dataset of size $m$, it achieves $(\varepsilon^{\prime},\delta^{\prime})$-DP on the subsampled dataset, where $\delta^{\prime}=\frac{m}{n}\delta$ and
 $\varepsilon^{\prime}=\log\left(1+\frac{m}{n}\left(e^{\varepsilon}-1\right)\right)=O\left(\frac{m}{n}\varepsilon\right).$ (3)
-This result is well-known (see, for example, (steinke2022composition, , Theorem 29)). Notably, if the DP mechanism is Gaussian, even tighter amplification bounds can be applied mironov2017renyi .
+This result is well-known (Steinke (2022, Theorem 29)), with tighter amplification bounds available for Gaussian mechanisms (Mironov, 2017).
 2.3 Differentially Private Adaptation of Diffusion Models
-Recent advancements in applying differential privacy (DP) to diffusion models have aimed to balance privacy preservation with the high utility of generative outputs.
+Recent advancements in applying DP to diffusion models have aimed to balance privacy preservation with the high utility of generative outputs.
 Dockhorn et al.
-dockhorn2023differentially proposed a differentially private diffusion model (DP diffusion) that enables privacy-preserving generation of realistic samples, setting a foundational approach for adapting diffusion processes under DP constraints.
-Another common strategy involves training a model on a large public dataset, followed by differentially private finetuning on a private dataset, as explored by Ghalebikesabi et al.
-ghalebikesabi2023differentially .
-While effective in certain contexts, this approach raises privacy concerns, particularly around risks of information leakage during the finetuning phase pmlr-v235-tramer24a .
+(Dockhorn et al., 2023) proposed a Differentially Private Diffusion Model (DPDM) that enables privacy-preserving generation of realistic samples, setting a foundational approach for adapting diffusion processes using DP-SGD.
+Another common strategy involves training a model on a large public dataset, followed by differentially private fine-tuning on a private dataset, as explored by Ghalebikesabi et al.
+(2023).
+While effective in certain contexts, this approach raises privacy concerns, particularly around risks of information leakage during the fine-tuning phase (Tramèr et al., 2024).
 In response to these limitations, various adaptation techniques have emerged.
-Although not specific to diffusion models, some methods focus on training models on synthetic data followed by DP-constrained finetuning, as in the VIP approach yu2024vip , which demonstrates the feasibility of applying DP in later adaptation stages.
-Other approaches explore differentially private learning of feature representations sander2024differentially , aiming to distill private information into a generalized embedding space while maintaining DP guarantees.
+Although not specific to diffusion models, some methods focus on training models on synthetic data followed by DP-constrained fine-tuning, as in the VIP approach (Yu et al., 2024), which demonstrates the feasibility of applying DP in later adaptation stages.
+Other approaches explore differentially private learning of feature representations (Sander et al., 2024), aiming to distill private information into a generalized embedding space while maintaining DP guarantees.
 Although these adaptations are not yet implemented for diffusion models, they lay essential groundwork for developing secure and efficient privacy-preserving generative models.
-3 Differentially Private Adaptation via Noisy Aggregated Embeddings
-3.1 General Approach
+Figure 2: Samples of images used in our style adaptation experiments. Left: artwork by @eveismyname ($n=158$). Right: Paris 2024 Olympic pictograms ($n=47$), © International Olympic Committee, 2023.
+3 Differentially Private Adaptation via Textual Inversion
+TI is inherently parameter-efficient and offers certain privacy benefits, as information from an entire dataset of images is compressed into a single token embedding vector.
+This compression limits the model’s capacity to memorize specific images, making data extraction attacks difficult.
+However, this privacy is merely heuristic and yet to be proven, so TI may still be vulnerable to privacy attacks such as membership inference.
+A similar adaptation technique with privacy guarantees may therefore be desirable.
 Let $x^{(1)},\dots,x^{(n)}$ represent a target dataset of images whose characteristics we wish to privately adapt our image generation towards.
-Assuming that we can encode each image $x^{(i)}$ as a directional embedding vector $u^{(i)}$, we can aggregate the embeddings $u^{(1)},\dots,u^{(n)}$ by calculating the centroid.
+Instead of training a single token embedding on the entire dataset as in regular TI, we train a separate embedding $u^{(i)}$ on each $x^{(i)}$ to obtain a set of embeddings $u^{(1)},\dots,u^{(n)}$, as illustrated in Figure 1.
+We can formalize the encoding process as follows:
+$u^{(i)}=\arg\min_{u}\mathbb{E}_{\epsilon\sim\mathcal{N}(0,I),t}[\|\epsilon-\epsilon_{\theta}(x_{t}^{(i)},y_{u},t)\|^{2}].$ (4)
+Then, we can aggregate the embeddings $u^{(1)},\dots,u^{(n)}$ by calculating the centroid.
 The purpose of this aggregation is to limit the sensitivity of the final output to each $x^{(i)}$.
@@
 We can therefore define the resulting embedding vector $u^{*}$ as follows:
-$u^{*}=\frac{1}{n}\sum_{i=1}^{n}u^{(i)}+\mathcal{N}(0,\sigma^{2}I)$ (7)
-where the minimum $\sigma$ required to provide $(\varepsilon,\delta)$-DP is given by the following expression based on (balle2018improving, , Theorem 1):
-$\sigma=\frac{\Delta}{n}\cdot\frac{\sqrt{2\log(1.25)/\delta}}{\varepsilon}.$ (8)
+$u^{*}=\frac{1}{n}\sum_{i=1}^{n}u^{(i)}+\mathcal{N}(0,\sigma^{2}I),$ (5)
+where the minimum $\sigma$ required to provide $(\varepsilon,\delta)$-DP is given by the following expression based on Balle & Wang (2018, Theorem 1):
+$\sigma=\frac{\Delta}{n}\cdot\frac{\sqrt{2\ln(1.25/\delta)}}{\varepsilon}.$ (6)
 In the context of our problem, $\Delta=\sup_{i,j}\|u^{(i)}-u^{(j)}\|$. Since our embedding vectors are directional, we can normalize each $u^{(i)}$, allowing us to set $\Delta=2$.
 The noisy centroid embedding $u^{*}$ can then be used to adapt the downstream image generation process.
-In Sections 3.3 and 3.4, we describe how to apply our method to TI and SG respectively, outlining the encoding of $x^{(i)}$ as $u^{(i)}$ and the role of $u^{*}$ in the image generation process.
+Similar to regular TI’s $u_{*}$, we can use $u^{*}$ to represent a new placeholder token $S^{*}$ that can be incorporated into prompts for personalized image generation.
+While $u^{*}$ may not fully solve the TI optimization problem presented in (2), it provides provable privacy guarantees, with only a minimal trade-off in accurately representing the style of the target dataset.
 To reduce the amount of noise needed to provide the same level of DP, we employ subsampling: instead of computing the centroid over all $n$ embedding vectors, we randomly sample $m\leq n$ embedding vectors without replacement and compute the centroid over only the sampled vectors.
@@
 where $\sigma$ can be computed numerically for any target $\varepsilon,\delta$ and subsampling rate $\frac{m}{n}$.
-3.2 Using Style Guidance
-Applying our method to style guidance is straightforward.
-To obtain the embeddings, we simply use the CLIP image encoder to obtain the embedding for each image.
-More formally, for each $x^{(i)}$, let $u^{(i)}=\textsf{CLIP}(x^{(i)})$.
-We can then aggregate the embeddings $u^{(i)}$ to obtain $u^{*}$ via Equation (7) or (9), depending on whether we wish to incorporate subsampling.
-Finally, we use $u^{*}$ to guide the reverse diffusion process, as outlined in Section 2.1.1, by computing $\hat{\epsilon}_{\theta}(x_{t},y,t)$ via a modification of Equation (2):
-$\hat{\epsilon}_{\theta}(x_{t},y,t)=\epsilon_{\theta}(x_{t},y,t)+w\sqrt{1-\alpha_{t}}\nabla_{x_{t}}\ell_{\cos}(u^{*},\textsf{CLIP}(\hat{x}_{0}))$ (10)
-3.3 Using Textual Inversion
-Textual Inversion (TI) is inherently parameter-efficient and offers certain privacy benefits, as information from an entire dataset of images is compressed into a single token embedding vector.
-This compression limits the model’s capacity to memorize specific images, making data extraction attacks difficult.
-However, this privacy is merely heuristic and yet to be proven, so TI may still be vulnerable to privacy attacks such as membership inference.
-A similar adapation technique with privacy guarantees may therefore be desirable.
-To incorporate our approach into TI, instead of training a single token embedding on the entire dataset, we instead train a separate embedding $u^{(i)}$ on each $x^{(i)}$ to obtain a set of embeddings $u^{(1)},\dots,u^{(n)}$.
-We can formalize the encoding process as follows:
-$u^{(i)}=\arg\min_{u}\mathbb{E}_{\epsilon\sim\mathcal{N}(0,I),t}[\|\epsilon-\epsilon_{\theta}(x_{t}^{(i)},y_{u},t)\|^{2}]$ (11)
-We can then aggregate the embeddings $u^{(i)}$ to obtain $u^{*}$ via Equation (7) or (9), depending on whether we wish to utilize subsampling.
-Similar to TI’s $u_{*}$, we can use $u^{*}$ to represent a new placeholder token $S^{*}$ that can be incorporated into prompts for personalized image generation.
-While $u^{*}$ may not fully solve the TI optimization problem presented in Equation (5), it provides provable privacy guarantees, with only a minimal trade-off in accurately representing the content or style of the training data.
 4 Experimental Results
 4.1 Datasets
-We compiled two datasets for our style adaptation experiments, designed to proxy for private datasets unrecognized by the base Stable Diffusion model.
-The first dataset consists of 158 artworks created by the artist @eveismyname, who has provided consent for non-commercial use.
-This dataset serves to examine how models can capture and replicate artistic styles without directly copying individual works.
-Although we acknowledge the possibility that a few of these artworks may have been inadvertently included in the pre-training of Stable Diffusion due their public presence on social media platforms, the artist’s limited recognition and modest portfolio size make it unlikely that their unique style could be effectively extracted from the pre-trained model.
-The second dataset consists of 47 pictograms from the Paris 2024 Olympics Paris2024Pictograms , approved strictly for non-commercial editorial use IOCProperties .
-These pictograms were unveiled in February 2023, several months after the release of Stable Diffusion v1.5, which serves as the base model for our experiments.
-Through this dataset, we hope to evaluate how effectively our approach adapts to novel and unfamiliar visual patterns.
-Representative samples from both datasets are presented in Figure 1.
-Figure 1: Samples of images used in our style adaptation experiments. Left: artwork by @eveismyname ($n=158$). Right: Paris 2024 Olympic pictograms ($n=47$), © International Olympic Committee, 2023.
-4.2 Style Guidance Results
-We applied our SG-based bansal2023universal approach, detailed in Section 3.2 to both datasets.
-While this method successfully provided privacy protection by obfuscating embedding details, the resulting image quality and style capture was notably less compared to the TI-based approach.
-The CLIP embeddings, even without noise or subsampling, led to images that retained only generalized stylistic elements, lacking the detailed fidelity and coherence achieved with TI.
-Results of our attempt can be seen in Figure 2.
-This result underscores the superiority of our TI method for maintaining both privacy and high-quality image generation.
-A possible explanation for the difference in style transfer effectiveness is that SG, as with other guidance methods, suffers from instability and sensitive dependency to hyperparameters such as guidance weight $w$.
-While Bansal et al.
-proposed using backward guidance and per-step self-recurrence to remedy these issues, we found that they were insufficient for our application.
-Another possible explanation is that the stylistic signal of the image embedding vectors $\textsf{CLIP}(x^{(i)})$ are simply not resilient to our aggregation process.
-Figure 2: Attempts of using universal guidance to generate drawings of Taylor Swift and icons of the Eiffel Tower in the styles of @eveismyname and Paris 2024 Pictograms respectively.
-Here, we apply no subsampling or DP-noise.
-4.3 Textual Inversion Results
+We compiled two datasets to evaluate our style adaptation method, specifically selecting content unlikely to be recognized by Stable Diffusion v1.5, our base model.
+The first dataset consists of 158 artworks by the artist @eveismyname, who has granted consent for non-commercial use.
+This dataset allows us to assess whether models can capture artistic styles without memorizing individual works.
+While some of these artworks may have been publicly accessible on social media, making incidental inclusion in Stable Diffusion’s pretraining possible, the artist’s limited recognition and relatively small portfolio reduce the likelihood that the model has internalized her unique style.
+This dataset serves as a controlled test for privacy-preserving style transfer on individual artistic collections.
+The second dataset contains 47 pictograms from the Paris 2024 Olympics (Paris 2024, ), permitted strictly for non-commercial editorial use (International Olympic Committee, ).
+These pictograms were officially released in February 2023, several months after the release of Stable Diffusion v1.5, ensuring they were absent from the model’s pretraining data.
+This dataset allows us to assess how well our approach adapts to newly introduced visual styles that the base model has never encountered.
+Both datasets are used to test the ability of our method to extract and transfer stylistic elements while preserving privacy. Representative samples are shown in Figure 2.
+4.2 Style Transfer Results
 Figure 3: Images generated by Stable Diffusion v1.5 using the prompt “A painting of Taylor Swift in the style of $<$@eveismyname$>$”, with the embedding $<$@eveismyname$>$ trained using different values of $m$ and $\varepsilon$.
-Figure 4: Images generated by Stable Diffusion v1.5 using the prompt “Icon of the Eiffel Tower in the style of <Paris 2024 Pictograms>”, with the embedding <Paris 2024 Pictograms> trained using different values of $m$ and $\varepsilon$.
-Using both the @eveismyname and Paris 2024 pictograms dataset, we trained TI gal2023an embeddings on Stable Diffusion v1.5 rombach2022high using our proposed method detailed in Section 3.3.
-Our primary goal is to investigate how DP configurations, specifically the $\varepsilon$ and subsampling sizes ($m$), affect the generated images’ quality and privacy resilience.
-For regular TI, we utilize Stable Diffusion’s default process to embed the private dataset without any additional noise.
-For the DP-enhanced experiments using our method, we test multiple configurations of $m$ and $\varepsilon$ to analyze the trade-off between image fidelity and privacy resilience.
-Figures 3 and  4 present generated images across three key configurations: (1) regular TI without DP, (2) centroid TI with DP at various epsilon values, and (3) centroid TI with DP at different subsampling sizes within the private dataset.
-We used the same random seed to generate embeddings and sample the images for ease of visual comparision between different configurations.
+Figure 4: Images generated by Stable Diffusion v1.5 using the prompt “Icon of a dragon in the style of $<$Paris 2024 Pictograms$>$”, with the embedding $<$Paris 2024 Pictograms$>$ trained using different values of $m$ and $\varepsilon$.
+Using both the @eveismyname and Paris 2024 pictograms dataset, we trained TI (Gal et al., 2023) embeddings on Stable Diffusion v1.5 (Rombach et al., 2022) using DPAgg-TI.
+Our primary goal is to investigate how DP configurations, specifically the privacy budget $\varepsilon$ and subsampling size $m$, affect the generated images quality and privacy resilience.
+For regular TI, we utilize the default process to embed the private dataset without any additional noise.
+For the DPAgg-TI, we test multiple configurations of $m$ and $\varepsilon$ to analyze the trade-off between image fidelity and privacy.
+Figures 3 and  4 present generated images across two key configurations: (1) regular TI without DP, (2) DPAgg-TI with DP at different values of $m$ and $\varepsilon$.
+We used the same random seed to generate embeddings, subsample images, and sample DP noise for ease of visual comparision between different configurations.
 As with common practice, we set $\delta=1/n$.
 Since $\sigma$ is undefined for $\varepsilon=0$, we demonstrate the results of $\varepsilon\approx 0$, in other words, infinite noise, by setting $\varepsilon=10^{-5}$.
+The purpose of this parameter value is to demonstrate the image generated when $u^{*}$ contains zero information about the target dataset.
 Images generated without DP closely resemble the unique stylistic elements of the target dataset.
@@
 In particular, $y_{u^{*}}$ tends towards $y$ (a conditioning vector independent of the learnable embedding).
-In our results, this manifests as a painting of Taylor Swift devoid of the artist-specific stylistic elements, or a generic icon of the Eiffel Tower (with color, as opposed to the black and white design of the pictograms).
+In our results, this manifests as a painting of Taylor Swift devoid of the artist-specific stylistic elements, or a generic icon of a dragon (with color, as opposed to the black and white design of the pictograms).
 With this in mind, $\varepsilon$ can be interpreted as a drift parameter, representing the progression from the optimal $u^{*}$ towards infinity, gradually steering the generated image away from the target style in exchange for stronger privacy guarantees.
+We also observe instances where there is a temporary drop in prompt fidelity (e.g., $m=16,\varepsilon\in[0.5,1]$ in Figure 3 and intermediate $\varepsilon$ values in Figure 4) which restores as $u^{*}$ drifts even further from its optimal value.
+We hypothesize that this is due to drifted $u^{*}$ capturing a different meaning unrelated to the prompt, before losing any meaning that could be interpreted by Stable Diffusion’s text encoder, causing $u^{*}$ to be disregarded from $y_{u^{*}}$ and the prompt fidelity to be restored.
+Another possible explanation is that the temporary drop in prompt fidelity is due to the drift path of $u^{*}$ passing through non-linear regions within embedding space.
+We leave further investigations into this observation for future work.
 Meanwhile, reducing $m$ also reduces the sensitivity of the generated image to $\varepsilon$, as evident by the observation that, on both datasets at $m=4$, (subsampling rate below $0.1$) image generation can tolerate $\varepsilon$ as low as $0.5$ without significant changes in visual characteristics, and retaining stylistic elements of the target dataset at $\varepsilon$ as low as $0.1$.
@@
 Mathematically, the subsample centroid is an unbiased estimate of the true centroid, and so the subsampling process itself defines a distribution centered at the true centroid.
-However, the amount of noise introduced by the subsampling process is limited by the individual image embeddings, as a subsample centroid can only stray from the true centroid as much as the most anomalous point in the dataset.
+However, the amount of noise introduced by the subsampling process is limited by the individual image embeddings, as a subsample centroid can only stray from the true centroid as much as the biggest outlier in the dataset.
+4.3 Quantitative Evaluation
+4.3.1 User Study
+To evaluate the utility of our approach under different DP and subsampling configurations, we conducted a user study with 25 participants.
+Each participant was shown reference images from the target dataset and asked to compare pairs of generated images, selecting the one that better captured the style of the reference images.
+Images were generated using 10 prompts and adapted TI embeddings for the @eveismyname and Paris 2024 Pictogram datasets, resulting in 20 groups of images.
+Each participant evaluated two groups, one randomly selected from each dataset, with comparisons focusing on model configurations differing by DP noise and subsampling size.
+Survey results, summarized in Table 3 in Appendix A, align with our design goals.
+Participants showed no clear preference between regular TI and DPAgg-TI, suggesting that our privacy-preserving approach maintains perceptual quality.
+As expected, both DP noise and reduced subsampling size degraded style fidelity, consistent with the trade-offs inherent in differential privacy.
+Preferences at $\varepsilon=1$ were split, but subsampling was generally favored, reinforcing its role in reducing noise impact while preserving style.
+4.3.2 Kernel Inception Distance
+The Kernel Inception Distance (KID) (Bińkowski et al., 2018) is a metric for evaluating generative models by measuring the difference between the distributions of generated and training images in an embedding space.
+To compute KID, images generated by the model and real training images are passed through an Inception network (Szegedy et al., 2015), and their distributional differences are estimated.
+Unlike the more commonly used Fréchet Inception Distance (FID) (Heusel et al., 2017), KID is an unbiased estimator of the true divergence between the learned and target distributions (Jayasumana et al., 2024), making it more suitable for smaller datasets, as in our case.
+We report KID scores for different parameters in Tables 2 and 2, showing that DPAgg-TI maintains the style transfer fidelity of TI while ensuring differential privacy.
+Further discussion of these results is provided in Appendix B.
+Table 1: KID scores of DPAgg-TI on @eveismyname dataset for various $\varepsilon$ values ranging from $\varepsilon=10^{-5},0.1,0.5,1.0,2.0,5.0$ (including no DP) under different subsampling levels ($m=4,8,16,32$) as well as regular TI (ctrl).
+$m$ No DP $\varepsilon=$ 5.0 $\varepsilon=$ 2.0 $\varepsilon=$ 1.0 $\varepsilon=$ 0.5 $\varepsilon=$ 0.1 $\varepsilon\approx$ 0
+0.0444 0.0794 0.0422 0.0529 0.0690 0.1117 0.0654
+32 0.0752 0.0845 0.0865 0.1167 0.0300 0.0649 0.0657
+16 0.0351 0.0379 0.0430 0.0663 0.1309 0.0438 0.0658
+8 0.0359 0.0366 0.0350 0.0366 0.0396 0.0530 0.0658
+4 0.0245 0.0250 0.0249 0.0250 0.0258 0.0314 0.0653
+ctrl 0.0318
+Table 2: KID scores of DPAgg-TI on Paris dataset for various $\varepsilon$ values ranging from $\varepsilon=1e-5,0.1,0.5,1.0,2.0,5.0$ (including no DP) under different subsampling levels ($m=4,8,16,32$) as well as regular TI (ctrl).
+$m$ No DP $\varepsilon=$ 5.0 $\varepsilon=$ 2.0 $\varepsilon=$ 1.0 $\varepsilon=$ 0.5 $\varepsilon=$ 0.1 $\varepsilon\approx$ 0
+0.1146 0.1202 0.1368 0.1314 0.1389 0.1209 0.1274
+32 0.1220 0.1036 0.1258 0.1377 0.1307 0.1245 0.1259
+16 0.1311 0.1424 0.1170 0.1311 0.1381 0.1335 0.1278
+8 0.1317 0.1307 0.1220 0.1117 0.1295 0.1313 0.1272
+4 0.1141 0.1094 0.1137 0.1190 0.1194 0.1583 0.1259
+ctrl 0.1388
+4.4 Ablation Study
+4.4.1 Textual Inversion with DP-SGD
+Figure 5: Comparing our approach to applying DP-SGD to regular TI using prompt “a painting of Taylor Swift in the style of @eveismyname”.
+A natural question that arises is how well our approach compares to the naive method of applying DP-SGD to regular TI training.
+We therefore integrated DP-SGD into the TI codebase using the Opacus library and trained similar embeddings on the @eveismyname and Paris 2024 datasets.
+We found that in most cases, notably the @eveismyname dataset, the amount of noise required for DP-SGD to achieve a reasonable value of $\varepsilon$ for DP is so high that the resulting embedding contains negligible information about the training dataset.
+In particular, the results for $\varepsilon=1$ are almost indistinguishable to $\varepsilon\approx 0$, as shown in Figure 5.
+We believe that this is simply because DP-SGD is not designed to handle such small datasets in the order of 100 images.
+Additional results can be found in Appendix D.
+4.4.2 Differentially Private Adaptation Using Style Guidance
+Figure 6: Attempts of using universal guidance to generate drawings of Taylor Swift and icons of the Eiffel Tower in the styles of @eveismyname and Paris 2024 Pictograms respectively.
+Here, we apply no subsampling or DP-noise.
+We extend our approach to style guidance (SG) by leveraging the framework of Universal Guidance (Bansal et al., 2024).
+Specifically, we focus on CLIP-based style guidance, which optimizes the similarity between the CLIP embeddings of a target image and the generated image.
+We encode each target image $x^{(i)}$ as $u^{(i)}$ via a CLIP image encoder, then aggregate the embeddings $u^{(1)},\dots,u^{(n)}$ into $u^{*}$ using (5) or (7), depending on whether subsampling is applied.
+The aggregated embedding $u^{*}$ is then incorporated into the reverse diffusion process as a style guide.
+Further implementation detailes are provided in Appendix C.
+We apply our SG-based approach to both datasets.
+While it provides privacy protection by obfuscating embedding details, the resulting images captured only generalized stylistic elements and lack the detailed fidelity and coherence achieved with the TI-based method.
+As shown in Figure 6, this highlights the superiority of TI in balancing privacy and high-quality image generation.
+The reduced effectiveness of SG for style transfer may stem from its sensitivity to hyperparameters such as the guidance weight $w$, leading to instability.
+Although Bansal et al.
+(2024) proposed remedies, namely backward guidance and per-step self-recurrence, these proved insufficient for our application.
+Additionally, the CLIP embeddings may not retain enough stylistic detail after the aggregation.
 5 Conclusion
-We presented a differentially private adaptation approach for diffusion models using embedding-based methods–Universal Guidance and Textual Inversion (TI)–to enable privacy-preserving style transfer from small, sensitive datasets.
-Through experiments on a private artwork dataset and the Paris 2024 Olympics pictograms, we demonstrated that our approach maintains high stylistic fidelity while providing strong privacy guarantees.
-Notably, TI proved more effective in capturing detailed stylistic elements compared to SG, especially with added noise and subsampling to protect privacy.
-Our findings suggest that embedding-driven methods can serve as efficient and scalable alternatives to DP-SGD for privacy-preserving model adaptation, minimizing computational demands and offering resilience against privacy attacks.
-This approach provides a balance between style quality and privacy, enabling generative models to adapt responsibly to new domains.
-6 Acknowledgments
-We sincerely thank Tatchamon Wongworakul (@eveismyname) for graciously providing her artwork for use in this study.
+We presented a differentially private adaptation method for diffusion models using Textual Inversion for privacy-preserving style transfer.
+Experiments on private artwork and Paris 2024 pictograms showed TI preserves stylistic fidelity and outperforms Style Guidance.
+Our results demonstrate embedding-driven methods as efficient, scalable alternatives to DP-SGD, balancing style quality and privacy.
+Impact Statement
+The use of images without owner consent raises significant ethical concerns, particularly regarding the exploitation of intellectual property.
+This work introduces a method for visual generative models to adapt to new styles and classes while ensuring privacy and copyright protection for data owners.
+By providing a framework for privacy-preserving adaptation, this technology aims to respect intellectual property and address ethical challenges in generative AI.
+While it does not eliminate the need for consent from data owners, we hope that it represents a step toward balancing innovation with ethical considerations in AI development.
+Beyond creative applications, the proposed method has broader potential uses, including synthetic data generation, privacy-preserving personalization, and fine-tuning diffusion models for private or domain-specific tasks.
+Acknowledgements
+We sincerely thank Tatchamon Wongworakul (@eveismyname) for providing her artwork for use in this study.
+We are also grateful to Anwar Hithnawi and Varun Chandrasekaran for their insightful discussions and feedback, as well as to all participants in our user study.
+Sanmi Koyejo acknowledges support by NSF 2046795 and 2205329, IES R305C240046, the MacArthur Foundation, Stanford HAI, OpenAI, and Google.
+Appendix A User Study
+A.1 Study Design and Objective
+The user study aimed to assess the utility of our approach under different DP and subsampling configurations by evaluating the models’ ability to adapt to novel styles.
+The study involved 25 participants, each of whom was tasked with comparing images generated using various configurations and selecting the one that better captured the style of reference images.
+A.2 Experimental Setup
+Participants were shown reference images from two datasets:
+  • The @eveismyname dataset of private artwork.
+  • The Paris 2024 Pictogram dataset.
+For each dataset, 10 prompts were used to generate images, resulting in 20 groups of images (10 prompts per dataset).
+Each group included images generated using the same prompt and dataset but with different model configurations.
+Configurations varied in the addition of DP noise and the size of subsampling.
+  • Original Textual Inversion (TI)
+  • DPAgg-TI ($\varepsilon=\infty$, no DP) w/o subsampling
+  • DPAgg-TI ($\varepsilon=1$) without subsampling
+  • No Adaptation
+  • DPAgg-TI ($\varepsilon=\infty$, no DP) with subsampling $(m=8)$
+  • DPAgg-TI ($\varepsilon=1$) with subsampling $(m=8)$
+  • Style Guidance (SG)
+A.3 Survey Procedure
+Participants were asked to evaluate two groups of images: one randomly selected from the @eveismyname dataset and one from the Paris 2024 Pictogram dataset. For each group:
+  1. 1.
+    Participants were shown reference images from the target dataset.
+  2. 2.
+    They were presented with pairs of images generated using different model configurations for the same prompt.
+  3. 3.
+    Participants selected the image they felt better captured the style of the reference images.
+A.4 Evaluation Metrics
+The study focused on assessing:
+  • Participants’ preference between regular TI and DPAgg-TI for style adaptation.
+  • The impact of DP noise and subsampling size on the perceived utility of style transfer.
+A.5 Results and Analysis
+The results are summarized in Table 3. Key observations include:
+  • Participants showed no clear preference between regular TI and DPAgg-TI in capturing styles for either dataset.
+  • Both DP noise and reduced subsampling size decreased the perceived quality of style transfer.
+  • Preferences were split between configurations with $\varepsilon=1$ with and without subsampling, though subsampling generally had favorable outcomes.
+These findings highlight the trade-off between increased DP robustness and reduced utility, suggesting that the optimal configuration may depend on subjective preferences and specific application requirements.
+Table 3: Survey Results.
+regular TI No Adaptation Unsure
+@eveismyname 19 4 2
+Paris 2024 16 6 3
+DPAgg-TI (no DP, no subsampling) No Adaptation Unsure
+@eveismyname 16 9 0
+Paris 2024 15 4 6
+regular TI DPAgg-TI (no DP, no subsamp.) Unsure
+@eveismyname 12 13 0
+Paris 2024 9 10 6
+regular TI DPAgg-TI (no DP, subsamp. $m=8$) Unsure
+@eveismyname 16 6 3
+Paris 2024 7 13 5
+DPAgg-TI (no DP, no subsampling) DPAgg-TI (no DP, subsamp. $m=8$) Unsure
+@eveismyname 18 4 3
+Paris 2024 10 8 7
+DPAgg-TI ($\varepsilon=1$) no subsampling DPAgg-TI ($\varepsilon=1$, subsamp. $m=8$) Unsure
+@eveismyname 14 10 1
+Paris 2024 3 16 6
+DPAgg-TI (no DP, no subsampling) Style Guidance Unsure
+@eveismyname 16 8 1
+Paris 2024 20 2 3
+DPAgg-TI ($\varepsilon=1$, subsamp. $m=8$) Style Guidance Unsure
+@eveismyname 16 8 1
+Paris 2024 19 2 4
+DPAgg-TI (no DP, subsamp. $m=8$) DPAgg-TI ($\varepsilon=1$, subsamp. $m=8$) Unsure
+@eveismyname 8 5 12
+Paris 2024 15 4 6
+Figure 7: Samples of image sets used in our user study. Participants are asked to compare 2 images at a time.
+Appendix B Kernel Inception Distance Discussion
+Our results indicate that DPAgg-TI preserves the style transfer fidelity of TI while also ensuring differential privacy.
+Notably, for @eveismyname ($m=4$) at low privacy budgets, we observe even lower KID values than standard TI, suggesting enhanced style alignment.
+Similarly, results for the Paris 2024 dataset follow a comparable trend, with DPAgg-TI achieving KID scores similar to TI at low privacy budgets.
+However, the overall KID scores for this dataset remain high within the context of diffusion model style transfer.
+Upon inspecting the generated images (Figure 8), we hypothesize that the abstract and out-of-distribution nature of the Paris 2024 images poses a challenge for the Inception network, leading to less meaningful feature embeddings.
+This likely inflates the measured embedding distances between generated and reference images, resulting in higher-than-expected KID values.
+For KID evaluations, we used prompts similar to those employed during TI training: “A painting/icon in the style of $S^{*}$”.
+Consistent with the training image captions, these prompts do not specify a subject.
+Figure 8: Sample of generated images for KID evaluations with respect to the Paris 2024 dataset.
+Appendix C Style Guidance
+C.1 Background: Denoising Diffusion Implicit Models
+Denoising Diffusion Implicit Models (DDIM) sampling (Song et al., 2021a) uses the predicted noise $\epsilon_{\theta}(x_{t},y,t)$ and a noise schedule represented by an array of scalars $\{\alpha_{t}\}_{t=1}^{T}$ to first predict a clean image $\hat{x}_{0}$, then makes a small step in the direction of $\hat{x}_{0}$ to obtain $x_{t-1}$.
+The reverse diffusion process for DDIM sampling can be formalized as follows:
+$\hat{x}_{0}=\frac{x_{t}-\sqrt{1-\alpha_{t}}\epsilon_{\theta}(x_{t},y,t)}{\sqrt{\alpha_{t}}}$ (8)
+$x_{t-1}=\sqrt{\alpha_{t-1}}\hat{x}_{0}+\sqrt{1-\alpha_{t-1}}\epsilon_{\theta}(x_{t},y,t).$ (9)
+C.2 Implementation
+We follow the style guidance process introduced by Bansal et al.
+(2024), modifying it to include differential privacy mechanisms.
+Let $x_{c}$ denote the target style image, $x_{t}$ the noisy image at step $t$, and $\mathcal{E}(\cdot)$ the CLIP image encoder.
+The forward guidance process is defined as follows:
+$\hat{\epsilon}_{\theta}(x_{t},y,t)=\epsilon_{\theta}(x_{t},y,t)+w\sqrt{1-\alpha_{t}}\nabla_{x_{t}}\ell_{\cos}(\mathcal{E}(x_{c}),\mathcal{E}(\hat{x}_{0})),$ (10)
+where $w$ is a guidance weight and $\ell_{\cos}$ is the negative cosine similarity loss.
+For a detailed description of Universal Guidance, including the backward guidance process and per-step self-recurrence, we refer the reader to the original paper.
+The reverse diffusion step replaces $\epsilon_{\theta}(x_{t},y,t)$ with $\hat{\epsilon}_{\theta}(x_{t},y,t)$, generating an image $x_{0}$ that aligns with the text conditioning $y$ while incorporating the stylistic characteristics of $x_{c}$.
+To integrate differential privacy, we encode each target image $x^{(i)}$ into $u^{(i)}=\mathcal{E}(x^{(i)})$ and aggregate these embeddings into $u^{*}$ using the centroid method.
+The aggregated $u^{*}$ guides the reverse diffusion process:
+$\hat{\epsilon}_{\theta}(x_{t},y,t)=\epsilon_{\theta}(x_{t},y,t)+w\sqrt{1-\alpha_{t}}\nabla_{x_{t}}\ell_{\cos}(u^{*},\mathcal{E}(\hat{x}_{0})).$ (11)
+This ensures privacy-preserving style transfer while maintaining high stylistic fidelity.
+C.3 Ablation
+Figure 9: Sample of paintings by Van Gogh used to generate style guidance embeddings.
+Figure 10: Images generated by Stable Diffusion v1.5 with style guidance towards Van Gogh’s Saint-Paul Asylum, Saint-Rémy collection using prompts “A painting of Taylor Swift (left) / the Eiffel Tower (center) / a tree (right)”.
+To better understand the limited effectiveness of style guidance in our experiments, despite its success in (Bansal et al., 2024), we applied our approach to a dataset of 143 paintings from Van Gogh’s Saint-Paul Asylum, Saint-Rémy collection(Innat, ) (Figure 9).
+Unlike the @eveismyname and Paris 2024 datasets, it is highly likely that Stable Diffusion has been trained on these images.
+Additionally,  Bansal et al.
+(2024) demonstrated successful adaptation towards the style of Van Gogh’s Starry Night as a single reference image, making this dataset a reasonable interpolation between their successful results and our more limited findings.
+Without DP noise or subsampling, we obtained reasonable style transfer results, as shown in Figure 10.
+This suggests that style guidance struggles when applied to previously unseen target styles, and that its effectiveness may depend on prior exposure within the pre-training data.
+Appendix D Additional Style Transfer and Ablation Results
+Figure 11: Images generated by Stable Diffusion v1.5 using the prompt “A painting of Taylor Swift in the style of $<$@eveismyname$>$”, with the embedding $<$@eveismyname$>$ trained using DPAgg-TI (with different subsample sizes $m$) and TI with DP-SGD using different values of $\varepsilon$.
+Figure 12: Images generated by Stable Diffusion v1.5 using the prompt “An icon of the Eiffel Tower in the style of $<$Paris 2024 Pictograms$>$”, with the embedding $<$Paris 2024 Pictograms$>$ trained using DPAgg-TI (with different subsample sizes $m$) and TI with DP-SGD using different values of $\varepsilon$.
+Figure 13: Images generated by Stable Diffusion v1.5 using the prompt “An icon of a dragon in the style of $<$Paris 2024 Pictograms$>$”, with the embedding $<$Paris 2024 Pictograms$>$ trained using DPAgg-TI (with different subsample sizes $m$) and TI with DP-SGD using different values of $\varepsilon$.

### v3

@@
 Abstract
-We introduce a novel method for adapting diffusion models under differential privacy (DP) constraints, enabling privacy-preserving style and content transfer without fine-tuning model weights.
-Traditional approaches to private adaptation, such as DP-SGD, incur significant computational and memory overhead when applied to large, complex models.
-In addition, when adapting to small-scale specialized datasets, DP-SGD incurs large amount of noise that significantly degrades the performance.
-Our approach instead leverages an embedding-based technique derived from Textual Inversion (TI) and adapted with differentially private mechanisms.
-We apply TI to Stable Diffusion for style adaptation using two private datasets: a collection of artworks by a single artist and pictograms from the Paris 2024 Olympics.
-Experimental results show that the TI-based adaptation achieves superior fidelity in style transfer, even under strong privacy guarantees.
-Machine Learning, ICML
+Personalizing large-scale diffusion models poses serious privacy risks, especially when adapting to small, sensitive datasets.
+A common approach is to fine-tune the model using differentially private stochastic gradient descent (DP-SGD), but this suffers from severe utility degradation due to the high noise needed for privacy, particularly in the small data regime.
+We propose an alternative that leverages Textual Inversion (TI), which learns an embedding vector for an image or set of images, to enable adaptation under differential privacy (DP) constraints.
+Our approach, Differentially Private Aggregation via Textual Inversion (DPAgg-TI), adds calibrated noise to the aggregation of per-image embeddings to ensure formal DP guarantees while preserving high output fidelity.
+We show that DPAgg-TI outperforms DP-SGD finetuning in both utility and robustness under the same privacy budget, achieving results closely matching the non-private baseline on style adaptation tasks using private artwork from a single artist and Paris 2024 Olympic pictograms.
+In contrast, DP-SGD fails to generate meaningful outputs in this setting.
+Figure 1: We compare our method (DPAgg-TI, top) to a baseline applying DP-SGD to Textual Inversion (bottom), using the prompt “an icon of the Eiffel Tower in the style of the Paris 2024 Olympic Pictograms.” While the baseline learns a single embedding over the dataset, our method privately aggregates per-image embeddings.
+At privacy budget $\varepsilon=1$, DPAgg-TI preserves visual fidelity much better than the baseline, and closely matches the non-private output (left), demonstrating a superior privacy-utility tradeoff.
 Figure 2: Overview of DPAgg-TI.
 We first apply Textual Inversion to extract embeddings for each image in the private dataset.
-These embeddings are then aggregated with differentially private mechansim, incorporating subsampling to produce a private embedding $u^{\star}$.
-Finally, images are generated using the corresponding token <$S^{\star}$>.
+These embeddings are then aggregated with differentially private mechansim, incorporating subsampling to produce a private embedding $u^{*}_{\text{DP}}$.
+Finally, images are generated using the corresponding token $S^{*}$>.
 1 Introduction
-In recent years, diffusion models (Ho et al., 2020; Song et al., 2021b), particularly latent diffusion models (Rombach et al., 2022), have spearheaded high quality text-to-image generation, and have been widely adopted by researchers and the general public alike.
-Trained on massive datasets like LAION-5B (Schuhmann et al., 2022), these models have developed a broad understanding of visual concepts, enabling new creative and practical applications.
-Notably, tools such as Stable Diffusion (Rombach et al., 2022; Podell et al., 2023) have been made readily accessible for general use.
-Building on this foundation, efficient adaptation methods such as parameter efficient fine-tuning (PEFT) methods (Hu et al., 2022; von Platen et al., 2022; Ruiz et al., 2023), guidance based approaches (Ho & Salimans, 2021; Kim et al., 2022; Bansal et al., 2024), and pseudo-word generation (Gal et al., 2023) enable users to leverage this extensive pretraining for customizing models that can specialize on downstream tasks with smaller datasets.
-The rapid adoption of diffusion models has raised significant privacy and legal concerns.
-These models are vulnerable to privacy attacks, such as membership inference (Duan et al., 2023), where attackers determine if a specific data point was used for training, and data extraction (Carlini et al., 2023), which enables reconstruction of training data.
+The rapid adoption of diffusion models  (Ho, Jain, and Abbeel 2020; Song et al.
+2021; Rombach et al.
+2022) has raised significant privacy and legal concerns.
+These models are vulnerable to privacy attacks, such as membership inference (Duan et al.
+2023), where attackers determine if a specific data point was used for training, and data extraction (Carlini et al.
+2023), which enables reconstruction of training data.
 This risk is amplified during fine-tuning on smaller, domain-specific datasets, where each record has a greater impact.
-Additionally, reliance on large datasets scraped without consent raises copyright concerns (Vyas et al., 2023), as diffusion models can reproduce original artworks without credit or compensation.
+Additionally, reliance on large datasets scraped without consent raises copyright concerns (Vyas, Kakade, and Barak 2023), as diffusion models can reproduce original artworks without credit or compensation.
 These issues highlight the urgent need for privacy-preserving technologies and clearer ethical and legal guidelines for generative models.
-Differential privacy (DP) (Dwork et al., 2006; Dwork, 2006) is a widely adopted approach to addressing these challenges, where controlled noise is added during training or inference to prevent information leakage from individual data points while still enabling the model to learn effectively from the overall dataset.
-One standard approach for ensuring DP in deep learning is Differentially Private Stochastic Gradient Descent (DP-SGD) (Abadi et al., 2016), which modifies traditional SGD by adding noise to clipped gradients.
+Differential privacy (DP) (Dwork 2006) is a widely adopted framework for addressing these challenges.
+One standard approach for ensuring DP in deep learning is Differentially Private Stochastic Gradient Descent (DP-SGD) (Abadi et al.
+2016), which modifies traditional SGD by adding noise to clipped gradients.
 However, applying DP-SGD to train diffusion models poses several challenges.
-It introduces significant computational and memory overhead due to per-sample gradient clipping (Hoory et al., 2021), which is essential for bounding gradient sensitivity (Dwork et al., 2006; Abadi et al., 2016).
+It introduces significant computational and memory overhead due to per-sample gradient clipping (Hoory et al.
+2021), which is essential for bounding gradient sensitivity (Dwork et al.
+2006; Abadi et al.
+2016).
 DP-SGD is also incompatible with batch-wise operations like batch normalization, as these link samples and hinder sensitivity analysis.
@@
 Consequently, existing diffusion models trained with DP-SGD are limited to relatively small-scale images.
-(Dockhorn et al., 2023; Ghalebikesabi et al., 2023).
-As a result, recent research has focused on privacy-preserving strategies for adapting diffusion models without the need for full DP-SGD training.
-One approach adapts large, publicly pre-trained models to new domains under DP constraints, leveraging their representational strengths while reducing computational and memory costs (Ghalebikesabi et al., 2023).
-Similarly, PEFT methods like DP-LoRA (Yu et al., 2022) fine-tune a small subset of parameters, enabling efficient adaptation with lower privacy costs.
-Methods like DP-RDM (Lebensold et al., 2024) avoid direct model updates by using retrieval mechanisms that condition image generation on private data retrieved during inference.
-However, these alternate approaches often fail to capture the detail of style, underscoring the challenges of balancing privacy, efficiency, and generative performance.
-Independent of privacy concerns, Textual Inversion (TI) (Gal et al., 2023) provides an effective method for adapting diffusion models to specific styles or content without modifying the model.
+(Dockhorn et al.
+2023; Ghalebikesabi et al.
+2023).
+Independent of privacy concerns, Textual Inversion (TI) (Gal et al.
+2023) provides an effective method for adapting diffusion models to specific styles or content without modifying the model.
 Instead, TI learns an external embedding vector that captures the style or content of a target image set, which is then incorporated into text prompts to guide the model’s outputs.
@@
 Additionally, since TI avoids direct model optimization, it remains efficient and compatible with DP constraints on smaller datasets.
-In this work, we propose a novel privacy-preserving adaptation method for smaller datasets, leveraging TI to avoid the extensive model updates required by DP-SGD or DP-PEFT approaches.
-Standard TI inherently compresses the target dataset into a single, low-dimensional vector, providing some obfuscation benefits, but it does not offer formal privacy guarantees.
-To address this limitation, we introduce a private variant of TI, called Differentially Private Aggregation via Textual Inversion (DPAgg-TI) and summarize it in Figure 1.
+In this work, we propose a novel privacy-preserving adaptation method for smaller datasets, leveraging TI to avoid the extensive model updates required by DP-SGD.
+Standard TI does not offer formal privacy guarantees, so to address this limitation, we introduce a private variant of TI, called Differentially Private Aggregation via Textual Inversion (DPAgg-TI) and summarize it in Figure 2.
 Our method decouples interactions among samples by learning a separate embedding for each target image, which are then aggregated into a noisy centroid.
@@
 Our experiments demonstrate the effectiveness of DPAgg-TI, showing that TI remains robust in preserving stylistic fidelity even under privacy constraints.
-Applying our method to a private artwork collection by @eveismyname and Paris 2024 Olympics pictograms (Paris 2024, ), we show that DPAgg-TI captures nuanced stylistic elements while ensuring privacy.
+Applying our method to a private artwork collection by @eveismyname and Paris 2024 Olympics pictograms (Paris 2024 2023), we show that DPAgg-TI captures nuanced stylistic elements while ensuring privacy.
 We observe a trade-off between privacy (controlled by DP parameter $\varepsilon$) and image quality: lower $\varepsilon$ reduces fidelity but maintains the target style under moderate noise.
@@
   • We validate our method on diverse datasets, demonstrating its effectiveness in capturing stylistic elements under privacy constraints.
-2 Preliminaries and Related Work
+2 Background and Related Work
 2.1 Diffusion Models
-Diffusion models (Ho et al., 2020; Song et al., 2021b, a; Rombach et al., 2022) leverage an iterative denoising process to generate high-quality images that align with a given conditional input from random noise.
+Diffusion models (Ho, Jain, and Abbeel 2020; Song et al.
+2021; Song, Meng, and Ermon 2021; Rombach et al.
+2022) leverage an iterative denoising process to generate high-quality images that align with a given conditional input from random noise.
 In text-to-image generation, this conditional input is based on a textual description (a prompt) that guides the model in shaping the image to reflect the content and style specified by the text.
@@
 These tokens are then converted into a sequence of embedding vectors $v_{i}$ that encapsulate the meaning of each token within the model’s semantic space.
-Next, these embeddings pass through a transformer text encoder, such as CLIP (Radford et al., 2021), outputting a single text-conditional vector $y$ that serves as the conditioning input.
+Next, these embeddings pass through a transformer text encoder, such as CLIP (Radford et al.
+2021), outputting a single text-conditional vector $y$ that serves as the conditioning input.
 This vector $y$ is then incorporated at each denoising step, guiding the model to align the output image with the specific details outlined in the prompt.
@@
 Textual Inversion.
-Textual Inversion (TI) (Gal et al., 2023) is an adaptation technique that enables personalization using a small dataset of typically 3-5 images.
+Textual Inversion (TI) (Gal et al.
+2023) is an adaptation technique that enables personalization using a small dataset of typically 3-5 images.
 This approach essentially learns a new token that encapsulates the semantic meaning of the training images, allowing the model to associate specific visual features with a custom token.
 To achieve this, TI trains a new token embedding, denoted as $u$, representing a placeholder token, denoted as $S$.
-During training, images are conditioned on phrases such as “A photo of $S$” or “A painting in the style of $S$”.
+During training, images are conditioned on phrases such as “A photo of SS” or “A painting in the style of SS”.
 However, unlike the fixed embeddings of typical tokens $v_{i}$, $u$ is a learnable parameter.
@@
 Through gradient descent, TI minimizes the diffusion model loss given in (1) with respect to $u$, while keeping the diffusion model parameters $\theta$ fixed, iteratively refining this embedding to capture the unique characteristics of the training images.
-The resulting optimal embedding $u_{*}$ is formalized as follows:
-$u_{*}=\arg\min_{u}\mathbb{E}_{x,\epsilon\sim\mathcal{N}(0,I),t\sim[T]}[\|\epsilon-\epsilon_{\theta}(x_{t},y_{u},t)\|^{2}].$ (2)
-Hence, $u_{*}$ represents an optimized placeholder token $S_{*}$, which can employed in prompts such as “A photo of $S_{*}$ floating in space” or “A drawing of a capybara in the style of $S_{*}$”, enabling the generation of personalized images that reflect the learned visual characteristics.
-2.2 Differential Privacy
-In this work, we adopt differential privacy (DP) (Dwork et al., 2006; Dwork, 2006) as our privacy framework.
+The resulting optimal embedding $u^{*}$ is formalized as follows:
+$u^{*}=\arg\min_{u}\mathbb{E}_{x,\epsilon\sim\mathcal{N}(0,I),t\sim[T]}[\|\epsilon-\epsilon_{\theta}(x_{t},y_{u},t)\|^{2}].$ (2)
+Hence, $u^{*}$ represents an optimized placeholder token $S^{*}$, which can employed in prompts such as “A photo of $S^{*}$ floating in space” or “A drawing of a capybara in the style of $S^{*}$”, enabling the generation of personalized images that reflect the learned visual characteristics.
+2.2 Differential Privacy.
+In this work, we adopt differential privacy (DP) (Dwork et al.
+2006; Dwork 2006) as our privacy framework.
 Over the past decade, DP has become the gold standard for privacy protection in both research and industry.
 It measures the stability of a randomized algorithm with respect to changes in an input instance, thereby quantifying the extent to which an adversary can infer the existence of a specific input based on the algorithm’s output.
-Definition 2.1 ((Approximate) Differential Privacy).
+Definition 1 ((Approximate) Differential Privacy).
 For $\varepsilon,\delta\geq 0$, a randomized mechanism $\mathcal{M}:\mathcal{X}^{n}\rightarrow\mathcal{Y}$ satisfies $(\varepsilon,\delta)$-DP if for all neighboring datasets $\mathcal{D},\mathcal{D}^{\prime}\in\mathcal{X}^{n}$ which differ in a single record (i.e., $\left\lVert\mathcal{D}-\mathcal{D}^{\prime}\right\rVert_{\mathsf{H}}\leq 1$ where $\left\lVert\cdot\right\rVert_{\mathsf{H}}$ is the Hamming distance) and all measurable $\mathcal{S}$ in the range of $\mathcal{M}$, we have that
@@
 When $\delta=0$, we say $\mathcal{M}$ satisfies $\varepsilon$-pure DP or ($\varepsilon$-DP).
-To achieve DP, the Gaussian mechanism is often applied (Dwork et al., 2014; Balle & Wang, 2018), adding Gaussian noise scaled by the sensitivity of the function $f$ and privacy parameters $\varepsilon$ and $\delta$.
-Specifically, noise with standard deviation $\sigma=\frac{\Delta_{f}\sqrt{2\ln(1.25/\delta)}}{\varepsilon}$ is added to the output [^1] In practice, we use numerical privacy accountant such as (Balle & Wang, 2018; Mironov, 2017) to calibrate the noise.
-(Balle & Wang, 2018), where $\Delta_{f}$ represents $\ell_{2}$-sensitivity of the target function $f(\cdot)$.
+To achieve DP, the Gaussian mechanism is often applied (Dwork, Roth et al.
+2014; Balle and Wang 2018), adding Gaussian noise scaled by the sensitivity of the function $f$ and privacy parameters $\varepsilon$ and $\delta$.
+Specifically, noise with standard deviation $\sigma=\frac{\Delta_{f}\sqrt{2\ln(1.25/\delta)}}{\varepsilon}$ is added to the output [^1] In practice, we use numerical privacy accountant such as (Balle and Wang 2018; Mironov 2017) to calibrate the noise.
+(Balle and Wang 2018), where $\Delta_{f}$ represents $\ell_{2}$-sensitivity of the target function $f(\cdot)$.
 When the context is clear, we may omit the subscript $f$.
-This mechanism enables a smooth privacy-utility tradeoff and is widely used in privacy-preserving machine learning, including in DP-SGD (Abadi et al., 2016), which applies Gaussian noise during model updates to achieve DP.
+This mechanism enables a smooth privacy-utility tradeoff and is widely used in privacy-preserving machine learning, including in DP-SGD (Abadi et al.
+2016), which applies Gaussian noise during model updates to achieve DP.
 Privacy Amplification by Subsampling.
@@
 $\varepsilon^{\prime}=\log\left(1+\frac{m}{n}\left(e^{\varepsilon}-1\right)\right)=O\left(\frac{m}{n}\varepsilon\right).$ (3)
-This result is well-known (Steinke (2022, Theorem 29)), with tighter amplification bounds available for Gaussian mechanisms (Mironov, 2017).
-2.3 Differentially Private Adaptation of Diffusion Models
+This result is well-known (Steinke (2022, Theorem 29)), with tighter amplification bounds available for Gaussian mechanisms (Mironov 2017).
+2.3 Private Adaptation of Diffusion Models
 Recent advancements in applying DP to diffusion models have aimed to balance privacy preservation with the high utility of generative outputs.
 Dockhorn et al.
-(Dockhorn et al., 2023) proposed a Differentially Private Diffusion Model (DPDM) that enables privacy-preserving generation of realistic samples, setting a foundational approach for adapting diffusion processes using DP-SGD.
+Dockhorn et al.
+(2023) proposed a Differentially Private Diffusion Model (DPDM) that enables privacy-preserving generation of realistic samples, setting a foundational approach for adapting diffusion processes using DP-SGD.
 Another common strategy involves training a model on a large public dataset, followed by differentially private fine-tuning on a private dataset, as explored by Ghalebikesabi et al.
 (2023).
-While effective in certain contexts, this approach raises privacy concerns, particularly around risks of information leakage during the fine-tuning phase (Tramèr et al., 2024).
+While effective in certain contexts, this approach raises privacy concerns, particularly around risks of information leakage during the fine-tuning phase (Tramèr, Kamath, and Carlini 2024).
 In response to these limitations, various adaptation techniques have emerged.
-Although not specific to diffusion models, some methods focus on training models on synthetic data followed by DP-constrained fine-tuning, as in the VIP approach (Yu et al., 2024), which demonstrates the feasibility of applying DP in later adaptation stages.
-Other approaches explore differentially private learning of feature representations (Sander et al., 2024), aiming to distill private information into a generalized embedding space while maintaining DP guarantees.
+Although not specific to diffusion models, some methods focus on training models on synthetic data followed by DP-constrained fine-tuning, as in the VIP approach (Yu et al.
+2024), which demonstrates the feasibility of applying DP in later adaptation stages.
+Other approaches explore differentially private learning of feature representations (Sander et al.
+2024), aiming to distill private information into a generalized embedding space while maintaining DP guarantees.
 Although these adaptations are not yet implemented for diffusion models, they lay essential groundwork for developing secure and efficient privacy-preserving generative models.
 Figure 3: Samples of images used in our style adaptation experiments. Left: artwork by @eveismyname ($n=158$). Right: Paris 2024 Olympic pictograms ($n=47$), © International Olympic Committee, 2023.
+Figure 4: Images generated by Stable Diffusion v1.5 using the prompt “A painting of Taylor Swift in the style of <<@eveismyname>>”, with the embedding <<@eveismyname>> trained using different values of $m$ and $\varepsilon$.
+Figure 5: Images generated by Stable Diffusion v1.5 using the prompt “Icon of a dragon in the style of <<Paris 2024 Pictograms>>”, with the embedding <<Paris 2024 Pictograms>> trained using different values of $m$ and $\varepsilon$.
 3 Differentially Private Adaptation via Textual Inversion
-TI is inherently parameter-efficient and offers certain privacy benefits, as information from an entire dataset of images is compressed into a single token embedding vector.
-This compression limits the model’s capacity to memorize specific images, making data extraction attacks difficult.
-However, this privacy is merely heuristic and yet to be proven, so TI may still be vulnerable to privacy attacks such as membership inference.
-A similar adaptation technique with privacy guarantees may therefore be desirable.
 Let $x^{(1)},\dots,x^{(n)}$ represent a target dataset of images whose characteristics we wish to privately adapt our image generation towards.
@@
 The purpose of this aggregation is to limit the sensitivity of the final output to each $x^{(i)}$.
-In order to provide DP guanrantees, we also add isotropic Gaussian noise to the centroid.
-We can therefore define the resulting embedding vector $u^{*}$ as follows:
-$u^{*}=\frac{1}{n}\sum_{i=1}^{n}u^{(i)}+\mathcal{N}(0,\sigma^{2}I),$ (5)
-where the minimum $\sigma$ required to provide $(\varepsilon,\delta)$-DP is given by the following expression based on Balle & Wang (2018, Theorem 1):
+In order to provide DP guarantees, we also add isotropic Gaussian noise to the centroid.
+We can therefore define the resulting embedding vector $u^{*}_{\text{DP}}$ as follows:
+$u^{*}_{\text{DP}}=\frac{1}{n}\sum_{i=1}^{n}u^{(i)}+\mathcal{N}(0,\sigma^{2}I),$ (5)
+where the minimum $\sigma$ required to provide $(\varepsilon,\delta)$-DP is given by the following expression based on Balle and Wang (2018, Theorem 1):
 $\sigma=\frac{\Delta}{n}\cdot\frac{\sqrt{2\ln(1.25/\delta)}}{\varepsilon}.$ (6)
 In the context of our problem, $\Delta=\sup_{i,j}\|u^{(i)}-u^{(j)}\|$. Since our embedding vectors are directional, we can normalize each $u^{(i)}$, allowing us to set $\Delta=2$.
-The noisy centroid embedding $u^{*}$ can then be used to adapt the downstream image generation process.
-Similar to regular TI’s $u_{*}$, we can use $u^{*}$ to represent a new placeholder token $S^{*}$ that can be incorporated into prompts for personalized image generation.
-While $u^{*}$ may not fully solve the TI optimization problem presented in (2), it provides provable privacy guarantees, with only a minimal trade-off in accurately representing the style of the target dataset.
+The noisy centroid embedding $u^{*}_{\text{DP}}$ can then be used to adapt the downstream image generation process.
+Similar to regular TI’s $u^{*}$, we can use $u^{*}_{\text{DP}}$ to represent a new placeholder token $S^{*}$ that can be incorporated into prompts for personalized image generation.
+While $u^{*}_{\text{DP}}$ may not fully solve the TI optimization problem presented in (2), it provides provable privacy guarantees, with only a minimal trade-off in accurately representing the style of the target dataset.
 To reduce the amount of noise needed to provide the same level of DP, we employ subsampling: instead of computing the centroid over all $n$ embedding vectors, we randomly sample $m\leq n$ embedding vectors without replacement and compute the centroid over only the sampled vectors.
@@
 Formally, we sample $D_{\text{sub}}\subseteq\{u^{(1)},\dots,u^{(n)}\}$ where $|D_{\text{sub}}|=m$, and compute the output embedding as follows:
-$u^{*}=\frac{1}{m}\sum_{u^{(i)}\in D_{\text{sub}}}u^{(i)}+\mathcal{N}(0,\sigma^{2}I),$ (7)
+$u^{*}_{\text{DP}}=\frac{1}{m}\sum_{u^{(i)}\in D_{\text{sub}}}u^{(i)}+\mathcal{N}(0,\sigma^{2}I),$ (7)
 where $\sigma$ can be computed numerically for any target $\varepsilon,\delta$ and subsampling rate $\frac{m}{n}$.
@@
 This dataset serves as a controlled test for privacy-preserving style transfer on individual artistic collections.
-The second dataset contains 47 pictograms from the Paris 2024 Olympics (Paris 2024, ), permitted strictly for non-commercial editorial use (International Olympic Committee, ).
+The second dataset contains 47 pictograms from the Paris 2024 Olympics (Paris 2024 2023), permitted strictly for non-commercial editorial use (International Olympic Committee 2025).
 These pictograms were officially released in February 2023, several months after the release of Stable Diffusion v1.5, ensuring they were absent from the model’s pretraining data.
@@
 4.2 Style Transfer Results
-Figure 3: Images generated by Stable Diffusion v1.5 using the prompt “A painting of Taylor Swift in the style of $<$@eveismyname$>$”, with the embedding $<$@eveismyname$>$ trained using different values of $m$ and $\varepsilon$.
-Figure 4: Images generated by Stable Diffusion v1.5 using the prompt “Icon of a dragon in the style of $<$Paris 2024 Pictograms$>$”, with the embedding $<$Paris 2024 Pictograms$>$ trained using different values of $m$ and $\varepsilon$.
-Using both the @eveismyname and Paris 2024 pictograms dataset, we trained TI (Gal et al., 2023) embeddings on Stable Diffusion v1.5 (Rombach et al., 2022) using DPAgg-TI.
+Using both the @eveismyname and Paris 2024 pictograms dataset, we trained TI (Gal et al.
+2023) embeddings on Stable Diffusion v1.5 (Rombach et al.
+2022) using DPAgg-TI.
 Our primary goal is to investigate how DP configurations, specifically the privacy budget $\varepsilon$ and subsampling size $m$, affect the generated images quality and privacy resilience.
@@
 For the DPAgg-TI, we test multiple configurations of $m$ and $\varepsilon$ to analyze the trade-off between image fidelity and privacy.
-Figures 3 and  4 present generated images across two key configurations: (1) regular TI without DP, (2) DPAgg-TI with DP at different values of $m$ and $\varepsilon$.
-We used the same random seed to generate embeddings, subsample images, and sample DP noise for ease of visual comparision between different configurations.
+Figures 4 and  5 present generated images across two key configurations: (1) regular TI without DP, (2) DPAgg-TI with DP at different values of $m$ and $\varepsilon$.
+We used the same random seed to generate embeddings, subsample images, and sample DP noise for ease of visual comparison between different configurations.
 As with common practice, we set $\delta=1/n$.
 Since $\sigma$ is undefined for $\varepsilon=0$, we demonstrate the results of $\varepsilon\approx 0$, in other words, infinite noise, by setting $\varepsilon=10^{-5}$.
-The purpose of this parameter value is to demonstrate the image generated when $u^{*}$ contains zero information about the target dataset.
+The purpose of this parameter value is to demonstrate the image generated when $u^{*}_{\text{DP}}$ contains zero information about the target dataset.
 Images generated without DP closely resemble the unique stylistic elements of the target dataset.
 In particular, images adapted using @eveismyname images displayed crisp details and nuanced color gradients characteristic of the artist’s work, while those of Paris 2024 pictograms captured the logo’s original structure.
-In contrast, DP configurations introduce a discernible degradation in image quality, with lower epsilon values and smaller subsampling sizes resulting in more noticeable noise and diminished stylistic fidelity.
-As $\varepsilon\rightarrow 0$, the resulting token embedding $u^{*}$ gradually loses its semantic meaning, leading to a loss of stylistic fidelity.
-In particular, $y_{u^{*}}$ tends towards $y$ (a conditioning vector independent of the learnable embedding).
+In contrast, DP configurations introduce a discernible degradation in image quality, with lower epsilon values and smaller subsampling sizes resulting in diminished stylistic fidelity.
+As $\varepsilon\rightarrow 0$, the resulting token embedding $u^{*}_{\text{DP}}$ gradually loses its semantic meaning, leading to a loss of stylistic fidelity.
+In particular, $y_{u^{*}_{\text{DP}}}$ tends towards $y$ (a conditioning vector independent of the learnable embedding).
 In our results, this manifests as a painting of Taylor Swift devoid of the artist-specific stylistic elements, or a generic icon of a dragon (with color, as opposed to the black and white design of the pictograms).
-With this in mind, $\varepsilon$ can be interpreted as a drift parameter, representing the progression from the optimal $u^{*}$ towards infinity, gradually steering the generated image away from the target style in exchange for stronger privacy guarantees.
-We also observe instances where there is a temporary drop in prompt fidelity (e.g., $m=16,\varepsilon\in[0.5,1]$ in Figure 3 and intermediate $\varepsilon$ values in Figure 4) which restores as $u^{*}$ drifts even further from its optimal value.
-We hypothesize that this is due to drifted $u^{*}$ capturing a different meaning unrelated to the prompt, before losing any meaning that could be interpreted by Stable Diffusion’s text encoder, causing $u^{*}$ to be disregarded from $y_{u^{*}}$ and the prompt fidelity to be restored.
-Another possible explanation is that the temporary drop in prompt fidelity is due to the drift path of $u^{*}$ passing through non-linear regions within embedding space.
+With this in mind, $\varepsilon$ can be interpreted as a drift parameter, representing the progression from the optimal $u^{*}_{\text{DP}}$ towards infinity, gradually steering the generated image away from the target style in exchange for stronger privacy guarantees.
+We also observe instances where there is a temporary drop in prompt fidelity (e.g., $m=16,\varepsilon\in[0.5,1]$ in Figure 4 and intermediate $\varepsilon$ values in Figure 5) which restores as $u^{*}_{\text{DP}}$ drifts even further from its optimal value.
+We hypothesize that this is due to drifted $u^{*}_{\text{DP}}$ capturing a different meaning unrelated to the prompt, before losing any meaning that could be interpreted by Stable Diffusion’s text encoder, causing $u^{*}_{\text{DP}}$ to be disregarded from $y_{u^{*}_{\text{DP}}}$ and the prompt fidelity to be restored.
+Another possible explanation is that the temporary drop in prompt fidelity is due to the drift path of $u^{*}_{\text{DP}}$ passing through non-linear regions within embedding space.
 We leave further investigations into this observation for future work.
@@
 This strong boost in robustness comes at a small price of base style capture fidelity.
-As observed in Figures 3 and 4, we can also treat subsampling as an introduction of noise.
+As observed in Figures 4 and 5, we can also treat subsampling as an introduction of noise.
 Mathematically, the subsample centroid is an unbiased estimate of the true centroid, and so the subsampling process itself defines a distribution centered at the true centroid.
@@
 4.3 Quantitative Evaluation
-4.3.1 User Study
+User Study
 To evaluate the utility of our approach under different DP and subsampling configurations, we conducted a user study with 25 participants.
@@
 Preferences at $\varepsilon=1$ were split, but subsampling was generally favored, reinforcing its role in reducing noise impact while preserving style.
-4.3.2 Kernel Inception Distance
-The Kernel Inception Distance (KID) (Bińkowski et al., 2018) is a metric for evaluating generative models by measuring the difference between the distributions of generated and training images in an embedding space.
-To compute KID, images generated by the model and real training images are passed through an Inception network (Szegedy et al., 2015), and their distributional differences are estimated.
-Unlike the more commonly used Fréchet Inception Distance (FID) (Heusel et al., 2017), KID is an unbiased estimator of the true divergence between the learned and target distributions (Jayasumana et al., 2024), making it more suitable for smaller datasets, as in our case.
-We report KID scores for different parameters in Tables 2 and 2, showing that DPAgg-TI maintains the style transfer fidelity of TI while ensuring differential privacy.
-Further discussion of these results is provided in Appendix B.
-Table 1: KID scores of DPAgg-TI on @eveismyname dataset for various $\varepsilon$ values ranging from $\varepsilon=10^{-5},0.1,0.5,1.0,2.0,5.0$ (including no DP) under different subsampling levels ($m=4,8,16,32$) as well as regular TI (ctrl).
-$m$ No DP $\varepsilon=$ 5.0 $\varepsilon=$ 2.0 $\varepsilon=$ 1.0 $\varepsilon=$ 0.5 $\varepsilon=$ 0.1 $\varepsilon\approx$ 0
-0.0444 0.0794 0.0422 0.0529 0.0690 0.1117 0.0654
-32 0.0752 0.0845 0.0865 0.1167 0.0300 0.0649 0.0657
-16 0.0351 0.0379 0.0430 0.0663 0.1309 0.0438 0.0658
-8 0.0359 0.0366 0.0350 0.0366 0.0396 0.0530 0.0658
-4 0.0245 0.0250 0.0249 0.0250 0.0258 0.0314 0.0653
-ctrl 0.0318
-Table 2: KID scores of DPAgg-TI on Paris dataset for various $\varepsilon$ values ranging from $\varepsilon=1e-5,0.1,0.5,1.0,2.0,5.0$ (including no DP) under different subsampling levels ($m=4,8,16,32$) as well as regular TI (ctrl).
-$m$ No DP $\varepsilon=$ 5.0 $\varepsilon=$ 2.0 $\varepsilon=$ 1.0 $\varepsilon=$ 0.5 $\varepsilon=$ 0.1 $\varepsilon\approx$ 0
-0.1146 0.1202 0.1368 0.1314 0.1389 0.1209 0.1274
-32 0.1220 0.1036 0.1258 0.1377 0.1307 0.1245 0.1259
-16 0.1311 0.1424 0.1170 0.1311 0.1381 0.1335 0.1278
-8 0.1317 0.1307 0.1220 0.1117 0.1295 0.1313 0.1272
-4 0.1141 0.1094 0.1137 0.1190 0.1194 0.1583 0.1259
-ctrl 0.1388
-4.4 Ablation Study
-4.4.1 Textual Inversion with DP-SGD
-Figure 5: Comparing our approach to applying DP-SGD to regular TI using prompt “a painting of Taylor Swift in the style of @eveismyname”.
+Kernel Inception Distance
+The Kernel Inception Distance (KID) (Bińkowski et al.
+2018) is a metric for evaluating generative models by measuring the difference between the distributions of generated and training images in an embedding space.
+To compute KID, images generated by the model and real training images are passed through an Inception network (Szegedy et al.
+2015), and their distributional differences are estimated.
+Unlike the more commonly used Fréchet Inception Distance (FID) (Heusel et al.
+2017), KID is an unbiased estimator of the true divergence between the learned and target distributions (Jayasumana et al.
+2024), making it more suitable for smaller datasets, as in our case.
+We report KID scores for different parameters in Tables  2 and 3 (see Appendix B), showing that DPAgg-TI maintains the style transfer fidelity of TI while ensuring differential privacy.
+Further discussion of these results is also provided in Appendix B.
+4.4 Ablation Study: Textual Inversion with DP-SGD
 A natural question that arises is how well our approach compares to the naive method of applying DP-SGD to regular TI training.
@@
 We believe that this is simply because DP-SGD is not designed to handle such small datasets in the order of 100 images.
-Additional results can be found in Appendix D.
-4.4.2 Differentially Private Adaptation Using Style Guidance
-Figure 6: Attempts of using universal guidance to generate drawings of Taylor Swift and icons of the Eiffel Tower in the styles of @eveismyname and Paris 2024 Pictograms respectively.
-Here, we apply no subsampling or DP-noise.
-We extend our approach to style guidance (SG) by leveraging the framework of Universal Guidance (Bansal et al., 2024).
-Specifically, we focus on CLIP-based style guidance, which optimizes the similarity between the CLIP embeddings of a target image and the generated image.
-We encode each target image $x^{(i)}$ as $u^{(i)}$ via a CLIP image encoder, then aggregate the embeddings $u^{(1)},\dots,u^{(n)}$ into $u^{*}$ using (5) or (7), depending on whether subsampling is applied.
-The aggregated embedding $u^{*}$ is then incorporated into the reverse diffusion process as a style guide.
-Further implementation detailes are provided in Appendix C.
-We apply our SG-based approach to both datasets.
-While it provides privacy protection by obfuscating embedding details, the resulting images captured only generalized stylistic elements and lack the detailed fidelity and coherence achieved with the TI-based method.
-As shown in Figure 6, this highlights the superiority of TI in balancing privacy and high-quality image generation.
-The reduced effectiveness of SG for style transfer may stem from its sensitivity to hyperparameters such as the guidance weight $w$, leading to instability.
-Although Bansal et al.
-(2024) proposed remedies, namely backward guidance and per-step self-recurrence, these proved insufficient for our application.
-Additionally, the CLIP embeddings may not retain enough stylistic detail after the aggregation.
+Additional results can be found in Appendix F.
+Figure 6: Comparing our approach to applying DP-SGD to regular TI using prompts “an icon of a dragon in the style of the Paris 2024 Olympic Pictograms” and “a painting of Taylor Swift in the style of @eveismyname” respectively.
+Note that our method aggregates individual TI embeddings for each training image, whereas the baseline trains a single TI embedding over the entire dataset.
+5 Copyright Protection Implications
+Our proposed mechanism can also be interpreted through the lens of copyright protection.
+This connection is grounded in the framework of Near Access-Freeness (NAF) (Vyas, Kakade, and Barak 2023), which evaluates whether a model’s outputs reveal undue influence from specific data points by comparing them to those from a safe model trained without access to the same data.
+Since DPAgg-TI satisfies $\varepsilon$-DP, it also satisfies $\varepsilon$-NAF, which means the adapted model behaves similarly to one that never saw the private images, under the NAF criterion.
+However, we emphasize that this guarantee holds only within the NAF framework; it does not constitute a general claim about content similarity or legal compliance.
+Crucially, DPAgg-TI is designed to adapt to the style of private images, not their specific content.
+Prior work and legal precedent suggest that style imitation is generally considered fair use and does not constitute infringement (Vyas, Kakade, and Barak 2023).
+Thus, our mechanism aligns with the intended protections of NAF: it avoids memorization while still enabling meaningful personalization and stylistic adaptation.
+We defer the details of copyright protection to Appendix D.
 6 Conclusion
-We presented a differentially private adaptation method for diffusion models using Textual Inversion for privacy-preserving style transfer.
-Experiments on private artwork and Paris 2024 pictograms showed TI preserves stylistic fidelity and outperforms Style Guidance.
-Our results demonstrate embedding-driven methods as efficient, scalable alternatives to DP-SGD, balancing style quality and privacy.
-Impact Statement
+We presented a differentially private adaptation method for diffusion models based on Textual Inversion, enabling privacy-preserving style transfer without the need for full model fine-tuning.
+By learning per-image embeddings and aggregating them with calibrated noise, our method, DPAgg-TI, achieves strong formal privacy guarantees while maintaining high output fidelity.
+Experiments on private artwork and Paris 2024 pictograms show that DPAgg-TI consistently outperforms DP-SGD, which fails to produce meaningful results under comparable privacy budgets.
+These results highlight the effectiveness of embedding-level adaptation as an efficient and scalable alternative to traditional gradient-based approaches, especially in low-data regimes.
+Unlike DP-SGD, which introduces significant computational overhead and utility degradation, DPAgg-TI is lightweight, modular, and compatible with existing diffusion backbones.
+Our findings suggest that embedding-centric approaches offer a promising direction for privacy-aware personalization, and motivate further research into cross-modal extensions, improved aggregation techniques, and integration with broader privacy-preserving frameworks.
+Acknowledgements
+We sincerely thank @eveismyname for providing her artwork for use in this study.
+We are also grateful to Anwar Hithnawi and Varun Chandrasekaran for their insightful discussions and feedback, as well as to all participants in our user study.
+Sanmi Koyejo acknowledges support by NSF 2046795 and 2205329, IES R305C240046, the MacArthur Foundation, Stanford HAI, OpenAI, and Google.
+Ethical Statement
 The use of images without owner consent raises significant ethical concerns, particularly regarding the exploitation of intellectual property.
@@
 Beyond creative applications, the proposed method has broader potential uses, including synthetic data generation, privacy-preserving personalization, and fine-tuning diffusion models for private or domain-specific tasks.
-Acknowledgements
-We sincerely thank Tatchamon Wongworakul (@eveismyname) for providing her artwork for use in this study.
-We are also grateful to Anwar Hithnawi and Varun Chandrasekaran for their insightful discussions and feedback, as well as to all participants in our user study.
-Sanmi Koyejo acknowledges support by NSF 2046795 and 2205329, IES R305C240046, the MacArthur Foundation, Stanford HAI, OpenAI, and Google.
 Appendix A User Study
+regular TI No Adaptation Unsure
+@eveismyname 19 4 2
+Paris 2024 16 6 3
+DPAgg-TI (no DP, no subsampling) No Adaptation Unsure
+@eveismyname 16 9 0
+Paris 2024 15 4 6
+regular TI DPAgg-TI (no DP, no subsamp.) Unsure
+@eveismyname 12 13 0
+Paris 2024 9 10 6
+regular TI DPAgg-TI (no DP, subsamp. $m=8$) Unsure
+@eveismyname 16 6 3
+Paris 2024 7 13 5
+DPAgg-TI (no DP, no subsampling) DPAgg-TI (no DP, subsamp. $m=8$) Unsure
+@eveismyname 18 4 3
+Paris 2024 10 8 7
+DPAgg-TI ($\varepsilon=1$) no subsampling DPAgg-TI ($\varepsilon=1$, subsamp. $m=8$) Unsure
+@eveismyname 14 10 1
+Paris 2024 3 16 6
+DPAgg-TI (no DP, no subsampling) Style Guidance Unsure
+@eveismyname 16 8 1
+Paris 2024 20 2 3
+DPAgg-TI ($\varepsilon=1$, subsamp. $m=8$) Style Guidance Unsure
+@eveismyname 16 8 1
+Paris 2024 19 2 4
+DPAgg-TI (no DP, subsamp. $m=8$) DPAgg-TI ($\varepsilon=1$, subsamp. $m=8$) Unsure
+@eveismyname 8 5 12
+Paris 2024 15 4 6
+Table 1: Survey Results.
+Figure 7: Samples of image sets used in our user study. Participants are asked to compare 2 images at a time.
 A.1 Study Design and Objective
@@
 These findings highlight the trade-off between increased DP robustness and reduced utility, suggesting that the optimal configuration may depend on subjective preferences and specific application requirements.
-Table 3: Survey Results.
-regular TI No Adaptation Unsure
-@eveismyname 19 4 2
-Paris 2024 16 6 3
-DPAgg-TI (no DP, no subsampling) No Adaptation Unsure
-@eveismyname 16 9 0
-Paris 2024 15 4 6
-regular TI DPAgg-TI (no DP, no subsamp.) Unsure
-@eveismyname 12 13 0
-Paris 2024 9 10 6
-regular TI DPAgg-TI (no DP, subsamp. $m=8$) Unsure
-@eveismyname 16 6 3
-Paris 2024 7 13 5
-DPAgg-TI (no DP, no subsampling) DPAgg-TI (no DP, subsamp. $m=8$) Unsure
-@eveismyname 18 4 3
-Paris 2024 10 8 7
-DPAgg-TI ($\varepsilon=1$) no subsampling DPAgg-TI ($\varepsilon=1$, subsamp. $m=8$) Unsure
-@eveismyname 14 10 1
-Paris 2024 3 16 6
-DPAgg-TI (no DP, no subsampling) Style Guidance Unsure
-@eveismyname 16 8 1
-Paris 2024 20 2 3
-DPAgg-TI ($\varepsilon=1$, subsamp. $m=8$) Style Guidance Unsure
-@eveismyname 16 8 1
-Paris 2024 19 2 4
-DPAgg-TI (no DP, subsamp. $m=8$) DPAgg-TI ($\varepsilon=1$, subsamp. $m=8$) Unsure
-@eveismyname 8 5 12
-Paris 2024 15 4 6
-Figure 7: Samples of image sets used in our user study. Participants are asked to compare 2 images at a time.
-Appendix B Kernel Inception Distance Discussion
+Appendix B Kernel Inception Distance
+$m$ No DP $\varepsilon=$ 5.0 $\varepsilon=$ 1.0 $\varepsilon=$ 0.5 $\varepsilon=$ 0.1 $\varepsilon\approx 0$
+0.0441 ± 0.0027 0.0798 ± 0.0032 0.0526 ± 0.0022 0.0688 ± 0.0020 0.1114 ± 0.0032 0.0654 ± 0.0027
+32 0.0753 ± 0.0047 0.0836 ± 0.0042 0.1166 ± 0.0037 0.0295 ± 0.0019 0.0644 ± 0.0021 0.0650 ± 0.0025
+16 0.0350 ± 0.0020 0.0381 ± 0.0018 0.0663 ± 0.0025 0.1303 ± 0.0033 0.0438 ± 0.0030 0.0660 ± 0.0029
+8 0.0359 ± 0.0018 0.0364 ± 0.0017 0.0366 ± 0.0019 0.0394 ± 0.0025 0.0527 ± 0.0033 0.0654 ± 0.0024
+4 0.0246 ± 0.0013 0.0251 ± 0.0016 0.0249 ± 0.0014 0.0256 ± 0.0012 0.0313 ± 0.0017 0.0653 ± 0.0023
+ctrl 0.0314 ± 0.0010
+Table 2: KID scores of DPAgg-TI on @eveismyname dataset for various $\varepsilon$ values ranging from $\varepsilon=10^{-5},0.1,0.5,1.0,5.0$ (including no DP) under different subsampling levels ($m=4,8,16,32$) as well as regular TI (ctrl).
+Reported values are the mean $\pm$ standard deviation over 100 random subsamples.
+$m$ No DP $\varepsilon=$ 5.0 $\varepsilon=$ 1.0 $\varepsilon=$ 0.5 $\varepsilon=$ 0.1 $\varepsilon\approx 0$
+0.1153 ± 0.0055 0.1194 ± 0.0054 0.1306 ± 0.0046 0.1395 ± 0.0057 0.1201 ± 0.0053 0.1274 ± 0.0055
+32 0.1222 ± 0.0066 0.1036 ± 0.0065 0.1375 ± 0.0047 0.1311 ± 0.0048 0.1248 ± 0.0060 0.1258 ± 0.0054
+16 0.1321 ± 0.0057 0.1411 ± 0.0077 0.1309 ± 0.0061 0.1380 ± 0.0047 0.1359 ± 0.0060 0.1273 ± 0.0057
+8 0.1303 ± 0.0084 0.1303 ± 0.0074 0.1112 ± 0.0062 0.1311 ± 0.0064 0.1318 ± 0.0052 0.1267 ± 0.0056
+4 0.1158 ± 0.0057 0.1085 ± 0.0056 0.1184 ± 0.0068 0.1194 ± 0.0065 0.1592 ± 0.0065 0.1268 ± 0.0055
+ctrl 0.1383 ± 0.0066
+Table 3: KID scores of DPAgg-TI on Paris dataset for various $\varepsilon$ values ranging from $\varepsilon=1e-5,0.1,0.5,1.0,5.0$ (including no DP) under different subsampling levels ($m=4,8,16,32$) as well as regular TI (ctrl).
+Reported values are the mean $\pm$ standard deviation over 100 random subsamples.
 Our results indicate that DPAgg-TI preserves the style transfer fidelity of TI while also ensuring differential privacy.
@@
 Consistent with the training image captions, these prompts do not specify a subject.
+For each parameter configuration, we generate 100 images and compute KID by repeatedly subsampling the larger of the real and generated sets to match the size of the smaller set, 100 times, then averaging the resulting KID scores.
 Figure 8: Sample of generated images for KID evaluations with respect to the Paris 2024 dataset.
-Appendix C Style Guidance
+Appendix C Differentially Private Adaptation via Style Guidance
 C.1 Background: Denoising Diffusion Implicit Models
-Denoising Diffusion Implicit Models (DDIM) sampling (Song et al., 2021a) uses the predicted noise $\epsilon_{\theta}(x_{t},y,t)$ and a noise schedule represented by an array of scalars $\{\alpha_{t}\}_{t=1}^{T}$ to first predict a clean image $\hat{x}_{0}$, then makes a small step in the direction of $\hat{x}_{0}$ to obtain $x_{t-1}$.
+Denoising Diffusion Implicit Models (DDIM) sampling (Song, Meng, and Ermon 2021) uses the predicted noise $\epsilon_{\theta}(x_{t},y,t)$ and a noise schedule represented by an array of scalars $\{\alpha_{t}\}_{t=1}^{T}$ to first predict a clean image $\hat{x}_{0}$, then makes a small step in the direction of $\hat{x}_{0}$ to obtain $x_{t-1}$.
 The reverse diffusion process for DDIM sampling can be formalized as follows:
@@
 C.2 Implementation
-We follow the style guidance process introduced by Bansal et al.
-(2024), modifying it to include differential privacy mechanisms.
-Let $x_{c}$ denote the target style image, $x_{t}$ the noisy image at step $t$, and $\mathcal{E}(\cdot)$ the CLIP image encoder.
-The forward guidance process is defined as follows:
-$\hat{\epsilon}_{\theta}(x_{t},y,t)=\epsilon_{\theta}(x_{t},y,t)+w\sqrt{1-\alpha_{t}}\nabla_{x_{t}}\ell_{\cos}(\mathcal{E}(x_{c}),\mathcal{E}(\hat{x}_{0})),$ (10)
+We extend our approach to style guidance (SG) by leveraging the framework of Universal Guidance (Bansal et al.
+2024).
+Specifically, we focus on CLIP-based style guidance, which optimizes the similarity between the CLIP embeddings of a target image and the generated image.
+We encode each target image $x^{(i)}$ as $u^{(i)}$ via a CLIP image encoder, then aggregate the embeddings $u^{(1)},\dots,u^{(n)}$ into $u^{*}_{\text{DP}}$ using (5) or (7), depending on whether subsampling is applied.
+The aggregated embedding $u^{*}_{\text{DP}}$ is then incorporated into the reverse diffusion process as a style guide.
+Let $x_{c}$ denote the target style image, $x_{t}$ the noisy image at step $t$, and $\mathcal{E}(\cdot)$ the CLIP image encoder. The forward guidance process is defined as follows:
+$\hat{\epsilon}_{\theta}(x_{t},y,t)=\epsilon_{\theta}(x_{t},y,t)+w\sqrt{1-\alpha_{t}}\nabla_{x_{t}}\ell_{\cos}(\mathcal{E}(x_{t}),\mathcal{E}(\hat{x}_{0})),$ (10)
 where $w$ is a guidance weight and $\ell_{\cos}$ is the negative cosine similarity loss.
@@
 The reverse diffusion step replaces $\epsilon_{\theta}(x_{t},y,t)$ with $\hat{\epsilon}_{\theta}(x_{t},y,t)$, generating an image $x_{0}$ that aligns with the text conditioning $y$ while incorporating the stylistic characteristics of $x_{c}$.
-To integrate differential privacy, we encode each target image $x^{(i)}$ into $u^{(i)}=\mathcal{E}(x^{(i)})$ and aggregate these embeddings into $u^{*}$ using the centroid method.
-The aggregated $u^{*}$ guides the reverse diffusion process:
-$\hat{\epsilon}_{\theta}(x_{t},y,t)=\epsilon_{\theta}(x_{t},y,t)+w\sqrt{1-\alpha_{t}}\nabla_{x_{t}}\ell_{\cos}(u^{*},\mathcal{E}(\hat{x}_{0})).$ (11)
+To integrate differential privacy, we encode each target image $x^{(i)}$ into $u^{(i)}=\mathcal{E}(x^{(i)})$ and aggregate these embeddings into $u^{*}_{\text{DP}}$ using the centroid method.
+The aggregated $u^{*}_{\text{DP}}$ guides the reverse diffusion process:
+$\hat{\epsilon}_{\theta}(x_{t},y,t)=\epsilon_{\theta}(x_{t},y,t)+w\sqrt{1-\alpha_{t}}\nabla_{x_{t}}\ell_{\cos}(u^{*}_{\text{DP}},\mathcal{E}(\hat{x}_{0})).$ (11)
 This ensures privacy-preserving style transfer while maintaining high stylistic fidelity.
-C.3 Ablation
-Figure 9: Sample of paintings by Van Gogh used to generate style guidance embeddings.
-Figure 10: Images generated by Stable Diffusion v1.5 with style guidance towards Van Gogh’s Saint-Paul Asylum, Saint-Rémy collection using prompts “A painting of Taylor Swift (left) / the Eiffel Tower (center) / a tree (right)”.
-To better understand the limited effectiveness of style guidance in our experiments, despite its success in (Bansal et al., 2024), we applied our approach to a dataset of 143 paintings from Van Gogh’s Saint-Paul Asylum, Saint-Rémy collection(Innat, ) (Figure 9).
+C.3 Style Transfer Results
+Figure 9: Attempts of using universal guidance to generate drawings of Taylor Swift and icons of the Eiffel Tower in the styles of @eveismyname and Paris 2024 Pictograms respectively.
+Here, we apply no subsampling or DP-noise.
+We apply our SG-based approach to both datasets.
+While it provides privacy protection by obfuscating embedding details, the resulting images captured only generalized stylistic elements and lack the detailed fidelity and coherence achieved with the TI-based method.
+As shown in Figure 9, this highlights the superiority of TI in balancing privacy and high-quality image generation.
+The reduced effectiveness of SG for style transfer may stem from its sensitivity to hyperparameters such as the guidance weight $w$, leading to instability.
+Although Bansal et al.
+(2024) proposed remedies, namely backward guidance and per-step self-recurrence, these proved insufficient for our application.
+Additionally, the CLIP embeddings may not retain enough stylistic detail after the aggregation.
+C.4 Ablation
+To better understand the limited effectiveness of style guidance in our experiments, despite its success in (Bansal et al.
+2024), we applied our approach to a dataset of 143 paintings from Van Gogh’s Saint-Paul Asylum, Saint-Rémy collection(Innat 2022) (Figure 10).
 Unlike the @eveismyname and Paris 2024 datasets, it is highly likely that Stable Diffusion has been trained on these images.
@@
 This suggests that style guidance struggles when applied to previously unseen target styles, and that its effectiveness may depend on prior exposure within the pre-training data.
-Appendix D Additional Style Transfer and Ablation Results
-Figure 11: Images generated by Stable Diffusion v1.5 using the prompt “A painting of Taylor Swift in the style of $<$@eveismyname$>$”, with the embedding $<$@eveismyname$>$ trained using DPAgg-TI (with different subsample sizes $m$) and TI with DP-SGD using different values of $\varepsilon$.
-Figure 12: Images generated by Stable Diffusion v1.5 using the prompt “An icon of the Eiffel Tower in the style of $<$Paris 2024 Pictograms$>$”, with the embedding $<$Paris 2024 Pictograms$>$ trained using DPAgg-TI (with different subsample sizes $m$) and TI with DP-SGD using different values of $\varepsilon$.
-Figure 13: Images generated by Stable Diffusion v1.5 using the prompt “An icon of a dragon in the style of $<$Paris 2024 Pictograms$>$”, with the embedding $<$Paris 2024 Pictograms$>$ trained using DPAgg-TI (with different subsample sizes $m$) and TI with DP-SGD using different values of $\varepsilon$.
+Figure 10: Sample of paintings by Van Gogh used to generate style guidance embeddings.
+Figure 11: Images generated by Stable Diffusion v1.5 with style guidance towards Van Gogh’s Saint-Paul Asylum, Saint-Rémy collection using prompts “A painting of Taylor Swift (left) / the Eiffel Tower (center) / a tree (right)”.
+Appendix D Copyright Protection
+Modern generative models typically produce outputs via randomized sampling.
+Leveraging this inherent randomness, Vyas, Kakade, and Barak (2023) introduced Near Access-Freeness (NAF) as a metric to quantify the similarity between a model’s output and copyrighted content.
+The key idea is to compare the output distribution of a potentially infringing model to that of a safe model – one trained without access to the target content.
+Formally, let $\mathsf{safe}$ be a mapping from a data point $x\in\mathcal{C}$ (where $\mathcal{C}$ is the collection of copyrighted samples) to a generative model $\mathsf{safe}(x)\in\mathcal{W}$ that is trained without using $x$.
+A canonical example is the leave-one-out-safe model, trained on the full dataset excluding $x$.
+Since $\mathsf{safe}(x)$ does not have access to $x$, the probability that it generates content resembling $x$ is exponentially small.
+Any such resemblance is considered fortuitous.
+Formally, the NAF criterion is defined as follows:
+Definition 2 (Near Access-Freeness (Vyas, Kakade, and Barak 2023)).
+Let $\mathcal{C}$ be a set of copyrighted samples and $\mathcal{W}$ a set of generative models.
+Given a mapping $\mathsf{safe}:\mathcal{C}\rightarrow\mathcal{W}$ and a divergence measure $\Delta$, we say a model $w$ is $k_{y}$-near access-free (or $k_{y}$-NAF) on prompt $y\in\mathsf{Y}$ if for every $x\in\mathcal{C}$,
+$\Delta\left(p(\cdot|y)\,\|\,\mathsf{safe}_{x}(\cdot|y)\right)\leq k_{y}.$
+If $k_{y}=0$, the model is indistinguishable from a safe model, meaning any resemblance to copyrighted material is by random chance.
+More generally, a small $k_{y}$ suggests the model is unlikely to generate outputs resembling $x$ with higher probability than a model that has never seen $x$.
+D.1 Connection to Differential Privacy
+NAF is closely related to concepts in Differential Privacy (DP) (Elkin-Koren et al.
+2023).
+Depending on the divergence measure $\Delta$, NAF resembles different DP variants – for example, $\varepsilon$-DP when $\Delta=\Delta_{\mathrm{max}}$ (Dwork et al.
+2006), and $(1,\varepsilon)$-Rényi DP when $\Delta=\Delta_{\mathrm{KL}}$.
+Translating DP to generative models yields the following definition:
+Definition 3 (Differentially Private Generation (DPG)).
+Let $S$ and $S^{\prime}$ be neighboring datasets.
+Denote by $P_{S}(\cdot|y)$ the distribution over outputs generated by a model trained on, or adapted from, $S$ with algorithm $\mathcal{A}$, where randomness includes both training and generation stages.
+The generation is said to satisfy $\varepsilon$-Differentially Private Generation ($\varepsilon$-DPG) if for every $y\in\mathcal{Y}$,
+$\Delta\left(P_{S}(\cdot|y)\,\|\,P_{S^{\prime}}(\cdot|y)\right)\leq\varepsilon.$
+Here, neighboring datasets differ by a single data point (or privacy unit).
+If the training process is $\varepsilon$-DP, then the outputs naturally satisfy $\varepsilon$-DPG via the data processing inequality.
+One benefit of DPG is the flexibility to add noise during generation rather than training, potentially improving the utility-privacy tradeoff.
+However, there are notable distinctions.
+$\varepsilon$-DP offers protection under arbitrary post-processing and multiple outputs, whereas $\varepsilon$-DPG only guarantees privacy for single outputs.
+Also, under DP, the trained model can be released, but under DPG, only the outputs are safe to share.
+Elkin-Koren et al.
+(2023) highlight further differences: NAF is one-sided—comparing a model to a fixed safe reference—whereas DPG is symmetric.
+This asymmetry in NAF can enable better utility.
+Additionally, NAF allows more flexibility in choosing the safe model, which can be exploited in algorithm design.
+Given these conceptual overlaps, both DP-SGD based training and our proposed private adaptation method DPAgg-TI satisfies $\varepsilon$-DP, so they naturally satisfy $\varepsilon$-NAF with the leave-one-out safe model.
+We emphasize that this guarantee is meaningful only within the formal framework of NAF.
+It does not imply broader legal immunity or empirical indistinguishability from the original content.
+However, within this framework, satisfying $\varepsilon$-NAF allows us to argue that any close resemblance between outputs and private training data is no more likely than would be expected from a model that never had access to that data.
+This theoretical grounding supports the privacy and safety claims of our adaptation method.
+Importantly, the goal of DPAgg-TI is to adapt to the style of a private image set—not its precise content.
+This distinction matters: style transfer is widely considered to fall under the doctrine of fair use, particularly in artistic and creative contexts.
+As discussed in Elkin-Koren et al.
+(2023) and further elaborated in legal analysis such as Carlini et al.
+(2023), generating new content in the style of a work, without reproducing its substantive elements, is generally not considered copyright infringement.
+Therefore, the use of DPAgg-TI to learn and reproduce stylistic attributes does not contradict the spirit or intent of the NAF framework.
+Instead, it offers a promising direction for responsibly fine-tuning generative models on private or copyrighted sources while respecting both privacy and intellectual property boundaries.
+Appendix E Computational Cost Comparisons
+Direct comparisons of computational cost across methods are inherently challenging due to differing training paradigms, optimization procedures, and parameter settings.
+Nonetheless, to provide a concrete sense of scale, we report representative computational costs for each method based on experiments conducted using a Stable Diffusion v1.5 model on a single NVIDIA A100 GPU.
+Below we summarize both training and inference overheads (the number of steps are optimized for each setup):
+Method Steps Batch Size Time Memory Usage
+TI (no DP) 10,000 (for 150 images) 1 25 min 7 GB
+8 2.5 hours 20 GB
+TI (DP-SGD) 30,000 (for 150 images) 1 80 min 7 GB
+8 7 hours 20 GB
+DPAgg-TI 2,000 per image N/A $\sim$5 min/image 7 GB
+SG N/A N/A N/A N/A
+Table 4: Training cost comparison across methods.
+Overhead from DP-SGD is relatively modest due to the low-dimensional embedding being optimized.
+N/A for SG means nothing is trained aside from the base model.
+Method Steps Batch Size Time Memory Usage
+TI (no DP, DP-SGD, DPAgg-TI) 50 1 1–2 sec 4 GB
+100 1 1–2 min 58 GB
+SG (no DP, DPAgg-SG) 500 1 $\sim$30 min 17 GB
+Table 5: Inference cost comparison across methods.
+Appendix F Additional Style Transfer and Ablation Results
+Figure 12: Images generated by Stable Diffusion v1.5 using the prompt “A painting of Taylor Swift in the style of <<@eveismyname>>”, with the embedding <<@eveismyname>> trained using DPAgg-TI (with different subsample sizes $m$) and TI with DP-SGD using different values of $\varepsilon$.
+Figure 13: Images generated by Stable Diffusion v1.5 using the prompt “An icon of the Eiffel Tower in the style of <<Paris 2024 Pictograms>>”, with the embedding <<Paris 2024 Pictograms>> trained using DPAgg-TI (with different subsample sizes $m$) and TI with DP-SGD using different values of $\varepsilon$.
+Figure 14: Images generated by Stable Diffusion v1.5 using the prompt “An icon of a dragon in the style of <<Paris 2024 Pictograms>>”, with the embedding <<Paris 2024 Pictograms>> trained using DPAgg-TI (with different subsample sizes $m$) and TI with DP-SGD using different values of $\varepsilon$.

### v4

@@
 These embeddings are then aggregated with differentially private mechansim, incorporating subsampling to produce a private embedding $u^{*}_{\text{DP}}$.
-Finally, images are generated using the corresponding token $S^{*}$>.
-1 Introduction
-The rapid adoption of diffusion models  (Ho, Jain, and Abbeel 2020; Song et al.
-2021; Rombach et al.
-2022) has raised significant privacy and legal concerns.
-These models are vulnerable to privacy attacks, such as membership inference (Duan et al.
-2023), where attackers determine if a specific data point was used for training, and data extraction (Carlini et al.
-2023), which enables reconstruction of training data.
+Finally, images are generated using the corresponding token $S^{*}$.
+I Introduction
+The rapid adoption of diffusion models  [1, 2, 3] has raised significant privacy and legal concerns.
+These models are vulnerable to privacy attacks, such as membership inference [4], where attackers determine if a specific data point was used for training, and data extraction [5], which enables reconstruction of training data.
 This risk is amplified during fine-tuning on smaller, domain-specific datasets, where each record has a greater impact.
-Additionally, reliance on large datasets scraped without consent raises copyright concerns (Vyas, Kakade, and Barak 2023), as diffusion models can reproduce original artworks without credit or compensation.
+Additionally, reliance on large datasets scraped without consent raises copyright concerns [6], as diffusion models can reproduce original artworks without credit or compensation.
 These issues highlight the urgent need for privacy-preserving technologies and clearer ethical and legal guidelines for generative models.
-Differential privacy (DP) (Dwork 2006) is a widely adopted framework for addressing these challenges.
-One standard approach for ensuring DP in deep learning is Differentially Private Stochastic Gradient Descent (DP-SGD) (Abadi et al.
-2016), which modifies traditional SGD by adding noise to clipped gradients.
+Differential privacy (DP) [7] is a widely adopted framework for addressing these challenges.
+One standard approach for ensuring DP in deep learning is Differentially Private Stochastic Gradient Descent (DP-SGD) [8], which modifies traditional SGD by adding noise to clipped gradients.
 However, applying DP-SGD to train diffusion models poses several challenges.
-It introduces significant computational and memory overhead due to per-sample gradient clipping (Hoory et al.
-2021), which is essential for bounding gradient sensitivity (Dwork et al.
-2006; Abadi et al.
-2016).
+It introduces significant computational and memory overhead due to per-sample gradient clipping [9], which is essential for bounding gradient sensitivity [10, 8].
 DP-SGD is also incompatible with batch-wise operations like batch normalization, as these link samples and hinder sensitivity analysis.
 Furthermore, training large models with DP-SGD often leads to substantial performance degradation, particularly under realistic privacy budgets since the required noise scales with the gradient norm.
-Consequently, existing diffusion models trained with DP-SGD are limited to relatively small-scale images.
-(Dockhorn et al.
-2023; Ghalebikesabi et al.
-2023).
-Independent of privacy concerns, Textual Inversion (TI) (Gal et al.
-2023) provides an effective method for adapting diffusion models to specific styles or content without modifying the model.
+Consequently, existing diffusion models trained with DP-SGD are limited to small-scale images [11, 12].
+Independent of privacy concerns, Textual Inversion (TI) [13] effectively adapts diffusion models to specific styles or content without modifying the model.
 Instead, TI learns an external embedding vector that captures the style or content of a target image set, which is then incorporated into text prompts to guide the model’s outputs.
-A key advantage of TI is its ability to compress a style into a compact vector, reducing computational and memory demands while simplifying the application of privacy-preserving mechanisms, as privacy constraints can be applied directly to embeddings rather than the full model.
+A key advantage of TI is its ability to compress a style into a compact vector, reducing computational and memory demands while simplifying privacy mechanisms, as privacy constraints can be applied directly to embeddings rather than the full model.
 Additionally, since TI avoids direct model optimization, it remains efficient and compatible with DP constraints on smaller datasets.
 In this work, we propose a novel privacy-preserving adaptation method for smaller datasets, leveraging TI to avoid the extensive model updates required by DP-SGD.
-Standard TI does not offer formal privacy guarantees, so to address this limitation, we introduce a private variant of TI, called Differentially Private Aggregation via Textual Inversion (DPAgg-TI) and summarize it in Figure 2.
+Standard TI does not offer formal privacy guarantees, so we introduce a private variant, Differentially Private Aggregation via Textual Inversion (DPAgg-TI), summarized in Figure 2.
 Our method decouples interactions among samples by learning a separate embedding for each target image, which are then aggregated into a noisy centroid.
 This approach ensures efficient and secure adaptation to private datasets.
-Our experiments demonstrate the effectiveness of DPAgg-TI, showing that TI remains robust in preserving stylistic fidelity even under privacy constraints.
-Applying our method to a private artwork collection by @eveismyname and Paris 2024 Olympics pictograms (Paris 2024 2023), we show that DPAgg-TI captures nuanced stylistic elements while ensuring privacy.
+Our experiments demonstrate the effectiveness of DPAgg-TI, showing that TI remains robust in preserving stylistic fidelity even under privacy constraints (Figure 1).
+Applying our method to a private artwork collection by @eveismyname and Paris 2024 Olympics pictograms [14], we show that DPAgg-TI captures nuanced stylistic elements while ensuring privacy.
 We observe a trade-off between privacy (controlled by DP parameter $\varepsilon$) and image quality: lower $\varepsilon$ reduces fidelity but maintains the target style under moderate noise.
@@
   • We validate our method on diverse datasets, demonstrating its effectiveness in capturing stylistic elements under privacy constraints.
-2 Background and Related Work
-2.1 Diffusion Models
-Diffusion models (Ho, Jain, and Abbeel 2020; Song et al.
-2021; Song, Meng, and Ermon 2021; Rombach et al.
-2022) leverage an iterative denoising process to generate high-quality images that align with a given conditional input from random noise.
-In text-to-image generation, this conditional input is based on a textual description (a prompt) that guides the model in shaping the image to reflect the content and style specified by the text.
-To convert the text prompt into a suitable conditional format, it is first broken down into discrete tokens, each representing a word or sub-word unit.
+II Background and Related Work
+II-A Diffusion Models
+Diffusion models [1, 2, 15, 3] leverage an iterative denoising process to generate high-quality images that align with a given conditioning input from random noise.
+In text-to-image generation, this conditioning input is based on a textual description (a prompt) that guides the model in shaping the image to reflect the content and style specified by the text.
+To convert the text prompt into a suitable conditioning format, it is first broken down into discrete tokens, each representing a word or sub-word unit.
 These tokens are then converted into a sequence of embedding vectors $v_{i}$ that encapsulate the meaning of each token within the model’s semantic space.
-Next, these embeddings pass through a transformer text encoder, such as CLIP (Radford et al.
-2021), outputting a single text-conditional vector $y$ that serves as the conditioning input.
+Next, these embeddings pass through a transformer text encoder, such as CLIP [16], outputting a single text-conditioning vector $y$ that serves as the conditioning input.
 This vector $y$ is then incorporated at each denoising step, guiding the model to align the output image with the specific details outlined in the prompt.
@@
 At each decreasing timestep $t$, the denoising model, which often utilizes a U-Net structure with cross-attention layers, takes a noisy image $x_{t}$ and text conditioning $y$ as inputs and predicts the noise component $\epsilon_{\theta}(x_{t},y,t)$, where $\theta$ denotes the denoising model’s parameters.
-The predicted noise is then used to make a reverse diffusion step from $x_{t}$ to $x_{t-1}$, iteratively refining the noisy image closer to a coherent output $x_{0}$ that aligns with the text conditional $y$.
-The objective function for a text-conditioned diffusion model, given both the noisy image $x_{t}$ and the text conditioning $y$, is typically a mean squared error (MSE) between the true noise $\epsilon$ and the predicted noise $\epsilon_{\theta}(x_{t},y,t)$.
-The denoising model is therefore trained over the following optimization problem:
+The predicted noise is then used to make a reverse diffusion step from $x_{t}$ to $x_{t-1}$, iteratively refining the noisy image closer to a coherent output $x_{0}$ conditioned on $y$.
+The objective function for a text-conditioned diffusion model, given both the noisy image $x_{t}$ and the text conditioning $y$, is typically a mean squared error between the true noise $\epsilon$ and the predicted noise $\epsilon_{\theta}(x_{t},y,t)$.
+The denoising model is therefore trained over the optimization problem
 $\theta^{*}=\arg\min_{\theta}\mathbb{E}_{x,\epsilon\sim\mathcal{N}(0,I),t\sim[T]}[\|\epsilon-\epsilon_{\theta}(x_{t},y,t)\|^{2}].$ (1)
-Textual Inversion.
-Textual Inversion (TI) (Gal et al.
-2023) is an adaptation technique that enables personalization using a small dataset of typically 3-5 images.
+II-B Textual Inversion
+Textual Inversion (TI) [13] is an adaptation technique that enables personalization using a small dataset of typically 3-5 images.
 This approach essentially learns a new token that encapsulates the semantic meaning of the training images, allowing the model to associate specific visual features with a custom token.
@@
 Let $y_{u}$ denote the text conditioning vector resulting from a prompt containing the token $S$.
-Through gradient descent, TI minimizes the diffusion model loss given in (1) with respect to $u$, while keeping the diffusion model parameters $\theta$ fixed, iteratively refining this embedding to capture the unique characteristics of the training images.
-The resulting optimal embedding $u^{*}$ is formalized as follows:
+Through gradient descent, TI minimizes the diffusion model loss with respect to $u$, instead of the diffusion model parameters $\theta$, which we keep fixed.
+By doing so, we iteratively refine this embedding to capture the unique characteristics of the training images.
+The resulting optimal embedding $u^{*}$ is formalized as
 $u^{*}=\arg\min_{u}\mathbb{E}_{x,\epsilon\sim\mathcal{N}(0,I),t\sim[T]}[\|\epsilon-\epsilon_{\theta}(x_{t},y_{u},t)\|^{2}].$ (2)
 Hence, $u^{*}$ represents an optimized placeholder token $S^{*}$, which can employed in prompts such as “A photo of $S^{*}$ floating in space” or “A drawing of a capybara in the style of $S^{*}$”, enabling the generation of personalized images that reflect the learned visual characteristics.
-2.2 Differential Privacy.
-In this work, we adopt differential privacy (DP) (Dwork et al.
-2006; Dwork 2006) as our privacy framework.
+II-C Differential Privacy
+In this work, we adopt differential privacy (DP) [10, 7] as our privacy framework.
 Over the past decade, DP has become the gold standard for privacy protection in both research and industry.
@@
 When $\delta=0$, we say $\mathcal{M}$ satisfies $\varepsilon$-pure DP or ($\varepsilon$-DP).
-To achieve DP, the Gaussian mechanism is often applied (Dwork, Roth et al.
-2014; Balle and Wang 2018), adding Gaussian noise scaled by the sensitivity of the function $f$ and privacy parameters $\varepsilon$ and $\delta$.
-Specifically, noise with standard deviation $\sigma=\frac{\Delta_{f}\sqrt{2\ln(1.25/\delta)}}{\varepsilon}$ is added to the output [^1] In practice, we use numerical privacy accountant such as (Balle and Wang 2018; Mironov 2017) to calibrate the noise.
-(Balle and Wang 2018), where $\Delta_{f}$ represents $\ell_{2}$-sensitivity of the target function $f(\cdot)$.
-When the context is clear, we may omit the subscript $f$.
-This mechanism enables a smooth privacy-utility tradeoff and is widely used in privacy-preserving machine learning, including in DP-SGD (Abadi et al.
-2016), which applies Gaussian noise during model updates to achieve DP.
-Privacy Amplification by Subsampling.
-Subsampling is a standard technique in DP, where a full dataset of size $n$ is first subsampled to $m$ records without replacement (typically with $m\ll n$) before the privatization mechanism (such as the Gaussian mechanism) is applied.
+To achieve DP, the Gaussian mechanism is commonly applied [17, 18], adding Gaussian noise scaled by the $\ell_{2}$-sensitivity $\Delta$ and privacy parameters $(\varepsilon,\delta)$.
+We add zero-mean isotropic Gaussian noise with standard deviation
+$\sigma\;=\;\frac{\Delta\,\sqrt{2\ln(1.25/\delta)}}{\varepsilon}.$ (3)
+In practice, we calibrate $\sigma$ using numerical privacy accountants (e.g., the analytic Gaussian mechanism and RDP) [18, 19].
+This mechanism enables a smooth privacy–utility trade-off and is widely used in privacy-preserving machine learning, including DP-SGD [8], which adds Gaussian noise to model updates to achieve DP.
+II-C1 Privacy Amplification by Subsampling
+Subsampling is a standard technique in DP, where a full dataset of size $n$ is first subsampled to $m$ records without replacement (typically with $m\ll n$) before the privatization mechanism (e.g.
+the Gaussian mechanism) is applied.
 Specifically, if a mechanism provides $(\varepsilon,\delta)$-DP on a dataset of size $m$, it achieves $(\varepsilon^{\prime},\delta^{\prime})$-DP on the subsampled dataset, where $\delta^{\prime}=\frac{m}{n}\delta$ and
 $\varepsilon^{\prime}=\log\left(1+\frac{m}{n}\left(e^{\varepsilon}-1\right)\right)=O\left(\frac{m}{n}\varepsilon\right).$ (4)
-This result is well-known (Steinke (2022, Theorem 29)), with tighter amplification bounds available for Gaussian mechanisms (Mironov 2017).
-2.3 Private Adaptation of Diffusion Models
+This result is well-known [20, Theorem 29], with tighter amplification bounds available for Gaussian mechanisms [19].
+II-D Private Adaptation of Diffusion Models
 Recent advancements in applying DP to diffusion models have aimed to balance privacy preservation with the high utility of generative outputs.
+Early work on differentially private generative models, such as Chen et al.’s [21] investigation of DP-GANs with model inversion defenses, established foundational principles for protecting generative models from privacy breaches during training.
 Dockhorn et al.
-Dockhorn et al.
-(2023) proposed a Differentially Private Diffusion Model (DPDM) that enables privacy-preserving generation of realistic samples, setting a foundational approach for adapting diffusion processes using DP-SGD.
-Another common strategy involves training a model on a large public dataset, followed by differentially private fine-tuning on a private dataset, as explored by Ghalebikesabi et al.
-(2023).
-While effective in certain contexts, this approach raises privacy concerns, particularly around risks of information leakage during the fine-tuning phase (Tramèr, Kamath, and Carlini 2024).
+[11] proposed a Differentially Private Diffusion Model (DPDM) that enables privacy-preserving generation of realistic samples, setting a foundational approach for adapting diffusion processes using DP-SGD.
+Another common strategy involves training a model on a large public dataset, followed by differentially private fine-tuning on a private dataset [12].
+While effective in certain contexts, this approach raises privacy concerns, particularly around risks of information leakage during the fine-tuning phase [22].
 In response to these limitations, various adaptation techniques have emerged.
-Although not specific to diffusion models, some methods focus on training models on synthetic data followed by DP-constrained fine-tuning, as in the VIP approach (Yu et al.
-2024), which demonstrates the feasibility of applying DP in later adaptation stages.
-Other approaches explore differentially private learning of feature representations (Sander et al.
-2024), aiming to distill private information into a generalized embedding space while maintaining DP guarantees.
+Although not specific to diffusion models, some methods focus on training models on synthetic data followed by DP-constrained fine-tuning, as in Yu et al.
+[23], which demonstrates the feasibility of applying DP in later adaptation stages.
+Other approaches explore differentially private learning of feature representations [24], aiming to distill private information into a generalized embedding space while maintaining DP guarantees.
 Although these adaptations are not yet implemented for diffusion models, they lay essential groundwork for developing secure and efficient privacy-preserving generative models.
 Figure 3: Samples of images used in our style adaptation experiments. Left: artwork by @eveismyname ($n=158$). Right: Paris 2024 Olympic pictograms ($n=47$), © International Olympic Committee, 2023.
-Figure 4: Images generated by Stable Diffusion v1.5 using the prompt “A painting of Taylor Swift in the style of <<@eveismyname>>”, with the embedding <<@eveismyname>> trained using different values of $m$ and $\varepsilon$.
-Figure 5: Images generated by Stable Diffusion v1.5 using the prompt “Icon of a dragon in the style of <<Paris 2024 Pictograms>>”, with the embedding <<Paris 2024 Pictograms>> trained using different values of $m$ and $\varepsilon$.
-3 Differentially Private Adaptation via Textual Inversion
+III Differentially Private Adaptation via Textual Inversion
 Let $x^{(1)},\dots,x^{(n)}$ represent a target dataset of images whose characteristics we wish to privately adapt our image generation towards.
 Instead of training a single token embedding on the entire dataset as in regular TI, we train a separate embedding $u^{(i)}$ on each $x^{(i)}$ to obtain a set of embeddings $u^{(1)},\dots,u^{(n)}$, as illustrated in Figure 2.
-We can formalize the encoding process as follows:
+We can formalize the encoding process as
 $u^{(i)}=\arg\min_{u}\mathbb{E}_{\epsilon\sim\mathcal{N}(0,I),t}[\|\epsilon-\epsilon_{\theta}(x_{t}^{(i)},y_{u},t)\|^{2}].$ (5)
-Then, we can aggregate the embeddings $u^{(1)},\dots,u^{(n)}$ by calculating the centroid.
-The purpose of this aggregation is to limit the sensitivity of the final output to each $x^{(i)}$.
-In order to provide DP guarantees, we also add isotropic Gaussian noise to the centroid.
-We can therefore define the resulting embedding vector $u^{*}_{\text{DP}}$ as follows:
-$u^{*}_{\text{DP}}=\frac{1}{n}\sum_{i=1}^{n}u^{(i)}+\mathcal{N}(0,\sigma^{2}I),$ (5)
-where the minimum $\sigma$ required to provide $(\varepsilon,\delta)$-DP is given by the following expression based on Balle and Wang (2018, Theorem 1):
-$\sigma=\frac{\Delta}{n}\cdot\frac{\sqrt{2\ln(1.25/\delta)}}{\varepsilon}.$ (6)
-In the context of our problem, $\Delta=\sup_{i,j}\|u^{(i)}-u^{(j)}\|$. Since our embedding vectors are directional, we can normalize each $u^{(i)}$, allowing us to set $\Delta=2$.
+Then, we aggregate the embeddings $u^{(1)},\dots,u^{(n)}$ by calculating their centroid and adding isotropic Gaussian noise.
+To ensure bounded sensitivity, we employ a purely directional token embedding (semantics depend only on direction), such as CLIP [16], and $\ell_{2}$-normalize each embedding vector prior to aggregation.
+We can therefore define the resulting centroid $u^{*}_{\text{DP}}$ as
+$u^{*}_{\text{DP}}=\frac{1}{n}\sum_{i=1}^{n}\frac{u^{(i)}}{\|u^{(i)}\|}+\mathcal{N}(0,\sigma^{2}I),$ (6)
+where $\sigma$ is given by (3).
+Under this normalization, the $\ell_{2}$-sensitivity of $u^{*}_{\text{DP}}$ is
+$\Delta=\frac{1}{n}\sup_{u,u^{\prime}}\left\|\frac{u}{\|u\|}-\frac{u^{\prime}}{\|u^{\prime}\|}\right\|=\frac{2}{n},$ (7)
 The noisy centroid embedding $u^{*}_{\text{DP}}$ can then be used to adapt the downstream image generation process.
@@
 Then the standard privacy amplification by subsampling bounds (such as (4)) can be applied.
-Formally, we sample $D_{\text{sub}}\subseteq\{u^{(1)},\dots,u^{(n)}\}$ where $|D_{\text{sub}}|=m$, and compute the output embedding as follows:
-$u^{*}_{\text{DP}}=\frac{1}{m}\sum_{u^{(i)}\in D_{\text{sub}}}u^{(i)}+\mathcal{N}(0,\sigma^{2}I),$ (7)
-where $\sigma$ can be computed numerically for any target $\varepsilon,\delta$ and subsampling rate $\frac{m}{n}$.
-4 Experimental Results
-4.1 Datasets
+Formally, we sample $D_{\text{sub}}\subseteq\{u^{(1)},\dots,u^{(n)}\}$ where $|D_{\text{sub}}|=m$, and compute $u^{*}_{\text{DP}}$ as
+[…]
+where $\sigma$ can be computed numerically using (3) with $\varepsilon^{\prime}$ and $\delta^{\prime}$ from Section II-C1.
+IV Experimental Results
+IV-A Datasets
 We compiled two datasets to evaluate our style adaptation method, specifically selecting content unlikely to be recognized by Stable Diffusion v1.5, our base model.
@@
 This dataset serves as a controlled test for privacy-preserving style transfer on individual artistic collections.
-The second dataset contains 47 pictograms from the Paris 2024 Olympics (Paris 2024 2023), permitted strictly for non-commercial editorial use (International Olympic Committee 2025).
+The second dataset contains 47 pictograms from the Paris 2024 Olympics [14], permitted strictly for non-commercial editorial use [25].
 These pictograms were officially released in February 2023, several months after the release of Stable Diffusion v1.5, ensuring they were absent from the model’s pretraining data.
@@
 Both datasets are used to test the ability of our method to extract and transfer stylistic elements while preserving privacy. Representative samples are shown in Figure 3.
-4.2 Style Transfer Results
-Using both the @eveismyname and Paris 2024 pictograms dataset, we trained TI (Gal et al.
-2023) embeddings on Stable Diffusion v1.5 (Rombach et al.
-2022) using DPAgg-TI.
+Figure 4: Images generated by Stable Diffusion v1.5 using the prompt “A painting of Taylor Swift in the style of <<@eveismyname>>”, with the embedding <<@eveismyname>> trained using different values of $m$ and $\varepsilon$.
+Figure 5: Images generated by Stable Diffusion v1.5 using the prompt “Icon of a dragon in the style of <<Paris 2024 Pictograms>>”, with the embedding <<Paris 2024 Pictograms>> trained using different values of $m$ and $\varepsilon$.
+IV-B Style Transfer Results
+Using both the @eveismyname and Paris 2024 pictograms dataset, we trained TI [13] embeddings on Stable Diffusion v1.5 [3] using DPAgg-TI.
 Our primary goal is to investigate how DP configurations, specifically the privacy budget $\varepsilon$ and subsampling size $m$, affect the generated images quality and privacy resilience.
@@
 Figures 4 and 5 present generated images across two key configurations: (1) regular TI without DP, (2) DPAgg-TI with DP at different values of $m$ and $\varepsilon$.
-We used the same random seed to generate embeddings, subsample images, and sample DP noise for ease of visual comparison between different configurations.
+To ensure reproducibility and fair comparison across all experimental conditions, we fixed the random seed for the entire generation pipeline.
+This design choice allows us to isolate the effect of our style transfer method while holding other sources of randomness constant.
 As with common practice, we set $\delta=1/n$.
-Since $\sigma$ is undefined for $\varepsilon=0$, we demonstrate the results of $\varepsilon\approx 0$, in other words, infinite noise, by setting $\varepsilon=10^{-5}$.
-The purpose of this parameter value is to demonstrate the image generated when $u^{*}_{\text{DP}}$ contains zero information about the target dataset.
 Images generated without DP closely resemble the unique stylistic elements of the target dataset.
-In particular, images adapted using @eveismyname images displayed crisp details and nuanced color gradients characteristic of the artist’s work, while those of Paris 2024 pictograms captured the logo’s original structure.
+In particular, images adapted using @eveismyname images displayed crisp details and nuanced color gradients characteristic of the artist’s work (Figure 4), while those of Paris 2024 pictograms captured the logo’s original structure (Figure 5).
 In contrast, DP configurations introduce a discernible degradation in image quality, with lower epsilon values and smaller subsampling sizes resulting in diminished stylistic fidelity.
+As a no-learning baseline, we consider the limit $\varepsilon\to 0$, under which $u^{*}_{\text{DP}}$ should convey zero information about the target dataset.
+Since $\sigma\propto 1/\varepsilon$ is undefined at $\varepsilon=0$, we approximate this regime by setting $\varepsilon=10^{-5}\approx 0$, which yields effectively infinite noise.
 As $\varepsilon\rightarrow 0$, the resulting token embedding $u^{*}_{\text{DP}}$ gradually loses its semantic meaning, leading to a loss of stylistic fidelity.
 In particular, $y_{u^{*}_{\text{DP}}}$ tends towards $y$ (a conditioning vector independent of the learnable embedding).
-In our results, this manifests as a painting of Taylor Swift devoid of the artist-specific stylistic elements, or a generic icon of a dragon (with color, as opposed to the black and white design of the pictograms).
-With this in mind, $\varepsilon$ can be interpreted as a drift parameter, representing the progression from the optimal $u^{*}_{\text{DP}}$ towards infinity, gradually steering the generated image away from the target style in exchange for stronger privacy guarantees.
+In our results, this manifests as a painting of Taylor Swift devoid of the artist-specific stylistic elements (Figure 4), or a generic icon of a dragon (with color, as opposed to the black and white design of the pictograms, Figure 5).
+To verify this interpretation, we generated images with the same prompts but without the special token $S^{*}$ and compared them to the $\varepsilon\approx 0$ generations.
+The images were visually identical, confirming that at $\varepsilon\approx 0$, the token becomes semantically meaningless and is ignored by the text encoder.
+With this in mind, given a fixed seed, $\varepsilon$ can be interpreted as a drift parameter, representing the progression from the optimal $u^{*}_{\text{DP}}$ towards a semantically meaningless embedding, gradually steering the generated image away from the target style in exchange for stronger privacy guarantees.
 We also observe instances where there is a temporary drop in prompt fidelity (e.g., $m=16,\varepsilon\in[0.5,1]$ in Figure 4 and intermediate $\varepsilon$ values in Figure 5) which restores as $u^{*}_{\text{DP}}$ drifts even further from its optimal value.
 We hypothesize that this is due to drifted $u^{*}_{\text{DP}}$ capturing a different meaning unrelated to the prompt, before losing any meaning that could be interpreted by Stable Diffusion’s text encoder, causing $u^{*}_{\text{DP}}$ to be disregarded from $y_{u^{*}_{\text{DP}}}$ and the prompt fidelity to be restored.
-Another possible explanation is that the temporary drop in prompt fidelity is due to the drift path of $u^{*}_{\text{DP}}$ passing through non-linear regions within embedding space.
-We leave further investigations into this observation for future work.
-Meanwhile, reducing $m$ also reduces the sensitivity of the generated image to $\varepsilon$, as evident by the observation that, on both datasets at $m=4$, (subsampling rate below $0.1$) image generation can tolerate $\varepsilon$ as low as $0.5$ without significant changes in visual characteristics, and retaining stylistic elements of the target dataset at $\varepsilon$ as low as $0.1$.
+Meanwhile, reducing $m$ also reduces the sensitivity of the generated image to with respect to $\varepsilon$, as evident by the observation that, on both datasets at $m=4$, (subsampling rate below $0.1$) image generation can tolerate $\varepsilon$ as low as $0.5$ without significant changes in visual characteristics, and retaining stylistic elements of the target dataset at $\varepsilon$ as low as $0.1$.
 This strong boost in robustness comes at a small price of base style capture fidelity.
@@
 However, the amount of noise introduced by the subsampling process is limited by the individual image embeddings, as a subsample centroid can only stray from the true centroid as much as the biggest outlier in the dataset.
-4.3 Quantitative Evaluation
-User Study
-To evaluate the utility of our approach under different DP and subsampling configurations, we conducted a user study with 25 participants.
-Each participant was shown reference images from the target dataset and asked to compare pairs of generated images, selecting the one that better captured the style of the reference images.
-Images were generated using 10 prompts and adapted TI embeddings for the @eveismyname and Paris 2024 Pictogram datasets, resulting in 20 groups of images.
-Each participant evaluated two groups, one randomly selected from each dataset, with comparisons focusing on model configurations differing by DP noise and subsampling size.
-Survey results, summarized in Table 1 in Appendix A, align with our design goals.
-Participants showed no clear preference between regular TI and DPAgg-TI, suggesting that our privacy-preserving approach maintains perceptual quality.
-As expected, both DP noise and reduced subsampling size degraded style fidelity, consistent with the trade-offs inherent in differential privacy.
-Preferences at $\varepsilon=1$ were split, but subsampling was generally favored, reinforcing its role in reducing noise impact while preserving style.
-Kernel Inception Distance
-The Kernel Inception Distance (KID) (Bińkowski et al.
-2018) is a metric for evaluating generative models by measuring the difference between the distributions of generated and training images in an embedding space.
-To compute KID, images generated by the model and real training images are passed through an Inception network (Szegedy et al.
-2015), and their distributional differences are estimated.
-Unlike the more commonly used Fréchet Inception Distance (FID) (Heusel et al.
-2017), KID is an unbiased estimator of the true divergence between the learned and target distributions (Jayasumana et al.
-2024), making it more suitable for smaller datasets, as in our case.
-We report KID scores for different parameters in Tables  2 and 3 (see Appendix B), showing that DPAgg-TI maintains the style transfer fidelity of TI while ensuring differential privacy.
-Further discussion of these results is also provided in Appendix B.
-4.4 Ablation Study: Textual Inversion with DP-SGD
-A natural question that arises is how well our approach compares to the naive method of applying DP-SGD to regular TI training.
-We therefore integrated DP-SGD into the TI codebase using the Opacus library and trained similar embeddings on the @eveismyname and Paris 2024 datasets.
-We found that in most cases, notably the @eveismyname dataset, the amount of noise required for DP-SGD to achieve a reasonable value of $\varepsilon$ for DP is so high that the resulting embedding contains negligible information about the training dataset.
-In particular, the results for $\varepsilon=1$ are almost indistinguishable to $\varepsilon\approx 0$, as shown in Figure 6.
-We believe that this is simply because DP-SGD is not designed to handle such small datasets in the order of 100 images.
-Additional results can be found in Appendix F.
-Figure 6: Comparing our approach to applying DP-SGD to regular TI using prompts “an icon of a dragon in the style of the Paris 2024 Olympic Pictograms” and “a painting of Taylor Swift in the style of @eveismyname” respectively.
-Note that our method aggregates individual TI embeddings for each training image, whereas the baseline trains a single TI embedding over the entire dataset.
-5 Copyright Protection Implications
-Our proposed mechanism can also be interpreted through the lens of copyright protection.
-This connection is grounded in the framework of Near Access-Freeness (NAF) (Vyas, Kakade, and Barak 2023), which evaluates whether a model’s outputs reveal undue influence from specific data points by comparing them to those from a safe model trained without access to the same data.
-Since DPAgg-TI satisfies $\varepsilon$-DP, it also satisfies $\varepsilon$-NAF, which means the adapted model behaves similarly to one that never saw the private images, under the NAF criterion.
-However, we emphasize that this guarantee holds only within the NAF framework; it does not constitute a general claim about content similarity or legal compliance.
-Crucially, DPAgg-TI is designed to adapt to the style of private images, not their specific content.
-Prior work and legal precedent suggest that style imitation is generally considered fair use and does not constitute infringement (Vyas, Kakade, and Barak 2023).
-Thus, our mechanism aligns with the intended protections of NAF: it avoids memorization while still enabling meaningful personalization and stylistic adaptation.
-We defer the details of copyright protection to Appendix D.
-6 Conclusion
-We presented a differentially private adaptation method for diffusion models based on Textual Inversion, enabling privacy-preserving style transfer without the need for full model fine-tuning.
-By learning per-image embeddings and aggregating them with calibrated noise, our method, DPAgg-TI, achieves strong formal privacy guarantees while maintaining high output fidelity.
-Experiments on private artwork and Paris 2024 pictograms show that DPAgg-TI consistently outperforms DP-SGD, which fails to produce meaningful results under comparable privacy budgets.
-These results highlight the effectiveness of embedding-level adaptation as an efficient and scalable alternative to traditional gradient-based approaches, especially in low-data regimes.
-Unlike DP-SGD, which introduces significant computational overhead and utility degradation, DPAgg-TI is lightweight, modular, and compatible with existing diffusion backbones.
-Our findings suggest that embedding-centric approaches offer a promising direction for privacy-aware personalization, and motivate further research into cross-modal extensions, improved aggregation techniques, and integration with broader privacy-preserving frameworks.
-Acknowledgements
-We sincerely thank @eveismyname for providing her artwork for use in this study.
-We are also grateful to Anwar Hithnawi and Varun Chandrasekaran for their insightful discussions and feedback, as well as to all participants in our user study.
-Sanmi Koyejo acknowledges support by NSF 2046795 and 2205329, IES R305C240046, the MacArthur Foundation, Stanford HAI, OpenAI, and Google.
-Ethical Statement
-The use of images without owner consent raises significant ethical concerns, particularly regarding the exploitation of intellectual property.
-This work introduces a method for visual generative models to adapt to new styles and classes while ensuring privacy and copyright protection for data owners.
-By providing a framework for privacy-preserving adaptation, this technology aims to respect intellectual property and address ethical challenges in generative AI.
-While it does not eliminate the need for consent from data owners, we hope that it represents a step toward balancing innovation with ethical considerations in AI development.
-Beyond creative applications, the proposed method has broader potential uses, including synthetic data generation, privacy-preserving personalization, and fine-tuning diffusion models for private or domain-specific tasks.
-Appendix A User Study
+Figure 6: Samples of image sets used in our user study. Participants were asked to compare 2 images at a time.
+TABLE I: Survey Results.
 Regular TI No Adaptation Unsure
@@
 Paris 2024 15 4 6
-Table 1: Survey Results.
-Figure 7: Samples of image sets used in our user study. Participants are asked to compare 2 images at a time.
-A.1 Study Design and Objective
-The user study aimed to assess the utility of our approach under different DP and subsampling configurations by evaluating the models’ ability to adapt to novel styles.
-The study involved 25 participants, each of whom was tasked with comparing images generated using various configurations and selecting the one that better captured the style of reference images.
-A.2 Experimental Setup
-Participants were shown reference images from two datasets:
-  • The @eveismyname dataset of private artwork.
-  • The Paris 2024 Pictogram dataset.
-For each dataset, 10 prompts were used to generate images, resulting in 20 groups of images (10 prompts per dataset).
-Each group included images generated using the same prompt and dataset but with different model configurations.
-Configurations varied in the addition of DP noise and the size of subsampling.
-  • Original Textual Inversion (TI)
-  • DPAgg-TI ($\varepsilon=\infty$, no DP) w/o subsampling
-  • DPAgg-TI ($\varepsilon=1$) without subsampling
-  • No Adaptation
-  • DPAgg-TI ($\varepsilon=\infty$, no DP) with subsampling $(m=8)$
-  • DPAgg-TI ($\varepsilon=1$) with subsampling $(m=8)$
-  • Style Guidance (SG)
-A.3 Survey Procedure
-Participants were asked to evaluate two groups of images: one randomly selected from the @eveismyname dataset and one from the Paris 2024 Pictogram dataset. For each group:
-  1. 1.
-    Participants were shown reference images from the target dataset.
-  2. 2.
-    They were presented with pairs of images generated using different model configurations for the same prompt.
-  3. 3.
-    Participants selected the image they felt better captured the style of the reference images.
-A.4 Evaluation Metrics
-The study focused on assessing:
-  • Participants’ preference between regular TI and DPAgg-TI for style adaptation.
-  • The impact of DP noise and subsampling size on the perceived utility of style transfer.
-A.5 Results and Analysis
-The results are summarized in Table 1. Key observations include:
-  • Participants showed no clear preference between regular TI and DPAgg-TI in capturing styles for either dataset.
-  • Both DP noise and reduced subsampling size decreased the perceived quality of style transfer.
-  • Preferences were split between configurations with $\varepsilon=1$ with and without subsampling, though subsampling generally had favorable outcomes.
-These findings highlight the trade-off between increased DP robustness and reduced utility, suggesting that the optimal configuration may depend on subjective preferences and specific application requirements.
-Appendix B Kernel Inception Distance
+IV-C User Study
+To evaluate our approach under different DP and subsampling configurations, we conducted a user study with 25 participants.
+The goal was to assess whether DPAgg-TI preserves perceptual quality while offering privacy guarantees.
+IV-C1 Study Design and Setup
+Participants were shown reference images from two datasets: the @eveismyname dataset of private artwork and the Paris 2024 Pictogram dataset.
+For each dataset, we used 10 prompts to generate images, resulting in 20 groups in total.
+Each group included images produced under different configurations, including regular TI, DPAgg-TI with and without DP noise, with and without subsampling ($m=8$), style guidance (see Appendix A), and a no-adaptation baseline.
+IV-C2 Survey Procedure
+Each participant evaluated two groups, one randomly selected from each dataset.
+For each group, participants were first shown the reference images, then asked to compare pairs of generated images produced with different configurations (see Figure 6).
+For each pair, they indicated which image better captured the reference style, or marked the choice as “unsure.”
+IV-C3 Results and Analysis
+Survey results are summarized in Table I.
+Participants showed no clear preference between regular TI and DPAgg-TI, suggesting that our privacy-preserving approach maintains perceptual quality.
+As expected, both DP noise and smaller subsampling size degraded style fidelity, consistent with the trade-offs inherent in differential privacy.
+At $\varepsilon=1$, preferences were split between configurations with and without subsampling, although the subsampling variant was generally favored.
+Overall, the findings highlight that DPAgg-TI achieves perceptual quality comparable to regular TI, while subsampling serves as an effective mechanism to balance privacy and stylistic fidelity.
+IV-D Kernel Inception Distance
+The Kernel Inception Distance (KID) [26] is a metric for evaluating generative models by measuring the difference between the distributions of generated and training images in an embedding space.
+To compute KID, images generated by the model and real training images are passed through an Inception network [27], and their distributional differences are estimated.
+Unlike the more commonly used Fréchet Inception Distance (FID) [28], KID is an unbiased estimator of the true divergence between the learned and target distributions [29], making it more suitable for smaller datasets, as in our case.
+TABLE II: KID scores of DPAgg-TI on @eveismyname dataset for various $\varepsilon$ values ranging from $\varepsilon=10^{-5},0.1,0.5,1.0,5.0$ (including no DP) under different subsampling levels ($m=4,8,16,32$) as well as regular TI (ctrl).
+Reported values are the mean $\pm$ standard deviation over 100 random subsamples.
 $m$ No DP $\varepsilon=$ 5.0 $\varepsilon=$ 1.0 $\varepsilon=$ 0.5 $\varepsilon=$ 0.1 $\varepsilon\approx 0$
@@
 ctrl 0.0314 ± 0.0010
-Table 2: KID scores of DPAgg-TI on @eveismyname dataset for various $\varepsilon$ values ranging from $\varepsilon=10^{-5},0.1,0.5,1.0,5.0$ (including no DP) under different subsampling levels ($m=4,8,16,32$) as well as regular TI (ctrl).
+TABLE III: KID scores of DPAgg-TI on Paris dataset for various $\varepsilon$ values ranging from $\varepsilon=10^{-5},0.1,0.5,1.0,5.0$ (including no DP) under different subsampling levels ($m=4,8,16,32$) as well as regular TI (ctrl).
 Reported values are the mean $\pm$ standard deviation over 100 random subsamples.
@@
 ctrl 0.1383 ± 0.0066
-Table 3: KID scores of DPAgg-TI on Paris dataset for various $\varepsilon$ values ranging from $\varepsilon=1e-5,0.1,0.5,1.0,5.0$ (including no DP) under different subsampling levels ($m=4,8,16,32$) as well as regular TI (ctrl).
-Reported values are the mean $\pm$ standard deviation over 100 random subsamples.
+We report KID scores for different parameters in Tables  II and III.
 Our results indicate that DPAgg-TI preserves the style transfer fidelity of TI while also ensuring differential privacy.
@@
 Upon inspecting the generated images (Figure 7), we hypothesize that the abstract and out-of-distribution nature of the Paris 2024 images poses a challenge for the Inception network, leading to less meaningful feature embeddings.
-This likely inflates the measured embedding distances between generated and reference images, resulting in higher-than-expected KID values.
+This likely inflates the measured embedding distances between generated and reference images, resulting in unusually high KID values.
 For KID evaluations, we used prompts similar to those employed during TI training: “A painting/icon in the style of $S^{*}$”.
 Consistent with the training image captions, these prompts do not specify a subject.
-For each parameter configuration, we generate 100 images and compute KID by repeatedly subsampling the larger of the real and generated sets to match the size of the smaller set, 100 times, then averaging the resulting KID scores.
+For each parameter configuration, we generate 100 images and compute KID by repeatedly subsampling the larger of the real and generated sets to match the size of the smaller set 100 times, then averaging the resulting KID scores.
 Figure 7: Sample of generated images for KID evaluations with respect to the Paris 2024 dataset.
-Appendix C Differentially Private Adaptation via Style Guidance
-C.1 Background: Denoising Diffusion Implicit Models
-Denoising Diffusion Implicit Models (DDIM) sampling (Song, Meng, and Ermon 2021) uses the predicted noise $\epsilon_{\theta}(x_{t},y,t)$ and a noise schedule represented by an array of scalars $\{\alpha_{t}\}_{t=1}^{T}$ to first predict a clean image $\hat{x}_{0}$, then makes a small step in the direction of $\hat{x}_{0}$ to obtain $x_{t-1}$.
-The reverse diffusion process for DDIM sampling can be formalized as follows:
+Figure 8: Comparing our approach to applying DP-SGD to regular TI using prompts “an icon of a dragon in the style of the Paris 2024 Olympic Pictograms” and “a painting of Taylor Swift in the style of @eveismyname” respectively.
+Note that our method aggregates individual TI embeddings for each training image, whereas the baseline trains a single TI embedding over the entire dataset.
+IV-E Ablation Study: Textual Inversion with DP-SGD
+A natural question that arises is how well our approach compares to the naive method of applying DP-SGD to regular TI training.
+We therefore integrated DP-SGD into the TI codebase using the Opacus library and trained similar embeddings on the @eveismyname and Paris 2024 datasets.
+We found that in most cases, notably the @eveismyname dataset, the amount of noise required for DP-SGD to achieve a reasonable value of $\varepsilon$ for DP is so high that the resulting embedding contains negligible information about the training dataset.
+In particular, the results for $\varepsilon=1$ are almost indistinguishable to $\varepsilon\approx 0$, as shown in Figure 8.
+We believe that this is simply because DP-SGD is not designed to handle such small datasets in the order of 100 images.
+Additional results can be found in Appendix C.
+V Discussion
+V-A Copyright Protection Implications
+Our proposed mechanism can also be interpreted through the lens of copyright protection.
+This connection is grounded in the framework of Near Access-Freeness (NAF) [6], which evaluates whether a model’s outputs reveal undue influence from specific data points by comparing them to those from a safe model trained without access to the same data.
+Modern generative models typically produce outputs via randomized sampling.
+Leveraging this inherent randomness, Vyas et al.
+[6] introduced NAF as a metric to quantify the similarity between a model’s output and copyrighted content.
+The key idea is to compare the output distribution of a potentially infringing model to that of a safe model—one trained without access to the target content.
+A canonical example is the leave-one-out-safe model, trained on the full dataset excluding $x$.
+Since $\mathsf{safe}(x)$ lacks access to $x$, the probability that it generates content resembling $x$ is expected to be small; any such resemblance is considered fortuitous.
+Definition 2 (Near Access-Freeness [6]).
+Let $\mathcal{C}$ be a set of copyrighted samples and $\mathcal{W}$ a set of generative models.
+Given a mapping $\mathsf{safe}:\mathcal{C}\rightarrow\mathcal{W}$ and a divergence measure $\Delta$, we say a model $w\in\mathcal{W}$ is $k_{y}$-near access-free (or $k_{y}$-NAF) on prompt $y\in\mathcal{Y}$ if for every $x\in\mathcal{C}$,
+$\Delta\!\left(P_{w}(\cdot|y)\,\|\,P_{\mathsf{safe}(x)}(\cdot|y)\right)\leq k_{y}.$
+If $k_{y}=0$, the model is indistinguishable from a safe model, meaning any resemblance to copyrighted material is by random chance.
+More generally, a small $k_{y}$ suggests the model is unlikely to generate outputs resembling $x$ with higher probability than a model that has never seen $x$.
+NAF is closely related to concepts in DP.
+Depending on the divergence measure $\Delta$, NAF resembles different DP variants—for example, $\varepsilon$-DP when $\Delta=\Delta_{\mathrm{max}}$ [10], and $(1,\varepsilon)$-Rényi DP when $\Delta=\Delta_{\mathrm{KL}}$.
+Translating DP to generative models yields this definition:
+Definition 3 (Differentially Private Generation (DPG)).
+Let $S$ and $S^{\prime}$ be neighboring datasets.
+Denote by $P_{S}(\cdot|y)$ the distribution over outputs generated by a model trained on, or adapted from, $S$ with algorithm $\mathcal{A}$, where randomness includes both training and generation.
+The generation satisfies $\varepsilon$-Differentially Private Generation ($\varepsilon$-DPG) if for every $y\in\mathcal{Y}$,
+$\Delta\!\left(P_{S}(\cdot|y)\,\|\,P_{S^{\prime}}(\cdot|y)\right)\leq\varepsilon.$
+Here, neighboring datasets differ by a single data point (or privacy unit).
+If the training process is $\varepsilon$-DP, then the outputs naturally satisfy $\varepsilon$-DPG via the data processing inequality.
+One benefit of DPG is the flexibility to add noise during generation rather than training, potentially improving the utility–privacy tradeoff.
+However, there are notable distinctions: $\varepsilon$-DP offers protection under arbitrary post-processing and multiple outputs, whereas $\varepsilon$-DPG only guarantees privacy for single outputs.
+Also, under DP, the trained model can be released, but under DPG, only the outputs are safe to share.
+Elkin-Koren et al.
+[30] highlight further differences: NAF is one-sided—comparing a model to a fixed safe reference—whereas DPG is symmetric.
+This asymmetry in NAF can enable better utility.
+Additionally, NAF allows more flexibility in choosing the safe model, which can be exploited in algorithm design.
+Since DPAgg-TI satisfies $\varepsilon$-DP, it also satisfies $\varepsilon$-NAF under the leave-one-out-safe model.
+Within the NAF framework, this means the adapted model behaves similarly to one that never saw the private images.
+Importantly, this guarantee is meaningful only within NAF; it does not imply broader legal immunity or empirical indistinguishability from the original content.
+However, it allows us to argue that any close resemblance between outputs and private training data is no more likely than would be expected from a model with no access to that data.
+Finally, the goal of DPAgg-TI is to adapt to the style of private image sets, not their precise content.
+This distinction matters: pure style imitation (without reproducing protectable expression or “substantive elements”) is often argued to be non-infringing in many creative contexts, though the legal status is jurisdiction- and fact-dependent; particularly in artistic and creative contexts.
+As discussed in Elkin-Koren et al.
+[30] and legal analyses such as Carlini et al.
+[5], generating new content in the style of a work, without reproducing its substantive elements, is generally not considered copyright infringement.
+Therefore, the use of DPAgg-TI to learn and reproduce stylistic attributes does not contradict the spirit or intent of the NAF framework.
+Instead, it offers a promising direction for responsibly fine-tuning generative models on private or copyrighted sources while respecting both privacy and intellectual property boundaries.
+Remark 1 (Scope of Protection and Artist-Level Extension).
+We provide record-level DP: it limits leakage or reconstruction of any individual private image, not an artist’s entire style.
+Consequently, it yields a corresponding NAF-style guarantee at the per-image level (under the chosen safe reference).
+This interpretation should be understood strictly within NAF and does not constitute a general copyright compliance claim.
+The DP guarantee continues to apply under targeted prompts: conditioning on detailed descriptions can increase the likelihood of reproducing a specific private work by at most an $e^{\varepsilon}$ factor (up to $\delta$).
+Artist-level (user-level) DP is conceptually possible by treating each artist as one unit and privately aggregating artist-level embeddings (e.g., via a DP mean mechanism), but typically requires stronger noise and may reduce utility; we leave a full exploration to future work.
+V-B Limitations
+DPAgg-TI is designed for the low-data, strong-privacy regime, where the number of private images is small ($n\approx 100$) and per-record protection with $\varepsilon<5$ matters.
+For large datasets with moderate subsampling, DP-SGD on the full model may become more efficient and could provide better utility.
+We explicitly position our method for scenarios where DP-SGD is known to struggle: strong privacy guarantees with limited training data.
+We acknowledge that in moderate privacy regimes ($5\leq\varepsilon\leq 10$) with larger batch sizes and careful tuning, DP-SGD with parameter-efficient fine-tuning methods might perform better than our approach.
+However, in our experiments, applying DP-SGD to regular TI even at $\varepsilon=20$ with carefully tuned hyperparameters (learning rate, batch size, scheduler) still required prohibitively high noise levels to satisfy the privacy accountant, preventing meaningful learning.
+Moreover, DP-SGD typically involves multiple training iterations, so the effective noise further accumulates due to composition across epochs, making convergence extremely difficult in the strong-privacy regime ($\varepsilon<5$) with approximately 100 images.
+VI Conclusion
+We presented a differentially private adaptation method for diffusion models based on Textual Inversion, enabling privacy-preserving style transfer without the need for full model fine-tuning.
+By learning per-image embeddings and aggregating them with calibrated noise, our method, DPAgg-TI, achieves strong formal privacy guarantees while maintaining high output fidelity.
+Experiments on private artwork and Paris 2024 pictograms show that DPAgg-TI consistently outperforms DP-SGD, which fails to produce meaningful results under comparable privacy budgets.
+These results highlight the effectiveness of embedding-level adaptation as an efficient and scalable alternative to traditional gradient-based approaches, especially in low-data regimes.
+Unlike DP-SGD, which introduces significant computational overhead and utility degradation, DPAgg-TI is lightweight, modular, and compatible with existing diffusion backbones.
+Our findings suggest that embedding-centric approaches offer a promising direction for privacy-aware personalization, and motivate further research into cross-modal extensions, improved aggregation techniques, and integration with broader privacy-preserving frameworks.
+Ethical Statement
+The use of images without owner consent raises significant ethical concerns, particularly regarding the exploitation of intellectual property.
+This work introduces a method for visual generative models to adapt to new styles and classes while ensuring privacy and copyright protection for data owners.
+By providing a framework for privacy-preserving adaptation, this technology aims to respect intellectual property and address ethical challenges in generative AI.
+While it does not eliminate the need for consent from data owners, we hope that it represents a step toward balancing innovation with ethical considerations in AI development.
+Beyond creative applications, the proposed method has broader potential uses, including synthetic data generation, privacy-preserving personalization, and fine-tuning diffusion models for private or domain-specific tasks.
+Acknowledgments
+We sincerely thank @eveismyname for providing her artwork for use in this study.
+We are also grateful to Anwar Hithnawi and Varun Chandrasekaran for their insightful discussions and feedback, as well as to all participants in our user study.
+Sanmi Koyejo acknowledges support by NSF 2046795 and 2205329, IES R305C240046, the MacArthur Foundation, Stanford HAI, OpenAI, and Google.
+This work was supported in part by the National Research Foundation of Korea (NRF) grant funded by the Korea government (MSIT) (No.
+RS-2025-23525649).
+Appendix A Differentially Private Adaptation via Style Guidance
+A-A Background: Denoising Diffusion Implicit Models
+Denoising Diffusion Implicit Models (DDIM) sampling [15] uses the predicted noise $\epsilon_{\theta}(x_{t},y,t)$ and a noise schedule represented by an array of scalars $\{\alpha_{t}\}_{t=1}^{T}$ to first predict a clean image $\hat{x}_{0}$, then makes a small step in the direction of $\hat{x}_{0}$ to obtain $x_{t-1}$.
+The reverse diffusion process for DDIM sampling can be formalized as
 $\hat{x}_{0}=\frac{x_{t}-\sqrt{1-\alpha_{t}}\epsilon_{\theta}(x_{t},y,t)}{\sqrt{\alpha_{t}}}$ (9)
 $x_{t-1}=\sqrt{\alpha_{t-1}}\hat{x}_{0}+\sqrt{1-\alpha_{t-1}}\epsilon_{\theta}(x_{t},y,t).$ (10)
-C.2 Implementation
-We extend our approach to style guidance (SG) by leveraging the framework of Universal Guidance (Bansal et al.
-2024).
+A-B Implementation
+We extend our approach to style guidance (SG) by leveraging the framework of Universal Guidance [31].
 Specifically, we focus on CLIP-based style guidance, which optimizes the similarity between the CLIP embeddings of a target image and the generated image.
@@
 The aggregated embedding $u^{*}_{\text{DP}}$ is then incorporated into the reverse diffusion process as a style guide.
-Let $x_{c}$ denote the target style image, $x_{t}$ the noisy image at step $t$, and $\mathcal{E}(\cdot)$ the CLIP image encoder. The forward guidance process is defined as follows:
+Let $x_{c}$ denote the target style image, $x_{t}$ the noisy image at step $t$, and $\mathcal{E}(\cdot)$ the CLIP image encoder. The forward guidance process is defined as
 $\hat{\epsilon}_{\theta}(x_{t},y,t)=\epsilon_{\theta}(x_{t},y,t)+w\sqrt{1-\alpha_{t}}\nabla_{x_{t}}\ell_{\cos}(\mathcal{E}(x_{t}),\mathcal{E}(\hat{x}_{0})),$ (11)
@@
 This ensures privacy-preserving style transfer while maintaining high stylistic fidelity.
-C.3 Style Transfer Results
+A-C Style Transfer Results
 Figure 9: Attempts of using universal guidance to generate drawings of Taylor Swift and icons of the Eiffel Tower in the styles of @eveismyname and Paris 2024 Pictograms respectively.
@@
 Although Bansal et al.
-(2024) proposed remedies, namely backward guidance and per-step self-recurrence, these proved insufficient for our application.
+[31] proposed remedies, namely backward guidance and per-step self-recurrence, these proved insufficient for our application.
 Additionally, the CLIP embeddings may not retain enough stylistic detail after the aggregation.
-C.4 Ablation
-To better understand the limited effectiveness of style guidance in our experiments, despite its success in (Bansal et al.
-2024), we applied our approach to a dataset of 143 paintings from Van Gogh’s Saint-Paul Asylum, Saint-Rémy collection(Innat 2022) (Figure 10).
+A-D Ablation
+To better understand the limited effectiveness of style guidance in our experiments, despite its success in Bansal et al.
+[31], we applied our approach to a dataset of 143 paintings from Van Gogh’s Saint-Paul Asylum, Saint-Rémy collection[32] (Figure 10).
 Unlike the @eveismyname and Paris 2024 datasets, it is highly likely that Stable Diffusion has been trained on these images.
 Additionally, Bansal et al.
-(2024) demonstrated successful adaptation towards the style of Van Gogh’s Starry Night as a single reference image, making this dataset a reasonable interpolation between their successful results and our more limited findings.
+[31] demonstrated successful adaptation towards the style of Van Gogh’s Starry Night as a single reference image, making this dataset a reasonable interpolation between their successful results and our more limited findings.
 Without DP noise or subsampling, we obtained reasonable style transfer results, as shown in Figure 11.
@@
 Figure 11: Images generated by Stable Diffusion v1.5 with style guidance towards Van Gogh’s Saint-Paul Asylum, Saint-Rémy collection using prompts “A painting of Taylor Swift (left) / the Eiffel Tower (center) / a tree (right)”.
-Appendix D Copyright Protection
-Modern generative models typically produce outputs via randomized sampling.
-Leveraging this inherent randomness, Vyas, Kakade, and Barak (2023) introduced Near Access-Freeness (NAF) as a metric to quantify the similarity between a model’s output and copyrighted content.
-The key idea is to compare the output distribution of a potentially infringing model to that of a safe model – one trained without access to the target content.
-Formally, let $\mathsf{safe}$ be a mapping from a data point $x\in\mathcal{C}$ (where $\mathcal{C}$ is the collection of copyrighted samples) to a generative model $\mathsf{safe}(x)\in\mathcal{W}$ that is trained without using $x$.
-A canonical example is the leave-one-out-safe model, trained on the full dataset excluding $x$.
-Since $\mathsf{safe}(x)$ does not have access to $x$, the probability that it generates content resembling $x$ is exponentially small.
-Any such resemblance is considered fortuitous.
-Formally, the NAF criterion is defined as follows:
-Definition 2 (Near Access-Freeness (Vyas, Kakade, and Barak 2023)).
-Let $\mathcal{C}$ be a set of copyrighted samples and $\mathcal{W}$ a set of generative models.
-Given a mapping $\mathsf{safe}:\mathcal{C}\rightarrow\mathcal{W}$ and a divergence measure $\Delta$, we say a model $w$ is $k_{y}$-near access-free (or $k_{y}$-NAF) on prompt $y\in\mathsf{Y}$ if for every $x\in\mathcal{C}$,
-$\Delta\left(p(\cdot|y)\,\|\,\mathsf{safe}_{x}(\cdot|y)\right)\leq k_{y}.$
-If $k_{y}=0$, the model is indistinguishable from a safe model, meaning any resemblance to copyrighted material is by random chance.
-More generally, a small $k_{y}$ suggests the model is unlikely to generate outputs resembling $x$ with higher probability than a model that has never seen $x$.
-D.1 Connection to Differential Privacy
-NAF is closely related to concepts in Differential Privacy (DP) (Elkin-Koren et al.
-2023).
-Depending on the divergence measure $\Delta$, NAF resembles different DP variants – for example, $\varepsilon$-DP when $\Delta=\Delta_{\mathrm{max}}$ (Dwork et al.
-2006), and $(1,\varepsilon)$-Rényi DP when $\Delta=\Delta_{\mathrm{KL}}$.
-Translating DP to generative models yields the following definition:
-Definition 3 (Differentially Private Generation (DPG)).
-Let $S$ and $S^{\prime}$ be neighboring datasets.
-Denote by $P_{S}(\cdot|y)$ the distribution over outputs generated by a model trained on, or adapted from, $S$ with algorithm $\mathcal{A}$, where randomness includes both training and generation stages.
-The generation is said to satisfy $\varepsilon$-Differentially Private Generation ($\varepsilon$-DPG) if for every $y\in\mathcal{Y}$,
-$\Delta\left(P_{S}(\cdot|y)\,\|\,P_{S^{\prime}}(\cdot|y)\right)\leq\varepsilon.$
-Here, neighboring datasets differ by a single data point (or privacy unit).
-If the training process is $\varepsilon$-DP, then the outputs naturally satisfy $\varepsilon$-DPG via the data processing inequality.
-One benefit of DPG is the flexibility to add noise during generation rather than training, potentially improving the utility-privacy tradeoff.
-However, there are notable distinctions.
-$\varepsilon$-DP offers protection under arbitrary post-processing and multiple outputs, whereas $\varepsilon$-DPG only guarantees privacy for single outputs.
-Also, under DP, the trained model can be released, but under DPG, only the outputs are safe to share.
-Elkin-Koren et al.
-(2023) highlight further differences: NAF is one-sided—comparing a model to a fixed safe reference—whereas DPG is symmetric.
-This asymmetry in NAF can enable better utility.
-Additionally, NAF allows more flexibility in choosing the safe model, which can be exploited in algorithm design.
-Given these conceptual overlaps, both DP-SGD based training and our proposed private adaptation method DPAgg-TI satisfies $\varepsilon$-DP, so they naturally satisfy $\varepsilon$-NAF with the leave-one-out safe model.
-We emphasize that this guarantee is meaningful only within the formal framework of NAF.
-It does not imply broader legal immunity or empirical indistinguishability from the original content.
-However, within this framework, satisfying $\varepsilon$-NAF allows us to argue that any close resemblance between outputs and private training data is no more likely than would be expected from a model that never had access to that data.
-This theoretical grounding supports the privacy and safety claims of our adaptation method.
-Importantly, the goal of DPAgg-TI is to adapt to the style of a private image set—not its precise content.
-This distinction matters: style transfer is widely considered to fall under the doctrine of fair use, particularly in artistic and creative contexts.
-As discussed in Elkin-Koren et al.
-(2023) and further elaborated in legal analysis such as Carlini et al.
-(2023), generating new content in the style of a work, without reproducing its substantive elements, is generally not considered copyright infringement.
-Therefore, the use of DPAgg-TI to learn and reproduce stylistic attributes does not contradict the spirit or intent of the NAF framework.
-Instead, it offers a promising direction for responsibly fine-tuning generative models on private or copyrighted sources while respecting both privacy and intellectual property boundaries.
-Appendix E Computational Cost Comparisons
-Direct comparisons of computational cost across methods are inherently challenging due to differing training paradigms, optimization procedures, and parameter settings.
-Nonetheless, to provide a concrete sense of scale, we report representative computational costs for each method based on experiments conducted using a Stable Diffusion v1.5 model on a single NVIDIA A100 GPU.
-Below we summarize both training and inference overheads (the number of steps are optimized for each setup):
+Appendix B Computational Cost Comparisons
+Direct comparisons of computational cost across methods are challenging due to differing training paradigms (per-image optimization vs.
+dataset-level training), optimization procedures, and privacy accounting.
+Nonetheless, to provide a concrete sense of scale, we report representative costs measured using Stable Diffusion v1.5 on a single NVIDIA A100 GPU (Tables IV and V).
+For each method, we tuned the number of optimization steps to reach its best utility under the target privacy budget.
+Sequential vs. batched execution
+The per-image runtime reported in Table IV corresponds to a sequential implementation that optimizes each textual-inversion embedding independently.
+In practice, we can optimize multiple embeddings jointly by batching several images at once (and optionally subsampling the private set per update), which reduces the effective wall-clock cost and avoids the naive linear scaling implied by “minutes per image $\times\;n$.”
+TABLE IV: Training cost comparison across methods.
+Overhead from DP-SGD is relatively modest due to the low-dimensional embedding being optimized.
+N/A for SG means nothing is trained aside from the base model.
 Method Steps Batch Size Time Memory Usage
@@
 SG N/A N/A N/A N/A
-Table 4: Training cost comparison across methods.
-Overhead from DP-SGD is relatively modest due to the low-dimensional embedding being optimized.
-N/A for SG means nothing is trained aside from the base model.
+TABLE V: Inference cost comparison across methods.
 Method Steps Batch Size Time Memory Usage
@@
 SG (no DP, DPAgg-SG) 500 1 $\sim$30 min 17 GB
-Table 5: Inference cost comparison across methods.
-Appendix F Additional Style Transfer and Ablation Results
+When DP-SGD can be faster
+DP-SGD amortizes computation across the dataset and can be faster wall-clock-wise for larger $n$ and/or moderate privacy budgets.
+In contrast, our method is designed for low-data personalization with strong per-record guarantees, and it offers a practical advantage in dynamic settings: it supports incremental updates to the private set (e.g.
+adding or removing images) without retraining a large set of model parameters, whereas DP-SGD-style training typically requires rerunning optimization to reflect such changes.
+We view these approaches as complementary, targeting different operating regimes.
+Appendix C Additional Style Transfer and Ablation Results
 Figure 12: Images generated by Stable Diffusion v1.5 using the prompt “A painting of Taylor Swift in the style of <<@eveismyname>>”, with the embedding <<@eveismyname>> trained using DPAgg-TI (with different subsample sizes $m$) and TI with DP-SGD using different values of $\varepsilon$.